### v21

# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

Abstract
[…]
Sinhala is the native language of the Sinhalese people who make up the largest ethnic group of Sri Lanka. The language belongs to the globe-spanning language tree, Indo-European. However, due to poverty in both linguistic and economic capital, Sinhala, in the perspective of Natural Language Processing tools and research, remains a resource-poor language which has neither the economic drive its cousin English has nor the sheer push of the law of numbers a language such as Chinese has. A number of research groups from Sri Lanka have noticed this dearth and the resultant dire need for proper tools and research for Sinhala natural language processing. However, due to various reasons, these attempts seem to lack coordination and awareness of each other. The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize contributions of their peers. As such, we shall be uploading this paper to arXiv and perpetually update it periodically to reflect the advances made in the field.
Index Terms:
Sinhala, Natural Language Processing, Resource Poor Language
1 Introduction
Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 16 million people [5, 6]. When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7]. To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that Sinhala language belongs to the same the Indo-European language tree [8, 9]. However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch. Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [10, 11, 12, 13, 14, 15, 16]. By extension, this makes Sinhala Script a member of the Aramaic family of scripts [17, 18]. Thus by inheritance, Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [19]. We show the evolutionary eras of the Sinhala script in Appendix A. It should be noted that the modern Sinhala language have loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [20]. Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [21, 22]), modern natural language processing tools for the Sinhala language are scarce [23].
Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages. To achieve this end, NLP systems operate at different levels [24, 25, 26]. A graphical representation of NLP layers and application domains are shown in Figure 1. On one hand, according to Liddy [25], these systems can be categorized into the following layers; phonological, morphological, lexical, syntactic, semantic, discourse, and pragmatic. The phonological layer deals with the interpretation of language sounds. As such, it consists of mainly speech-to-text and text-to-speech systems. In cases where one is working with the written text of the language rather than speech, it is possible to replace this layer with tools which handle Optical Character Recognition (OCR) and language rendering standards (such as Unicode [27]). The morphological layer analyses words at their smallest units of meaning. As such, analysis on word lemmas and prefix-suffix-based inflection are handled in this layer. Lexical layer handles individual words. Therefore tasks such as Part of Speech (PoS) tagging happens here. The next layer, syntactic, takes place at the phrase and sentence level where grammatical structures are utilized to obtain meaning. Semantic layer attempts to derive the meanings from the word level to the sentence level. Starting with Named Entity Recognition (NER) at the word level and working its way up by identifying the contexts they are set in until arriving at overall meaning. The discourse layer handles meaning in textual units larger than a sentence. In this, the function of a particular sentence maybe contextualized within the document it is set in. Finally, the pragmatic layer handles contexts read into contents without having to be explicitly mentioned [24, 25]. Some forms of anaphora (coreference) resolution [28, 29, 30, 31, 32] fall into this application.
Figure 1: NLP layers and tasks [24]
On the other hand, Wimalasuriya and Dou [26] categorize NLP tools and research by utility. They introduce three categories with increasing complexity; Information Retrieval (IR), Information Extraction (IE), and Natural Language Understanding (NLU). Information Retrieval covers applications, which search and retrieve information which are relevant to a given query. For pure IR, tools and methods up-to and including the syntactic layer in the above analysis are used. Information Extraction, on the other hand, extracts structured information. The difference between IR and IE is the fact that IR does not change the structure of the documents in question. Be them structured, semi-structured, or unstructured, all IR does is fetching them as they are. In comparison, IE, takes semi-structured or unstructured text and puts them in a machine readable structure. For this, IE utilizes all the layers used by IR and the semantic layer. Natural Language Understanding is purely the idea of cognition. Most NLU tasks fall under AI-hard category and remain unsolved [24]. However, with varying accuracy, some NLU tasks such as machine translation [^2] This is, however, not without the criticism of being nothing more than a Chinese room [33] rather than true NLU. are being attempted. The pragmatic layer of the above analysis belongs to the NLU tasks while the discourse layer straddles information extraction and natural language understanding [24].
The objective of this paper is to serve as a comprehensive survey on the state of natural language processing resources for the Sinhala language. The initial structure and content of this survey are heavily influenced by the preliminary surveys carried out by de Silva [23] and Wijeratne et al. [24]. However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [34, 35] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available. We also discuss how the non-compliance of policies to put data and code online [36], after the research is concluded and the paper is published, have negatively impacted the growth and sustainability of Sinhala NLP. Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively. For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^3] https://github.com/lknlp/lknlp.github.io. This manuscript is at version $5.3.0$. The latest version of the manuscript can be obtained from arXiv [^4] https://arxiv.org/abs/1906.02358 or ResearchGate [^5] http://bit.ly/31AhvvR.
Figure 6 in Appendix B shows the most prolific researchers in the domain of Sinhala NLP. The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that researcher has authored. The edges between two researchers are labeled with the number of Sinhala NLP papers the relevant pair of researchers have co-authored. When selecting authors, we have applied a threshold of 3 Sinhala NLP publications. Given that the objective of the visualization is to portray the co-operation between researchers, we have also added the strongest edge that connects each researcher to the rest of the researchers in the graph. The few isolated nodes are researchers that have authored at least 3 Sinhala NLP publications but do not have any coauthored papers with anyone else on the graph. We have also added labels to clusters in cases where all or the majority of researchers in those clusters have the same affiliation. It is observable that the cluster from the Department of Computer Science & Engineering, University of Moratuwa is the most prolific in Sinhala NLP research.
Figure 7 in Appendix B shows the probability of studies from the institutions to which the most prolific authors from Figure 6 are affiliated citing institutions of the same set. We observe both interesting and disturbing trends which we discuss in the figure caption
Figure 8 in Appendix B shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper. If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication. For example, the paper Building a wordnet for Sinhala [37] is counted for both Nisansa de Silva and Gihan Dias. If a single paper contributes to more than one research area, that paper is counted for all of the research areas to which it contributes. For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [23] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
The remainder of this survey is organized as follows; Section 2 introduces some important properties and conventions of the Sinhala language which are important for the development and understanding of Sinhala NLP. Section 3 discusses the various tools and research available for Sinhala NLP. In this section we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work. We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka. Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work. Finally, Section 5, concludes the survey.
2 Properties of the Sinhala Language
TABLE I: Examples for grammatical cases and inflection of animate common nouns
TABLE II: Examples for grammatical cases and inflection of inanimate common nouns
Note that the grammatical cases of Auxiliary and Vocative do not exist for inanimate nouns in Sinhala.
Before moving on to discussing Sinhala NLP resources, we shall give a brief introduction to some of the important properties of Sinhala language, which impact the development of Sinhala NLP resources. Sinhala grammar has two forms: written (literary) and spoken. These forms differ from each-other in their core grammatical structures [38, 1, 39]. The written form strictly adheres to the SOV (Subject, Object, and Verb) configuration [40, 41]. Further, in the written form, subject-verb agreement is enforced [42] such that, in order to be grammatically correct, the subject and the verb must agree in terms of: gender (male/female), number (singular/plural) and person (1st/2nd/3rd). However, in spoken Sinhala, the SOV order can be neglected [43] and male singular 3rd person verb can be used for all nouns [42]. Sinhala is also a head-final language, where the complements and modifiers would appear before their heads [44] this is similar to that of English and dissimilar to that of French. In total, according to Abhayasinghe [45], there are 25 types of simple sentence structures in Sinhala. Similar to many Indo Aryan languages, animacy plays a major role in Sinhala grammar in syntactic and semantic roles [46, 47, 48]. Comparative studies done by Noguchi [49] and by Miyagishi [50, 39] have found that animacy extends its influence from phrase level to sentence level in Sinhala (e.g., Usage of post-positions [40, 51]). On this matter, Table I explains grammatical cases and inflections of animate common nouns while Table II explains grammatical cases and inflections of inanimate common nouns. We provide a comparative analysis of parsing the very simple English sentence “I eat a red apple” and its Sinhala, Hindi, and French translations in Fig 2. English and French parsing was done using the Stanford Parser [^6] http://nlp.stanford.edu:8080/parser/. Hindi parsing was done using the IIIT-Hyderabad Parser [^7] http://ltrc.iiit.ac.in/analyzer/ and the study by Singh et al. [52].
(a) English
(b) Sinhala
(c) Hindi
(d) French
Figure 2: Parse trees for the sentence “I eat a red apple” in four languages.
Herath et al. [53, 54] argue that pure Sinhala words did not have suffixes and that adding suffixes was incorporated to Sinhala after 12th century BC with the influx of Sanskrit words. With this, they declare Sinhala to have to following types of words:
  1. 1.
    Suffixes
  2. 2.
    Nouns
  3. 3.
    Cases
  4. 4.
    Verbs
  5. 5.
    Conjunctions and articles
  6. 6.
    Adjectives
  7. 7.
    Demonstratives, Interrogatives, and negatives
  8. 8.
    Particles and prefixes
They further divide nouns into five groups: material, agentive, common, abstract, and proper. In addition to these, they also introduce compound nouns.We show the noun categorization proposed by Herath et al. [53] in Table III.
TABLE III: Noun categorization by Herath et al. [53]
Herath et al. [54] categorize Sinhala suffixes along the attributes of: gender, number, definiteness, case, and conjunctive. They further claim that there are 3 types of suffixes: Suf1 adds gender, number, and definiteness; Suf2 adds case; and Suf3 adds conjunctive. Conjunctive is claimed to be equivalent to too and and in English. We show an extension of the suffix structure proposed by Herath et al. [54] in Table IV. In their analysis on register variation (vocabulary) of 60 languages, Li et al. [55] observes that Sinhala exhibits homogeneity between 0.5 and 1.0 in the three corpora considered: CC (macro-web register - Common Crawl), TW (social media register - Twitter), and WK (Wikipedia register - March 2020).
TABLE IV: Extension of the suffix structure proposed by Herath et al. [54]
Number = Singular (S) / Plural (P)
Definite = Definite (D) / Indefinite (I) / Undecided (U)
Case = Nominative (N) / Accusative (A) / Dative (Da) / Genitive (G) / Instrumentive (In) / Auxiliary (Au) / Locative (L) / Ablative (Ab)
Conjun = with suffix th (Y) / without suffix th (N)
3 Sinhala NLP resources
In this section we generally follow the structure shown in Figure 1 for sectioning. However, in addition to that, we also discuss topics such as available corpora, other data sets, dictionaries, and WordNets. We focus on NLP tools and research rather than the mechanics of language script handling [56, 57, 58, 59, 60, 61]. One of the earliest attempts on Sinhala NLP was done by Herath et al. [62]. However, progress on that project has been minimal due to the limitations of their time. The later work by Nandasara [63] has not caught much of the advances done up to the time of its publication. Given that it was a decade old by the time the first edition of this survey was compiled, we observe the existence of many new discoveries in Sinhala NLP which have not been taken into account by it. A review on some challenges and opportunities of using Sinhala in computer science was done by Nandasara and Mikami [64]. At this point, it is worth noting that the largest number of studies in Sinhala NLP has been on optical character recognition (OCR) rather than on higher levels of the hierarchy shown in Figure 1. On the other hand, the most prolific single project of Sinhala NLP we have observed so far is an attempt to create an end-to-end Sinhala-to-English translator [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 19, 76, 77, 78, 79, 80, 81, 82]. The sinling [^8] https://github.com/ysenarath/sinling/ on GitHub contains a collection of tools for handling Sinhala NLP including a Tokenizer, Stemmer, PoS Tagger, Morphological joiner, and a Morphological splitter. The package is available for Python via pypi.
Tamil, the other official language of Sri Lanka is also a resource-poor language. However, due to the existence of larger populations of Tamil speakers worldwide, including but not limited to economic powerhouses such as India, there are more research and tools available for Tamil NLP tasks [24]. Therefore, it is rational to notice that Sinhala and Tamil NLP endeavours can help each other. Especially, given the above fact, that these are official languages of Sri Lanka, results in the generation of parallel data sets in the form of official government documents and local news items. A number of researchers make use of this opportunity. We shall be discussing those applications in this paper as well. Further, there have been some fringe implementations, which bridge Sinhala with other languages such as Japanese [22, 83, 84, 85, 9].
3.1 Corpora
For any language, the key for NLP applications and implementations is the existence of adequate corpora. On this matter, a relatively substantial Sinhala text corpus [^9] https://osf.io/a5quv/ was created by Upeksha et al. [86, 87] by web crawling. It was later extended by adding Jathaka Stories [^10] https://bit.ly/JathakaTxt and more web-crawled news articles [^11] https://bit.ly/3osodBj. Later a smaller Sinhala news corpus [^12] https://osf.io/tdb84/ was created by de Silva [23]. Both of the above corpora are publicly available. However, none of these come close to the massive capacity and range of the existing English corpora. A word corpus of approximately 35,000 entries was developed by Weerasinghe et al. [88]. But it does not seem to be online anymore. Guzmán et al. [89] provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^13] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^14] http://bit.ly/2ZaQFZo.  Wijeratne and de Silva [90] have publicly released [^15] https://bit.ly/2GEI4d6 a massive corpus of text and stop words taken from a decade of Sinhala Facebook posts. While the stop word extraction of Wijeratne and de Silva [90] algorithmic, Lakmal et al. [91] has introduced a manually curated stop word list [^16] http://bit.ly/3wFi0Wf. A parallel corpus of Sinhala and English was collected by Bañón et al. [92] containing 217,407 sentences and available to download from their website [^17] https://www.paracrawl.eu/. However, the later audit by Caswell et al. [93] raised issues on the quality of that data set. A parallel corpus [^18] https://github.com/kdissa/comparable-corpus of aligned Sinhala-English documents and sentences obtained from crawling the web was released by Sachintha et al. [94]. The study by Warusawithana et al. [95] created a refined version [^19] https://bit.ly/3JF77t9 of the OpenSLR-52 speech corpus for Sinhala [^20] https://openslr.org/52 by Kjartansson et al. [96] [^21] A later further overview on the corpus is also available [97].Dhananjaya et al. [98] created sin-cc-15M corpus2222223https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus. A text corpus collected from the Sihala blog Kalaya [^23] http://www1.kalaya.org/ along with the relevant code is available on Github [^24] https://bit.ly/KalayaCorpus.
As for Sinhala-Tamil corpora, Hameed et al. [99] claim to have built a sentence-aligned Sinhala-Tamil parallel corpus and Mohamed et al. [100] claim to have built a word aligned Sinhala-Tamil parallel corpus. However, at the time of writing this paper, neither of them was publicly available. A very small Sinhala-Tamil aligned parallel corpus created by Farhath et al. [101] using order papers of government of Sri Lanka is available to download [^25] http://bit.ly/2HTMEmeVasantharajan and Thayasivam [102] used Printed Character Recognition (PCR) to create a large-scale Tamil-Sinhala-English parallel corpus which they claimed to be available on Github  [^26] https://github.com/Chaarangan/tamizhinet-corpus. Their follow-up work [103] on adapting the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents resulted in an improved version of the corpus which also can be found on Github [^27] https://bit.ly/3VIK2Lj.
3.2 Data Sets
Specific data sets for Sinhala, as expected, is scarce. However, a Sinhala PoS tagged data set [104, 105, 106] is available to download from github [^28] http://bit.ly/2Krhrbv. Further, a Sinhala NER data set created by Manamini et al. [107] is also available to download from github [^29] http://bit.ly/2XrwCoKLiyanage et al. [108] analyzed Sinhala fastText and Word2Vec in the context of cross-lingual embedding spaces.
Facebook has released FastText [109, 110, 111] models for the Sinhala language trained using the Wikipedia corpus. They are available as both text models [^30] http://bit.ly/2JXAyL8 and binary files [^31] http://bit.ly/2JY5J9c. Using the above models by Facebook, Lakmal et al. [91] have created an extended FastText model trained on Wikipedia, News, and official government documents. The binary file [^32] http://bit.ly/2WowH0h of the trained model is available to be downloaded. Herath et al. [112], Herath and Medagoda [113] have compiled a report on the Sinhala lexicon for the purpose of establishing a basis for NLP applications. A comparative analysis of Sinhala word embedding has been conducted by Lakmal et al. [91]. A similar study was then conducted by Silva [114] analyising the progress that had been made in NLP in the intervening years. A dataset [^33] https://github.com/LIRNEasia/MisinformationCorpusSinhala consisting of 3576 Sinhala documents drawn from Sri Lankan news websites and tagged (CREDIBLE, FALSE, PARTIAL or UNCERTAIN) was published by Jayawickrama et al. [115] A benchmark data set for Sinhala spell correction was created by  Sonnadara et al. [116] which they put online on Github [^34] https://github.com/chason94/SinNeuSpellCorrector. De Saa and Ranathunga [117] has released a data set [^35] https://bit.ly/3FTNaMZ for Sinhala hate speech detection which consists of comments from youtube. Another data set consisting of Sinhala hate speech comments pulled from Facebook is available [^36] https://bit.ly/3j9r3Za on Kaggle but it does not have an accompanying paper. Similarly, Perera et al. [118] have released a hate annotated dataset [^37] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. A text-to-speech data set [^38] https://github.com/pnfo/sinhala-tts-dataset with 3300 Sinhala sentences with 7.5 hours of recordings was released by the Path Nirvana Foundation. The open parallel corpus (OPUS) [^39] https://opus.nlpl.eu/ by Tiedemann [119] curates the largest collection of parallel data between Sinhala and other languages. Zhang et al. [120] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100404040https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT). Thier code is also available on GitHub [^41] https://github.com/bzhangGo/zero. Further, Tiedemann [121] created an NMT benchmark named Tatoeba MT Challenge [^42] https://github.com/Helsinki-NLP/Tatoeba-Challenge using OPUS data based on Tatoeba [^43] https://tatoeba.org/. The English-Sinhala data set can also be directly downloaded from a link on Github [^44] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data.
A benchmark Sinhala-English translation data set named FLORES [^45] https://tinyurl.com/flores200dataset was created by [89]. This includes a 600k+ Sinhala-English subtitle pairs [^46] http://bit.ly/2KsFQxm initially collected by [122], 45k+ Sinhala-English sentence pairs from GNOME [^47] http://bit.ly/2Z8q0fo, KDE [^48] http://bit.ly/2WLY6bI, and Ubuntu [^49] http://bit.ly/2wLVZGt. They further provided two monolingual corpora for Sinhala. Those were a 155k+ sentences of filtered Sinhala Wikipedia [^50] http://bit.ly/2EQZ7oM and 5178k+ sentences of Sinhala common crawl [^51] http://bit.ly/2ZaQFZo. In addition to the data set, they also have made their code [^52] https://github.com/facebookresearch/flores publicly available. This work was then extended by NLLB Team et al. [123] of which the data and code is available on github under the No Language Left Behind (NLLB) project  [^53] https://github.com/facebookresearch/fairseq/tree/nllb. Very importantly, they also provide a toxicity data set for Sinhala [^54] https://github.com/facebookresearch/flores/tree/main/toxicity which can be used for hate speech detection in or by translation or otherwise.
Jenarthanan et al. [124] introduced the ACTSEA dataset [^55] https://bit.ly/3oZoziI which contains Sinhala and Tamil tweets annotated with emotions. They have 318,308 Sinhala tweets annotated.  Dhananjaya et al. [98] have made publicly available three data sets developed for their study: 1) They have taken the corpus created by Sachintha et al. [94] and derived a Sinhala News source classification data set [^56] https://bit.ly/3PBTyfo, They have taken the corpus created by de Silva [23] and derived a Sinhala News source classification data set [^57] https://bit.ly/3PBTyfo, and 3) They have taken the corpus created by Upeksha et al. [86] and derived a Sinhala writing style classification data set [^58] https://bit.ly/3PARuUS. A data set of Sinhala swear and/or obscene words is publicly available in both Unicode [^59] https://bit.ly/3R1bdOc and Singlish [^60] https://bit.ly/3S3Z1O2 formats.  Fernando et al. [125] has created aligned corpora for Sinhala–English, Tamil–English, and Sinhala–Tamil language pairs. The corpora are available on github [^61] https://bit.ly/3DB50no as well as the code [^62] https://bit.ly/3FGcYOL for their document and sentence alignment task. Buddhika et al. [126] introduce a crowd sourcing tool that they named Voicer to collect speech data. They claim that the tool is open source and that they had created a Sinhala Speech corpus of 10 hours with 39 different sentences in the banking domain. Neither the code nor the data is publicly linked in the research paper. However, a subsequent work by Hellarawa and Thayasivam [127] uses this data set. Thus, it can be assumed that this data set may be available if the authors are contacted.
Ranasinghe et al. [128] created a benchmark dataset they named SOLD [^63] https://huggingface.co/datasets/sinhala-nlp/SOLD: Sinhala Offensive Language Dataset which contains 10,000 posts from Twitter annotated both at sentence-level and token-level with the two classes offensive and not offensive. In the same paper, they also introduce the dataset SemiSOLD [^64] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach. The relevant code is also available on Github  [^65] https://github.com/Sinhala-NLP/SOLD. Sinhala is included in the 2800+ language metadata set composed by van Esch et al. [129]. This can be used for comparative analysis of Sinhala against other languages in the data set. Ruder et al. [130] created the multi-language dataset named XTREME-UP [^66] https://github.com/google-research/xtreme-up which contains Sinhala data sets for OCR and Transliteration tasks. Pratap et al. [131] created the The Massively Multilingual Speech (MMS) [^67] https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set which contains Sinhala for the Spoken Language Identification (LID) task. A data set [^68] https://github.com/laurieburchell/open-lid-dataset and a model for the Language Identification (LID) task including Sinhala were created by Burchell et al. [132]. A large Sinhala-English dictionary with $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs has been made publicly available on github [^69] https://github.com/kasunw22/sinhala-para-dict/tree/main by Wickramasinghe and De Silva [133]. In their follow-up work [134] they used that dictionary data to obtain Sinhala-English embedding alignment. The relevant code and data can be accessed on GitHub [^70] https://bit.ly/3t3SKu7. Nguyen et al. [135] released CulturaX [^71] https://huggingface.co/datasets/uonlp/CulturaX a cleaned dataset for training Large Language Models. It contains 753,655 Sinhala documents with 880,289,097 tokens. However, it should be noted that this only amounts to 0.01% of their multilingual data set. In comparison, Tamil has a 0.07% share, Hindi has a 0.27% share, and English has a 45.13% share. Kudugunta et al. [136] released MADLAD-400727272https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages. For Sinhala, they report that they have 788K noisy docs, 349.2K clean docs, 22.1M noisy sentences, 16M clean sentences, 3.4B noisy tokens, and 1.9B clean tokens. Further, for the translation tasks they report that they have 7363378 sentence pairs where one of the languages in each pair is Sinhala.
XL-sum [^73] https://github.com/csebuetnlp/xl-sum/tree/master Hasan et al. [137] and M3LS [^74] https://github.com/anubhav-jangra/M3LS Verma et al. [138] are text summarizing datasets which contains Sinhala data collected from the BBC website [^75] https://www.bbc.com/sinhala. Respectively, they have 3,414 and 10,148 Sinhala text documents and their summaries. Being a multi-modal dataset, M3LS additionally contains the relevant images collected from the same source. A dynamic word-level Sinhala Sign Language video dataset of 50 classes has been published by Charuka et al. [139]. It contains 1110 videos. The OSCAR [^76] https://oscar-project.org/ dataset [140, 141] contains web mined Sinhala data consisting of 172,755,385 words in 301,066 documents.
3.3 Dictionaries
A necessary component for the purpose of bridging Sinhala and English resources are English-Sinhala dictionaries. The earliest and most extensive Sinhala-English dictionary available for consumption was by Malalasekera [142]. However, this dictionary is locked behind copyright laws and is not available for public research and development. This copyright issue is shared with other printed dictionaries [143, 144, 145, 146, 147, 148] as well. The dictionary by Kulatunga [149] is publicly available for usage through an online web interface but does not provide API access or means to directly access the data set. The largest publicly available English-Sinhala dictionary data set is from a discontinued FireFox plug-in EnSiTip [150] which bears a more than passing resemblance to the above dictionary by Kulatunga [149]Hettige and Karunananda [68] claim to to have created a lexicon to help in their attempt to create a system capable of English-to-Sinhala machine translation. A review on the requirements for English-Sinhala smart bilingual dictionary was conducted by Samarawickrama and Hettige [151]. The study by Wickramasinghe and De Silva [133] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs. Both their code and the two versions of the dictionary are publicly available on github [^77] https://github.com/kasunw22/sinhala-para-dict/tree/main. A Sinhala-Sinhala dictionary named Wahara [^78] http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
There exists the government-sponsored trilingual dictionary [152], which matches Sinhala, English, and Tamil. However, other than a crude web interface on the ministry website, there is no efficient API or any other way for a researcher to access the data on this dictionary. Weerasinghe and Dias [153] have created a multilingual place name database for Sri Lanka which may function both as a dictionary and a resource for certain NER tasks.
3.4 WordNets
WordNets [154] are extremely powerful and act as a versatile component of many NLP applications. They encompass a number of linguistic properties which exist between the words in the lexicon of the language including but not limited to: hyponymy, hypernymy, synonymy, and meronymy. Their uses range from simple gazetteer listing applications [26] to information extraction based on semantic similarity [155, 156] or semantic oppositeness [157]. An attempt has been made to build a Sinhala Wordnet by Wijesiri et al. [37]. For a time it was hosted on [158] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al. [159] have cloned to use in their application uploaded on github [^79] https://github.com/jseanm1/aruthSWSD. However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product. Another effort to build a Sinhala Wordnet was initiated by Welgama et al. [160] independently from above; but it too have stopped progression even before achieving the completion level of Wijesiri et al. [37].
3.5 Morphological Analyzers
As shown in Fig 1, morphological analysis is a ground level necessary component for natural language processing. Given that Sinhala is a highly inflected language [43, 42, 23], a proper morphological analysis process is vital. The earliest attempt on Sinhala morphological analysis we have observed are the studies by Herath et al. [53, 54]. They are more of an analysis of Sinhala morphology rather than a working tool. As such we discussed the observations and conclusions of these works at Section 2. It is also worth to note that these works predates the introduction of Sinhala unicode and thus use a transliteration of Sinhala in the Latin alphabet.
The next attempt by Herath et al. [161] creates a modular unit structure for morphological analysis of Sinhala. Much later, as a step on their efforts to create a system with the ability to do English-to-Sinhala machine translation, Hettige and Karunananda [65] claim to have created a morphological analyzer (void of any public data or code), which links to their studies of a Sinhala parser [66] and computational grammar [19]Hettige et al. [77] further propose a multi-agent System for morphological analysis. Welgama et al. [162] attempted to evaluate machine learning approaches for Sinhala morphological analysis. Yet another independent attempt to create a morphological parser for Sinhala verbs was carried out by Fernando and Weerasinghe [163]. Later, another study, which was restricted to morphological analysis of Sinhala verbs was conducted by Dilshani and Dias [164]. There was no indication on whether this work was continued to cover other types of words. Further, other than this singular publication, no data or tools were made publicly accessible. Nandathilaka et al. [165] proposed a rule based approach for Sinhala lemmatizing. The work by Welgama et al. [166] claim to have set a set of gold standard definitions for the morphology of Sinhala Words; but given that their results are not publicly available, further usage or confirmation of these claims cannot not be done. The table V provides a comparative summery of the discussion above. The combined study introduced a rule-based stemmer [167] and a tokenizer [168] for Sinhala. A later work by Kumarasinghe et al. [169] named SinMorphy used a comprehensive vocabulary of Sinhala words to conduct rule-based morphological analysis on Sinhala. Ekanayaka et al. [170] compared the effectiveness of using RNN, LSTM, and GRU for Sinhala morphological analysis of Sinhala and Sinhlish deatasets and found the BiGRU gives the highest accuracy of 87.96%.
TABLE V: Morphological Analyzers comparison
Base: Rule-based (RB) / Machine Learning (ML)
Able to Handle Part of Speech (Handles): Yes (Y) / No (N)
Outputs: Yes (Y) / No (N) / No Information (O)
Abbreviations: Nouns (Nu), Verbs (Ve), Adjectives (Aj), Adverbs (Av), Function Words (Fn), Root (R), Person (P), Number (Nb), Gender (G), Article (A), Case (C)
Base Modus Operandi Handles Output
Nu Ve Aj Av Fn R P Nb G A C
 Hettige and Karunananda [65] RB Finite State Automata Y Y N N N Y Y Y Y Y Y
 Hettige et al. [77] RB Agent-based Y Y Y Y Y N N N N N N
 Nandathilaka et al. [165] RB N/A Y N N N N Y N N N N N
 Welgama et al. [162] ML Morfessor algorithm Y Y Y Y Y Y N N N N N
 Fernando and Weerasinghe [163] RB Finite State Transducer N Y N N N Y Y Y Y N Y
 Dilshani and Dias [164] RB N/A N Y N N N O O O O O O
3.6 Part of Speech Taggers
The next step after morphological analysis is Part of Speech (PoS) tagging. The PoS tags differ in number and functionality from language to language. Therefore, the first step in creating an effective PoS tagger is to identifying the PoS tag set for the language. This work has been accomplished by Fernando et al. [104] and Dilshani et al. [105]. Expanding on that, Fernando et al. [104] has introduced an SVM Based PoS Tagger for Sinhala and then Fernando and Ranathunga [106] give an evaluation of different classifiers for the task of Sinhala PoS tagging. While here it is obvious that there has been some follow up work after the initial foundation, it seems, all of that has been internal to one research group at one institution as neither the data nor the tools of any of these findings have been made available for the use of external researchers. Several attempts to create a stochastic PoS tagger for Sinhala has been done with the studies by Herath and Weerasinghe [171]Jayaweera and Dias [172], and Jayasuriya and Weerasinghe [173] being the most notable.
Within a single group which did one of the above stochastic studies [172], yet another set of studies was carried out to create a Sinhala PoS tagger starting with the foundation of Jayaweera and Dias [174] which then extended to a Hidden Markov Model (HMM) based approach [175] and an analysis of unknown words [176, 177]. Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [178]. A RESTFul PoS tagging web service created by Jayaweera and Dias [179] using the above research can still be accessed [^80] http://bit.ly/2F0jKid via POST and GET. A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al. [180]. The study by Kothalawala et al. [181] discussed the data availability problem in NLP with a Sinhala POS tagging experiment among others.  Withanage and Silva [182] proposed a stochastic POS tagger based on a small 10,000 word corpus drawn from Facebook and Twitter.  Wijerathna [183] used Support Vector Machines (SVM) to tag Sinhala text with 30 the PoS tags that were proposed by Fernando et al. [104]. The study by Sathsarani et al. [184] compared rule-based and stochastic models against deep learning models in the task of Sinhala PoS tagging.
3.7 Parsers
The PoS tagged data then needs to be handed over to a parser. This is an area which is not completely solved even in English due to various inherent ambiguities in natural languages. However, in the case of English, there are systems which provide adequate results [185] even if not perfect yet. The Sinhala state of affairs, is that, the first parser for the Sinhala language was proposed by Hettige and Karunananda [66] with a model for grammar [19]. The study by Liyanage et al. [43] is concentrated on the same given that they have worked on formalizing a computational grammar for Sinhala. While they do report reasonable results, yet again, do not provide any means for the public to access the data or the tools that they have developed. Kanduboda [42] have worked on Sinhala differential object markers relevant for parsing.
The first attempt at a Sinhala parser, as mentioned above, was by Hettige and Karunananda [66] where they created prototype Sinhala morphological analyzer and a parser as part of their larger project to build an end-to-end translator system. The function of the parser is based on three dictionaries: Base Dictionary, Rule Dictionary, and Concept Dictionary. They are built as follows:
  • The Base Dictionary: prakurthi (base words), nipatha (prepositions), upasarga (prefixes), and vibakthi (Irregular Verbs).
  • The Rule Dictionary: inflection rules used to generate various forms of verbs and nouns from the base words.
  • The Concept Dictionary: synonyms and antonyms for the words found in the base dictionary.
Parsers are, in essence, a computational representation of the grammar of a natural language. As such, in building Sinhala parsers, it is crucial to create a computational model for Sinhala grammar. The first such attempt was taken by Hettige and Karunananda [19] with special consideration given to Morphology and the Syntax of the Sinhala language as an extension to their earlier work [66]. Here, it is worthy to note that, unlike in their earlier attempt [66], where they explicitly mentioned that they are building a parser, in this study [19], they use the much conservative claim of building a computational grammar. Under Morphology, they again handled Sinhala inflection. Their system is based on a Finite State Transducer (FST) and Context-Free Grammar (CFG) where they they modeled 85 rules for nouns and 18 rules for verbs. The specific implementation is more partial to a rule-based composer rather than parser. It is also worthy to note that this system could only handle simple sentences which only contained the following 8 constituents: Attributive Adjunct of Subject, Subject, Attributive Adjunct of Object, Object, Attributive Adjunct of Predicate, Attributive Adjunct of the Complement of Predicate, Complement of Predicate, and Predicate. With these, they propose the following grammar rules for Sinhala:
S = Subject Akkyanaya
Subject = SimpleSubject | ComplexSubject
ComplexSubject = SimpleSubject ConSub
SimpleSubject = Noun | Adjective Noun
ConSub = Conjunction SimpleSubject
Akkyanaya = VerbP | Object VerbP
Object = SimpleObject | ComplexObject
ComplexObject = Conjunction SimpleObject
SimpleObject = Noun | Adjective Noun
VerbP = Verb | Adverb Verb
The later work by Liyanage et al. [43] also involves formalizing a computational grammar for Sinhala. They claim that Sinhala can have any order of words in practice. However, they do not note that this is happening because practices of the spoken language, which does not share the strong SOV conventions of the written language, are slowly seeping into written text. However, they do make note of how Sinhala grammar is modeled as a head-final language [44]. They propose the Sinhala Noun Phrase ($NP$) to be defined as shown in equation 1 where $NN$ is a noun which can be of types: common noun ($N$), pronoun ($PrN$) or proper noun ($PropN$). The adjectival phrase ($ADJP$) is then defined as as shown in equation 2 where: $Det$ is a Determiner, $Adj$ is the adjective, and $Deg$ is an optional operator Degrees which can be used to intensify the meaning of the adjective in cases where the adjective is qualitative. While they note that according to Gunasekara [186], there has to three classes of adjectives (qualitative, quantitative, and demonstrative), they do not implement this distinction in their system. Similarly, they propose Sinhala Verb Phrase ($VP$) to be defined as shown in equation 3 where $V$ is a single verb. They here note that they are ignoring compound verbs and auxiliary verbs in their grammar. The adverbial phrases ($ADVP$) are then recursively defined as as shown in equation 4.
$NP=[ADJP][NN]$ (1)
$ADJP=\bigg{[}[Det]\Big{[}[Deg][Adj]\Big{]}\bigg{]}$ (2)
$VP=[ADVP][V]$ (3)
$ADVP=\Bigg{[}[NP]\bigg{[}[ADVP]\Big{[}[Deg][ADV]\Big{]}\bigg{]}\Bigg{]}$ (4)
Similar to Hettige and Karunananda [19], the work by Liyanage et al. [43] also builds a CFG for Sinhala covering 10 out of the 25 types of simple sentence structures in Sinhala reported by Abhayasinghe [45]. This parser is unable to parse sentences where inanimate subjects do not consider the number. Further, sentences which contain, compound verbs, auxiliary verbs, present participles, or past participles cannot be handled by this parser. If the verbs have imperative mood or negation those too cannot be handled by this. Non-verbal sentences which end with adjectives, oblique nominals, locative predicates, adverbials, or any other language entity which is not a verb cannot be handled by this parser.
The study by Kanduboda [42] covers not the whole of Sinhala parsing but analyzes a very specific property of Sinhala observed by Aissen [187] which states that it is possible to notice Differential Object Marking (DOM) in Sinhala active sentences. Kanduboda [42] define this as the choice of /wa/ and /ta/ object markers. They further observe three unique aspects of DOM in Sinhala: (a) it is only observed in active sentences which contain transitive verbs, (b) it can occur with accusative marked nouns but not with any other cases, (c) it exists only if the sentence has placed an animate noun in the accusative position. They do a statistical analysis and provide a number of short gazetteer lists as appendixes. However, they observe that further work has to be done for this particular language rule in Sinhala given that they found some examples which proved to be exceptions to the general model which they proposed.
3.8 Named Entity Recognition Tools
TABLE VI: NER system comparison
* Denotes a baseline.
$F_{1}$ to $F_{11}$ denotes Context Words, Word Prefixes and Suffixes, Length of the Word, Frequency of the Word, First Word/ Last Word of a Sentence, (POS) Tags, Gazetteer Lists, Clue Words, Outcome Prior, Previous Map, and Cutoff Value
CRF ME Features
$F_{1}$ $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$ $F_{6}$ $F_{7}$ $F_{8}$ $F_{9}$ $F_{10}$ $F_{11}$
 Dahanayaka and Weerasinghe [188] Yes Yes* Yes Yes No No No No No No No No No
 Senevirathne et al. [189] Yes No Yes Yes Yes No Yes No No Yes No Yes No
 Manamini et al. [107] Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes Yes
As shown in Fig 1, once the text is properly parsed, it has to be processed using a Named Entity Recognition (NER) system. The first attempt of Sinahla NER was done by Dahanayaka and Weerasinghe [188]. Given that they were conducting the first study for Sinhala NER, they based their approach on NER research done for other languages. In this, they gave prominent notice to that of Indic languages. On that matter, they were the first to make the interesting observation that NER for Indic languages (including, but not limited to Sinhala) is more difficult than that of English by the virtue of the absence of a capitalization mechanic. Following prior work done on other languages, they used Conditional Random Fields (CRF) as their main model and compared it against a baseline of a Maximum Entropy (ME) model. However, they only use the candidate word, Context Words around the candidate word, and a simple analysis of Sinhala suffixes as their features.
The follow up work by Senevirathne et al. [189] kept the CRF model with all the previous features but did not report comparative analysis with an ME model. The innovation introduced by this work is a richer set of features. In addition to the features used by Dahanayaka and Weerasinghe [188], they introduced, Length of the Word as a threshold feature. They also introduced First Word feature after observing certain rigid grammatical rules of Sinhala. A feature of clue Words in the form of a subset of Context Words feature was first proposed by this work. Finally, they introduced a feature for Previous Map which is essentially the NE value of the preceding word. Some of these feature extractions are done with the help of a rule-based post-processor which utilizes context-based word lists.
The third attempt at Sinhala NER was by Manamini et al. [107] who dubbed their system Ananya. They inherit the CRF model and ME baseline from the work of Dahanayaka and Weerasinghe [188]. In addition to that, they take the enhanced feature list of Senevirathne et al. [189] and enrich it further more. They introduce a Frequency of the Word feature based on the assumption that most commonly occurring words are not NEs. Thus, they model this as a Boolean value with a threshold applied on the word frequency. They extend the  First Word feature proposed by Senevirathne et al. [189] to a First Word/ Last Word of a Sentence feature noting that Sinhala grammar is of SOV configuration. They introduce a (PoS) Tag feature and a gazetteer lists based feature keeping in line with research done on NER in other languages. They formally introduce clue Words, which was initially proposed as a sub-feature by Dahanayaka and Weerasinghe [188], as an independent feature. Utilizing the fact that they have the ME model unlike Dahanayaka and Weerasinghe [188], they introduce a complementary feature to Previous Map named Outcome Prior, which uses the underlying distribution of the outcomes of the ME model. Finally, they introduce a Cutoff Value feature to handle the over-fitting problem.
The table VI provides a comparative summary of the discussion above. It should be noted that all three of these models only tag NEs of types: person names, location names and organization names. The Ananya system by Manamini et al. [107] is available to download at GitHub  [^81] http://bit.ly/2XrwCoK. The data and code for the approaches by Dahanayaka and Weerasinghe [188] and by Senevirathne et al. [189] are not accessible to the public. Azeez and Ranathunga [190] proposed a fine-grained NER model for Sinhala building on their earlier work on NER [107] and PoS tagging [106]Anuruddha [191] proposed a method based on reinforcement learning for Sinhala NER. A Sinhala NER system restricted to the sports domain was introduced by Wijesinghe and Tissera [192], where they attempted to utilise classical machine learning models. The work by Mallikarachchi et al. [193] used support vector machines to detect Sinhala named entities of types person, location and, organization.
3.9 Semantic Similarity
A Sinhala semantic similarity measure has been developed for short sentences by Kadupitiya et al. [194]. This work has been then extended by Kadupitiya et al. [195] for the application use case of short answer grading. Data and tools for these projects are not publicly available. A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [196] where they demonstrate their results for Sinhala and Tamil. A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al. [197].
3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [198]. A smaller implementation of Sinhala news classification has been attempted by de Silva [23]. As mentioned in Section 3.2, their news corpus is publicly available [^82] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [199]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [200] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [201] for human interpretability of Sinhala document classification. However, they too do not provide access corpus. Nanayakkara and Ranathunga [202] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [203, 204] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [205]Demotte and Ranathunga [206] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [207]. The attempt by Sameemdeen and Selvanthan [208] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [202]Gunasekara and Haddela [203]Lakmali and Haddela [199], and Buddhika et al. [209]. Then they propose active learning [210, 211] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [212] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [26]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [213]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [214]. The study by [215] used Genetic Algorithm on Lucene [^83] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [216] uses adapter-based [217, 218, 219, 220, 221, 222] fine-tuning on XLM-R [223], for code-mixed and code-switched text. Kirindage and Godewithana [224] used LDA [225] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [226] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [98]. Both their code [^84] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus8585853https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^86] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^87] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^88] https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [227] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [228] used capsule-based methods recommended by Senevirathne et al. [207] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [229] compared Word2Vec, FastText and Doc2vec [230] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [231] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [232] based on their prior work [233]. A word2vec based tool [^89] http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [234]Demotte et al. [235] proposed a sentiment analysis system based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [236], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [207] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [237] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [238] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [239] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [240]. The work by Karunarathne [241] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [242] conducted a multi model analysis on classifying Sinhala songs by emotion.
The work by Jayawickrama et al. [243, 244] used the data set released by Wijeratne and de Silva [90] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [245] and compared the results obtained with their data set against that of Senevirathne et al. [207]Aththanayaka and Naleer [246] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset [^90] https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [124]. The dataset contains 318,308 Sinhala tweets annotated with emotions.
3.12 Hate Speech Detection
Given the low-resource nature of the Sinhala language, there does not seem to be a distinction between the Hate Speech Detection task, Offensive Speech Detection task, and Inappropriate Speech Detection task in the literature for Sinhala. More often, the terms are used interchangeably even though there is a clear distinction between the definitions of them in the literature focused on high-resource languages [247]. Given the lack of differentiation in the Sinhala literature, we too list studies conducted on one or multiple of those tasks in this section. As mentioned in the Section 3.2, a large annotated data set for Sinhala hate speech detection was created by NLLB Team et al. [123].
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [248]. A feature model and a data set [^91] https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [117]Sandaruwan et al. [249] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [250] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [251] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [252] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [253] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [254] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [255] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [256] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [257]. The study by Shalinda and Munasinghe [258] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [259] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [260] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [261] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [118] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset [^92] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [262] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [263] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala.
After the release of the SOLD dataset by Ranasinghe et al. [128] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [264] used traditional machine learning algorithms to predict Sinhala hate speech. Ojo et al. [265] used mBERT to identify hate and offensive content in Sinhala under HASOC 2023, Task 1A [^93] https://hasocfire.github.io/hasoc/2023/task1.html which included the SOLD dataset. Ranasinghe and Zampieri [266] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [267] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. The study by Narayan et al. [268] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%.
3.13 Fake News Detection
A dataset consisting of Sinhala documents drawn from Sri Lankan news websites was published by Jayawickrama et al. [115] along with the benchmark misinformation classification models. A hybrid approach to detect Sinhala fake news on Social media was proposed by Wijayarathna and Jayalal [269] where the text content of the post is checked against credible sources and the authenticity of the user account posting the relevant post is evaluated by a rule-based points allocation schema. Wijayarathna and Jayalal [270] collected a set of 120 fake news tweets and 250 non-fake news tweets which they then converted to vectors by taking the fasttext vector for words and averaging. The vector representations of the tweets were then compared to predict whether the news containing it is real or fake. The study by Udurawana et al. [271] proposes to use an accuracy score (obtained by analysing the news content) and a credibility score (obtained by using a scoring mechanism) to detect fake news in Sinhala text. They also incorporate a module that classifies on the basis of passive aggressiveness. The study by Adihetti and Jayalal [272] used autoencoders to detect Sinhala fake news from social media posts.
3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [273, 274, 275, 276, 277] for Sinhala. For this, Arukgoda et al. [159] have proposed a system named Aruth based on the Lesk Algorithm[278]. An online tool [^94] http://aruth.herokuapp.com/, an API [^95] https://bit.ly/3sJEYbS of the algorithm, and code along with data on github [^96] https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [279] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [280]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [281]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [282]. The study by Wimalasuriya [283], which has the same name as the above work by Welgama [282], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [284]. The study by Rathnayake et al. [285] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [286] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [137] introduced the XL-sum [^97] https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website [^98] https://www.bbc.com/sinhala. The M3LS [^99] https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [138] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website100100100https://www.bbc.com/sinhala. Further, they claim that MT5 [287] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
3.16 Other Semantic Tools
Applications of the semantic layer are more advanced than the ones below it in Figure 1. But even with the obvious lack of resources and tools, a number of attempts have been made on semantic-level applications for the Sinhala Language. The earliest attempt on semantic analysis was done by Herath et al. [288] using their earlier work which dealt with Sinhala morphological analysis [53].
3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [289]Wickramasinghe et al. [290] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)101101101https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [131] has Sinhala data for the Spoken Language Identification (LID) task.
3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [291]Wasala et al. [292] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [293]Weerasinghe et al. [294] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [295]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [296]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [297, 298, 299]Sodimana et al. [300] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [301] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [302] and Mishangi [303] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [304]. The works by Lakmal et al. [305] and Senarathna et al. [306] adapted MaryTTS [307] to synthesize Sinhala speech. The study by Jayawardhana et al. [308] used Deep Voice [309] for Sinhala and English TTS. Gamage et al. [310] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system.  Anuradha and Thelijjagoda [311] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [312].
3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [313] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [314], active learning [315], and speaker adaptation [316]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [317] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [318]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [319]Priyadarshani [320] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [321], genetic algorithms [322], and syllable segmentation method utilizing acoustic envelopes [323]. The method proposed by Gunasekara and Meegama [324] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [325]. The work by Karunanayake et al. [326] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [327] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [328] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [329] and Subspace Gaussian Mixture Model (SGMM) [329] in Sinhala speech recognition. In the follow-up work, Gamage et al. [330] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [331] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [332]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result.  Karunathilaka et al. [333] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [334] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [310] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [335]. They have used the Kaldi speech recognition toolkit [336] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [337], proposed a Sinhala speech recognition system. The study by Azir et al. [338] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [339] is based on Tacotron [340] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [341] used DeepSpeech102102102https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [342] used an earlier version of the CMUSphinx toolkit103103103https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [343] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [344, 345] to automatically generate Sinhala subtitles from Speech.
3.17.3 Speech-to-Speech
Layansan et al. [346] created a speech-to-speech translation system for Sinhala on the Android platform. The system developed by Rajapakshe et al. [347] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end.
3.17.4 Speech-to-Intent
The work by Karunanayake et al. [348] used English phoneme-Based Automatic Speech Recognition (ASR) for intent identification in Sinhala and Tamil. Ignatius and Thayasivam [349] proposed a speaker-invariant speech-to-intent classification model with i-vector based speaker normalization, which was then evaluated on Sinhala, and Tamil speech intent data sets. The later work by Yadav et al. [350] used pre-trained embeddings for Sinhala speech intent classification. Hellarawa and Thayasivam [127] proposes a BiLSTM-based ASR system for intent classification which they have tested on the banking domain Sinhala speech dataset created by Buddhika et al. [126]. For this, they report an accuracy of 98.53%.
3.17.5 Speech classification
The Sinhala speech classification system proposed by Buddhika et al. [209] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [351] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [352] presented a Sinhala speech command classification system which can be used for downstream applications. The work by Welarathna et al. [353] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. The voice assistant system created by Senarathne et al. [354] is capable of handling Sinhala voice commands.
3.17.6 Lip Synchronization
The study by Weerathunga et al. [355] worked on lip synchronization for Sinhala speech where videos of people speaking Sinhala were mapped to a visemes alphabet created by them. Further of this line of study, Wakkumbura et al. [356] came up with Phoneme-Viseme mapping for Sinhala speech that they intended to be used for future applications of robotics.
3.18 Optical Character Recognition Applications
While it is not necessarily a component of the NLP stack shown in Fig 1, which follows the definition by Liddy [25], it is possible to swap out the bottom most phonological layer of the stack in favour of an Optical Character Recognition (OCR) and text rendering layer.
3.18.1 Printed Text
The XTREME-UP104104104https://github.com/google-research/xtreme-up data set created by Ruder et al. [130] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [357] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [358, 359]. They then used hidden Markov model-based optimization on the recognized Sinhala script [360]. Similarly, Hewavitharana et al. [361] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [112], Herath and Medagoda [113] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [362], by Ajward et al. [363], and by Madushanka et al. [364].  Karunanayaka et al. [365] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [366] mainly involving the nearest-neighbor method [367, 368]. A study by Ediriweera [369] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [370] has been extended to be online and made available to use via desktops [371] and hand-held devices [372] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [373]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [374]Premachandra et al. [375] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [376]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [377].  Jayawickrama et al. [378] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [379] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [380]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [102] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [381].  Anuradha et al. [382] used Tesseract 3105105105https://tesseract-ocr.github.io/ [383] for Sinhala OCR. The later study by [384] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [385] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [386], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [304]. A study on Sinhala text extraction from social media images (memes) was conducted by Samarajeewa and Ranathunga [387]. They specifically handled the character-touching issue. The study by Walawage and Ranathunga [388] attempted to devise a feature set to separately identify Sinhala and English text on social media images (memes). The study by de Silva and Liyanage [389] used Convolutional Spiking Neural Networks to extract Sinhala text from YouTube thumbnails.  Chanda et al. [390] proposed a Gaussian kernel SVM based method for word-wise Sinhala, Tamil, and English script identification. The work by Vasantharajan et al. [103] adapted the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents to create a Tamil-Sinhala-English parallel corpus.
3.18.2 Handwritten Text
Fernando et al. [391] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [392] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [393]Nilaweera et al. [394] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [395] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [396]Silva et al. [397] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [398]. The study by Walawage and Ranathunga [399] and its follow up study by Walawage [400] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [401] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [402] and Wasalthilake and Kartheeswaran [403], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [404]Ifhaam and Jayalal [405] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [406]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [407] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [337], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [408] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia.
3.18.3 Ancient Text
Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al. [409]. The work of Peiris [410] also focused on OCR for ancient Sinhala inscriptions. Building upon the architecture proposed by Ruwanmini et al. [411], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al. [412]. The study by Wickramarathna and Ranathunga [413] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings. Heenkenda and Fernando [414] used Inception-v3 [415], VGG-19 [416], and ResNet-50 [417] to classify Sinhala inscriptions to historical time periods.
3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [418]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [419]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [420] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [421] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [119] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conduced by Sheshadri et al. [422] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [423] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [424] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. As mentioned in Section 3.1, the study by [93] raised questions on the quality of the existing Sinhala-English corpora. Sen et al. [425] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair. A study on the viability of using Google Translate for the legal domain English-Sinhala and Sinhala-English translation was conducted by [426]. They used human experts to extensively analyse the end result of the translation with many concrete examples of legal phrases that needed to be translated. The multilingual lexicon for low-resource machine translation dataset GATITOS introduced by Jones et al. [427] does not include Sinhala among their 170 languages set. However, they use Sinhala data from other sources and include several observations for Sinhala in this task. They observe that PanLex [428] and GATITOS improve results in the En-Si direction for smaller models (Transformer Big, 475M) but weaken the results in the Si-En direction. For bigger models (Transformer 1.6B), they do not observe an improvement or degradation of En-Si or Si-En direction.
3.19.1 Sinhala-English Non-NMT
A series of work has been done by a group towards English to Sinhala translation as mentioned in some of the above subsections. This work includes; building a morphological analyzer [65], lexicon databases [68], a transliteration system [69], an evaluation model [74], a computational model of grammar [19], and a multi-agent solution [80]. After working on human-assisted machine translation [70],  Hettige and Karunananda [73, 75] have attempted to establish a theoretical basics for English to Sinhala machine translation. A very simplistic web based translator was proposed [71, 72]. The same group have worked on a Sinhala ontology generator for the purpose of machine translation [79] and a phrase level translator [81] based on the previous work on a multi-agent system for translation [78]. Further, an application of the English to Sinhala translator on the use case of selected text for reading was implemented [76]. They later continued their work on multi-agent English to Sinhala translation with the AGR organizational model [82].
Another group independently attempted English-to-Sinhala machine translation [429] with a statistical approach [430]Wijerathna et al. [431] and De Silva et al. [432] have proposed simple rule based translators. An example-based method applied on the English-Sinhala sentence aligned government domain corpus was proposed by Silva and Weerasinghe [433]. A translator based on a look-up system was proposed by Vidanaralage et al. [434]. In a preprint, Joseph et al. [435] proposes an evolutionary algorithm for Sinhala to English translation with a basis of Point-wise Mutual Information (PMI) and claims that the code will be shared once the paper is accepted. However, they do not report any quantitative results to be compared and the reported qualitative results are also superficial. Pushpananda et al. [436] utilized statistical machine translation to translate between Sinhala and Tamil. Fernando et al. [437] tries to solve the Out of vocabulary (OOV) problem for Sinhala in the context of Sinhala-English-Tamil statistical machine translation. The approach proposed by Rajitha et al. [438] uses statistical machine translation and transliteration to align Sinhala and English documents.
3.19.2 Sinhala-English NMT
Fonseka et al. [439] used Byte Pair Encoding (BPE) for English to Sinhala neural machine translation. As another solution to the OOV problem, an analysis of subword techniques to improve English to Sinhala Neural Machine Translation (NMT) was conducted by Naranpanawa et al. [440]. A data augmentation method to expand bilingual lexicon terms based on case markers for the purpose of solving the OOV problem in the domain of NMT was proposed by Fernando et al. [441] which they later extended further [442]Epaliyana et al. [443] proposed iterative filtering and data selection be used to improve Sinhala-English NMT. Perera et al. [444] used English Part-of-Speech (PoS) tags to improve English to Sinhala NMT. Lin et al. [445] used a model based on fairseq [446] to improve machine translation between English and Sinhala. The second half of the study by Arafath [334] dealt with translating Sinhala speech to other languages. Kugathasan and Sumathipala [447] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [448]. Nguyen et al. [449] introduced a new fine-tuning objective LAgSwAV (Language-Agnostic Constraint for SwAV loss), using which they obtained 5.4 BLEU for English-Sinhala. Later, this method was further discussed by Nguyen [450]. Attigala and Weerasinghe [451] conducted an analysis on the effectiveness of ChatGPT [452] in translating Sinhala songs to English.
3.19.3 Singlish to Sinhala (Transliteration)
The XTREME-UP106106106https://github.com/google-research/xtreme-up data set created by Ruder et al. [130] contains a Sinhala data set for the transliteration task. The early work by Goonetilleke et al. [453] attempted Sinhala transliteration through the Latin alphabet. However, their work does not use the word transliteration and instead focuses on the predictive aspect. Priyadarshani et al. [454] used statistical machine learning for transliteration of names between Sinhala, Tamil, and English. A rule-based system on trigrams was proposed by Liwera and Ranathunga [455] for Singlish to Sinhala transliteration of social media text. A Singlish to Sinhala converter which uses an LSTM was proposed by De Silva [456]. A rule-based approach for the same was proposed by de Silva and Ahangama [457]. The study by Nanayakkara et al. [458] introduced an English-to-Sinhala transliteration system. Sumanathilaka et al. [459, 460] proposed a Trie [461] data structure-based algorithm for word suggestion in Sinhala transliterations. An extended analysis of the same work was presented in a later work [462]. Amarasekara et al. [463] proposed a rule-based method supported by N-gram analysis and a corpus dataset to transliterate Singlish tweets to Sinhala. The study by Rajapaksha et al. [464] claims to have created a translation system, however, what they have created is a transliteration system between Singlish and Sinhala. While they also claim to have trained ASR and TTS systems, the paper does not identify them by name or citation. The same is true for the data sets they have used to train the various modules in their pipeline. The studies by Kumaravithana et al. [465] and Jayawardhana et al. [466] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
3.19.4 Singlish and English (Translation via Transliteration)
An LSTM-based sequence-to-sequence model was used by Sandaruwan et al. [467] for Singlish to English NMT task. Nalinka et al. [468] used simple transformers with positional embedding to translate Singlish into English. The study by De Silva et al. [469] claims to have achieved Singlish to English translation by simple stemming.
3.19.5 Between Sinhala and Non-English Languages
Most of the cross-Sinhala and Tamil work has been done in the domain of machine translation. A neural machine translation for Sinhala and Tamil languages was initiated by Tennage et al. [470, 471]. Then they further enhanced it with transliteration and byte pair encoding [472] and used synthetic training data to handle the rare word problem [473]. This project produced Si-Ta [474] a machine translation system of Sinhala and Tamil official documents. In the statistical machine translation front, Farhath et al. [475] worked on integrating bilingual lists. The attempts by Weerasinghe [476] and Sripirakas et al. [477] were also focused on statistical machine translation while Jeyakaran and Weerasinghe [478] attempted a kernel regression method. A yet another attempt was made by Pushpananda et al. [479] which they later extended with some quality improvements [480]. An attempt at real-time direct translation between Sinhala and Tamil was done by Rajpirathap et al. [481]Dilshani et al. [482] have done a study on the linguistic divergence of Sinhala and Tamil languages with respect to machine translation. Mokanarangan [483] claims to have built a named entity translator between Sinhala and Tamil for official government documents. But this work is locked behind an institutional repository wall and thus is not accessible by other researchers. Arukgoda et al. [484] studied the possibility of using deep learning techniques to improve Sinhala-Tamil translation which they further improved later [485]Pramodya et al. [486] compared Transformers, Recurrent Neural Networks, and Statistical Machine Translation (SMT) in the context of Tamil to Sinhala machine translation. The work by Nissanka et al. [487] used monolingual word embedding to improve NMT between Sinhala and Tamil. Thillainathan et al. [488] uses pre-trained mBART [489] models for six directional translations between Sinhala, Tamil, and English. Yashothara and Uthayasanker [490] discussed the use of the Hierarchical Phrase-Based Model for Tamil to Sinhala and Sinhala to Tamil translations. Pramodya [491] presented a comparison of SMT [436] and NMT models for Sinhala-Tamil translation. For NMT they have used base transformer, turned transformer, and mT5 [287].
There have been attempts to link Sinhala NLP with Japanese by Herath et al. [22, 84, 83]Thelijjagoda et al. [85]Thelijjagoda et al. [492], and Kanduboda [9]. Jayasinghe [493] discusses the importance of translation between Sinhala and the liturgical language of Buddhism, Pali [494, 495, 496]. [497] attempted to use a dictionary-based machine-translation method for this task. The study by Anuradha and Thelijjagoda [311] uses machine translation on the unique application of converting Sinhala and English Braille documents, which they have run OCR on, into voice.
3.20 Spelling and Grammar
The open-source data driven approach proposed by Wasala et al. [498, 499] claims to be able to check and correct spelling errors in Sinhala. The approach by Jayalatharachchi et al. [500] attempts to obtain synergy between two algorithms for the same purpose. These efforts [498, 500] were then extended by Subhagya et al. [501]. A rule-based Sinhala spell checker named SinSpell based on Hunspell107107107http://hunspell.github.io/ was introduced by Liyanapathirana et al. [502]. They have also made the tool available108108108http://nlp-tools.uom.lk/sinspell/ online for use. The study by Sithamparanathan and Uthayasanker [503] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil. Sonnadara et al. [116] created a benchmark data set for Sinhala spell correction along with a neural model. A multi agent-based spell checker, named LaSi Spell for Sinhala spell checking was introduced by Samarawickrama et al. [504]. The study by Udagedara et al. [505] specifically solved the problem of spell-checking Sri Lankan names and addresses. A system named Erroff was proposed by Sudesh et al. [506] to correct real-word errors in Sinhala text.
A model for detecting grammatical mistakes in Sinhala was developed by Pabasara and Jayalal [507]. They followed this up with a grammatical error detection and correction model [508]Gunasekara et al. [509] used annotation projection for semantic role labelling for Sinhala. A Sinhala grammar checker based on Hidden Markov models was developed by Fernando and Arudchelvam [510]. Widyaratna [511] used a sequence-to-sequence model with attention, which is generally used for translation tasks, to translate sentences with common grammatical errors to their corrected counterparts. The work by Jayasuriya et al. [512] used a rule-based approach along with Google Translation to correct grammar is Sinhala text. A rule-based system to convert Sinhala sentences from active voice to passive voice while correcting grammatical errors was proposed by Ilukkumbura and Rupasinghe [513]. The study by Goonawardena et al. [514] a rule-based system to spell-check Sinhala text as well as detect and correct grammatical errors. Navoda et al. [515] also claim to have created an automated tool to check Sinhala spelling and grammar.
3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [67]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [516] where they used, the fasttext model trained by Facebook [109, 110, 111], on a RASA109109109https://rasa.com/ [517] chat bot. A Sinhala chat bot for train information was proposed by Harshani [518]. Similarly, the tool proposed by Chandrasena et al. [519] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [520] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [521] used off the shelf RASA NLU Engine [517] and Microsoft Bot Network [522] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [523]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [524] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [347] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
3.22 News and/or Social Media Recommendation
A trending topic detection model for Sinhala tweets using simple clustering and ranking algorithms was proposed by Jayasekara and Ahangama [525]. Sandamini et al. [526] proposed a post recommendation system, which supports Singlish, for social media. Tennakoon and Gamlath [527] proposes a hybrid system which uses skip-gram and collaborative Filtering on Multi-Layer Perceptron for recommending categorized Sinhala news articles. Tennakoon et al. [528] then extends the the system to also be able identify grey sheep users while preforming the aforementioned hybrid recommendation using LDA [225] and SVM. Following the above work, a news aggregator with news categorization, comment filtering, and two types of recommendation systems was proposed by Malsha et al. [529].  Madhushika et al. [530] analysed Twitter trending topics to understand how Sinhala Twitter data affects news dissemination on mass media. They proposed calculating a news value to a tweet which can be utilized to sort tweets by their news-worthiness in order to give better recommendations.
3.23 Sinhala Sign Language
Meyler [531] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language. Further, [532] claim that SSL was derived from British Sign Language (BSL). However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL. In this section, we will discuss works that have used either term. But when relevant we will point out the works that claimed to work on Sri Lankan Sign Language as opposed to Sinhala Sign Language. As far as we can observe, there does not seem to be a technical or functional difference between the two.
In the domain of Sinhala sign language, Liyanaarachchi et al. [533] claims to have created a signing dataset for the Sinhala sign language however only the abstract of their work can be publicly accessible. DISSANAYAKE et al. [534] also claims to have created a database of gestures that are included in Sinhala sign language. Charuka et al. [139] have published a dynamic word-level Sinhala Sign Language video dataset of 50 classes. The dataset contains 1110 videos.
Priyankara et al. [535] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [536] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [537], the approach by Dilakshan and Priyadarshana [538], as well as the approach by Peiris [539] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [540] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [541] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [542]. The study by Charuka et al. [139] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [543] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [544] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. Strides have been made in the domains of computer interpreting for written Sinhala [545] and animation of finger-spelled words and number signs [546].
Kumar et al. [547] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. Both the approaches by Perera and Kumarika [548] and Herath and Ishanka [549] integrated MediaPipe [550] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [548] only used LSTM while Herath and Ishanka [549] used LSTM, CNN, and CNN-LSTM. Rishan et al. [532] uses Leap Motion Controller110110110https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [551] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
3.24 Sinhala Braille
The work by de Silva and Liyanage [389] uses KNN, SVM, and a simple ANN system to recognize Sinhala Braille text. Vithanage [552] also claims to have created a conversion engine to easily convert the Braille text into the corresponding Sinhala text. The study by Madubashana [553], the study by Ariyarathna et al. [554] as well as the study by Weerasinghe [555] also focused on creating an automated Braille to Sinhala recognition system.  Anuradha and Thelijjagoda [311] uses OCR on Sinhala and English Braille documents on which they then run a machine translation system in order to convert them into voice.
3.25 Plagiarism Detection
An extremely simple plagiarism detection tool which only uses n-grams of simply tokenized text was proposed by Basnayake et al. [556]. Another simple plagiarism detection tool that uses synonymy and Hyponymy-Hypernymy (which they call Generalization in the paper) was attempted by Rajamanthri and Thelijjagoda [557]. They later extended this work [558] to propose a more advanced plagiarism detection tool which uses Internet resources. KasthuriArachchi and Charles [559] proposed using Word2Vec vector cosign similarity to detect plagiarism. A multi-document Sinhala similarity detection tool to detect plagiarism was proposed byPiyarathna [560]Punchihewa et al. [561] developed a character-level model which can identify the author for Sinhala text in student answers.
3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [562] and later by Smith [563] using an XGB classifier and a CRF model building on their previous work [564], which analysed such data. Shanmugalingam and Sumathipala [565] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [448]. They later used it for NMT in Sinhala-English Code-Mixed text [447]. As discussed in Section 3.10Rathnayake et al. [216] used adapter-based [217, 218, 219, 220, 221, 222] fine-tuning on XLM-R [223], for classifying code-mixed and code-switched Sinhala text. Hettigoda [226] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [246] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [228] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [523] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [388] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [566] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook.
3.27 Miscellaneous Applications
In this section, we discuss NLP tools and research which are either hard to categorize under above sections or reasonably involve multiples of them.
Dissanayake and Hettige [567] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [568] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha. On the other hand, Amarasinghe and Ranathunga [569] used an ontology-based approch to generate Sinhala essay questions. The work by Liyanage and Ranathunga [570, 571] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages. The follow-up work by Niyarepola et al. [572] used pre-trained mBART [489] and mT5 [287, 573] models for the same task and showed better results. The study by Kao and Ilmini [574] used LSTM to generate Sinhala lyrics. Bandara et al. [575] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala. However, they did not train a GTP-2 model in Sinhala. Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API.
Fernando [576] proposed a method for inexact matching of Sinhala proper names. A study on determining canonical word order of colloquial Sinhala sentences using priority information was conducted by Kanduboda and Tamaoka [577] which they later extended [578, 579, 580]. A cross-language information retrieval system where Sinhala search queries are converted to English search queries by mapping Sinhala word embeddings to English word embeddings was proposed by Hisan et al. [581].
Sandathara et al. [582] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [583] assists children to learn to speak without stuttering. Nethmi et al. [584] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. Ranasinghe et al. [585] trained a Sinhala image caption generator model on Google translated MS COCO [586] dataset and tested it on Google translated Flickr8K [587] and MS COCO datasets.
Rajitha et al. [438] has proposed a statistical machine translation and transliteration method to align Sinhala and English documents. Rajitha et al. [588] used the data set that they introduced in their previous work [94] to prove that task specific supervised distance learning metrics outperform their unsupervised counterparts, for document alignment. Fernando et al. [125] have used pre-trained multilingual language models (PMLMs) to improve document and sentence alignment between Sinhala–English, Tamil–English, and Sinhala–Tamil language pairs.
Kumari and Hettiarachchi [589] proposed an algorithm for Sinhala topic modelling based on LDA [225] and RAKE [590]Batawalaarachchi [591] proposed two methods using statistical features to select words to be included in Sinhala document titles. Arambewela et al. [592] proposed a Sinhala writing assistant tool utilising CNNS.  Jayaweera et al. [593] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [594]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [98] has published two Sinhala BERT models: SinBERT-large111111111https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small112112112https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo113113113https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [595] on the OSCAR [140, 141] dataset. Wickramasinghe and de Silva [134] used Sinhala-English dictionary data [133] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub114114114https://bit.ly/3t3SKu7. Liyanage et al. [596] have created a small Sinhala treebank of 100 sentences. Burchell et al. [132].show that they obtain better results for the LID task for most of the languages compared to NLLB [123]. However, for Sinhala the results by Burchell et al. [132] are the same as that of NLLB. Minixhofer et al. [597] have used Sinhala from OPUS100 [120] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [598] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [573] and mT5 [287, 573] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [599]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [600]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [601].
4 Primary Sources
Even though the main objective of this survey is to cover NLP tools and research, we noticed that much of these NLP tools and research depend on primary sources of Sinhala language such as printed books in the role of knowledge sources and ground truth. Therefore, for the benefit of other researchers who venture into Sinhala NLP, we decided to add a short introduction to the available primary sources of Sinhala language used by their peers. We note that the body of work by a single scholar, Disanayake [2, 602, 603, 604, 605, 606, 607, 608, 40, 609, 610, 611, 612], is quite prominent in the case of being used for NLP applications. For formal introduction of the language, the books by Disanayake [2] and Perera [3] are commonly used. In cases which deal with the Sinhala alphabet, the introduction by Indrasena [613] and by Disanayake [602, 603] have been used. An analysis of modern Sinhala linguistics has been done by Jayathilake [614] and by Pallatthara and Weihene [41]. The early study by Henadeerage [615] covers a number of topics on the Sinhala language such as grammatical relations, argument structure, phrase structure and focus constructions.
As we discussed in Section 3.3, a number of printed Sinhala dictionaries exist, Malalasekera [142] being the most prominent English-Sinahala dictionary among them. In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [148, 144, 147], Sinhala-Sinhala [143, 146], and English-Sinhala-Tamil [145].
A number of NLP applications have utilized first sources intended to teach children [616, 617, 618, 619, 620] or foreigners [621, 186, 622]115115115Note that [622] is an extension of [186].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
For applications where a rule based approach for Sinhala spelling correction is utilized, the books by Disanayake [609, 610], by Koparahewa [623], and by Gair and Karunatillake [624] are used to provide a basis. A number of NLP applications which handle spoken Sinhala in the capacity of phonological layer (Section 3.17) applications or otherwise, make note of the fact that spoken Sinhala is considerably distinguishable from written Sinhala, as such, they refer primary sources which explicitly deal with spoken Sinhala [621, 608, 40, 625, 626, 618, 612].
Primary sources used in NLP application for Sinhala grammar are varied. A number of them provide overviews of the entirety of Sinhala grammar [627, 41, 20, 628, 629, 630, 631, 44, 632, 633, 634, 635, 636, 611]. There are specific primary sources focusing on verbs [637, 619, 607], nouns [620, 606], prepositions [618], compounds [605], derivation [604], case system [638], and sentence structure [45] of the Sinhala language. The book by Rajapaksha [639] is commonly used in NLP applications as a guide for word tagging and punctuation mark handling. NLP studies that tackle the hard problem of handling questions expressed in Sinhala often refer to the book by Kariyakarawana [640]Kekulawala [641] has aptly discussed the much controversial topic of the situation of future tense of Sinhala.
5 Conclusion
At this point, a reader might think, there seems to be a significant number of implementations of NLP for Sinhala. Therefore, how can one justify listing Sinhala as a resource poor language? The important point which is missing in that assumption is that in the cases of almost all of the above listed implementations and findings, the only thing that is publicly available for a researcher is a set of research papers. The corpora, tools, algorithm, and anything else that were discovered through these research are either locked away as properties of individual research groups or worse lost to the time with crashed ancient servers, lost hard drives, and expired web hosts. This reason and probably academic/research rivalry have caused these separate research groups not to cite or build upon the works of each-other. In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig. 7 in Appendix B). This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP work-flow. In their analysis of low resourced languages, Ranathunga and de Silva [642] observed that only 11.43% of Sinhala NLP papers have released the relevant data sets. Further, according to them, code being released sits at 9.71% while tools being released sits at 5.71%. While the 11.43% figure of data release may induce a feeling of availability, Ranathunga and de Silva [642] further observes that it is only 1.14% that has been released in public repositories. Research publications promising access to data and code only to be found lacking later is a common academic shortfall according to Gabelica et al. [36] however, given that Sinhala NLP is already having minimal work done as a whole compared to some other languages, we simply cannot afford to lose any of the generated data or code. In conclusion, what can be said is that, even though there are islands of implementations done for Sinhala NLP, they are of very small scale and/or are usually not readily accessible for further use and research by other researchers. Thus, so far, sadly, Sinhala stays a resource poor language.
Acknowledgments
The authors would like to thank Romain Egele for checking the examples we have provided in French for their accuracy. Similarly, the authors would also like to thank Shravan Kale for checking the examples we have provided in Hindi for their accuracy.
Appendix A Evolutionary Eras of the Sinhala Script
Figure 3: Evolutionary eras of the Sinhala alphabet as proposed by Lankage [643, 644], Mudiyanse [645], Thennakoon [646], Karunarathne [647], Deraniyagala [648], and Nandasara and Mikami [64]
The evolutionary eras of the Sinhala script have raised much scholarly debate. We have shown some prominent categorizations in Fig 3. Other than Mudiyanse [645], scholars generally agree with the fact that the oldest records start at around 300 BCE. Other than Deraniyagala [648], all others agree that an era border should exist at 1AD. However, this is a peculiar observation. Lankage [643, 644], Mudiyanse [645], Karunarathne [647], and Nandasara and Mikami [64] place an era border at 300 CE while Lankage [643, 644], Mudiyanse [645], Thennakoon [646], and Nandasara and Mikami [64] place an era border at 500 CE. The next most common border placement is at 800 CE by Lankage [643, 644], Mudiyanse [645], Thennakoon [646], and Karunarathne [647]. The last relatively common border placement is agreed between Lankage [643, 644], Mudiyanse [645], Thennakoon [646], and Nandasara and Mikami [64]. Given that the widest coverage is provided by Nandasara and Mikami [64], we have used their era definitions and examples in Fig 4 and Fig 5. The sources from which Nandasara and Mikami [64] have extracted the ancient script are given by the relevant superscripts as follows:
300 BCE - 1 CE
  1. ${}^{{1}}$
    Periyankulama (207-197 BCE)
  2. ${}^{{2}}$
    Mihintale (207-197 BCE)
  3. ${}^{{3}}$
    Situlpawwa (161-137 BCE)
  4. ${}^{{4}}$
    Korawakgala (77-63 BCE)
  5. ${}^{{5}}$
    Ritigala Weweltanne (22-7 BCE)
  6. ${}^{{6}}$
    Yatahalena Vihara (22-7 BCE)
  7. ${}^{{7}}$
    Gallena Vihara (22-7 BCE)
  8. ${}^{{8}}$
    Nuwaragala (22-7 BCE)
  9. ${}^{{9}}$
    Ritigala Andiayakanna (22-7 BCE)
  10. ${}^{{10}}$
    Boowattegala (22-7 BCE)
  11. ${}^{{11}}$
    Rajagala (44-22 BCE)
1 CE - 300 CE
  1. ${}^{{12}}$
    Anuradhapura (1-7 CE)
  2. ${}^{{13}}$
    Situlpawwa (1-7 CE)
  3. ${}^{{14}}$
    Maharatmale (7-18 CE)
  4. ${}^{{15}}$
    Wallipuram (67-111 CE)
  5. ${}^{{16}}$
    Viharagala (60-67 CE)
  6. ${}^{{17}}$
    Pahala Kainattama (60-67 CE)
300 CE - 500 CE
  1. ${}^{{18}}$
    Tonigala (301-328 CE)
  2. ${}^{{19}}$
    Ruwanweliseya (337-365 CE)
  3. ${}^{{20}}$
    Thissamaharama (406-428 CE)
  4. ${}^{{21}}$
    Anuradhapura (437-452 CE)
500 CE - 700 CE
  1. ${}^{{22}}$
    Kandanadu (517-518 CE)
  2. ${}^{{23}}$
    Dhakshinathupa (639-650 CE)
  3. ${}^{{24}}$
    Baron Paviliyan (639-650 CE)
  4. ${}^{{25}}$
    Kuchchaweli (639-650 CE)
  5. ${}^{{26}}$
    Murutawa (639-650 CE)
700 CE - 1000 CE
  1. ${}^{{27}}$
    Thiriyaya (733-771 CE)
  2. ${}^{{28}}$
    Viyaulpotha (853-887 CE)
  3. ${}^{{29}}$
    Dorabewila (915-923 CE)
  4. ${}^{{30}}$
    Baddulla (946-954 CE)
  5. ${}^{{31}}$
    Polonnaruwa (982-1029 CE)
  6. ${}^{{32}}$
    Indikatuseya (982-1029 CE)
Figure 4: The evolution of the Sinhala script under the era classification proposed by Nandasara and Mikami [64] - Part 1.
Note: The IPA for (*) anusvāraya [649] and (**) visargaya [650] are approximations as there is no consensus on their representation. 1200 AD and 1500 AD from inscriptions and pillars [64]. 1737 from the first printed Sinhala book collected by Nandasara [651]. 1876 from “Alfabete des gesammten Erdkreises” (Alphabets of the entire world[652] reported as CINGALESISCH [651]. 1891 from the alphabet shown by Gunasekara [186]. 1996 from the character set Sarasavi developed by S T Nandasara [651]. Nandasara and Mikami [64] report that the gray cells are not included due to being possible to be produced using consonant modifiers. 1998 from the Iskoola Pota UNICODE font by Microsoft [653].
Figure 5: The evolution of the Sinhala script under the era classification proposed by Nandasara and Mikami [64] - Part 2
Note: 1200 AD and 1500 AD from inscriptions and pillars [64]. 1737 from the first printed Sinhala book collected by Nandasara [651]. 1876 from “Alfabete des gesammten Erdkreises” (Alphabets of the entire world[652] reported as CINGALESISCH [651]. 1891 from the alphabet shown by Gunasekara [186]. 1996 from the character set Sarasavi developed by S T Nandasara [651]. Nandasara and Mikami [64] report that the gray cells are not included due to being possible to be produced using consonant modifiers. 1998 from the Iskoola Pota UNICODE font by Microsoft [653].
Appendix B Author Meta Analysis
Figure 6: Co-author graph of the most prolific researchers in the Sinhala NLP domain (Selected at the threshold of at least 3 publications)
Figure 7: The Probability of research from an institution citing that of other institutions. Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext117117117https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up. With all those limitations in mind, we still can make a few Interesting observations. The institute with the highest number of publications, UCSC - University of Colombo seems to be getting the most citations from most sources. However, they themselves seem to almost exclusively cite their own papers (0.7543). The only exception is the smaller number of works they cite (0.2114) from the Department of CSE - University of Moratuwa, the institute with the second highest number of publications. This leaves them with a 0.0343 probability of citing anyone else. Comparatively, Department of CSE - University of Moratuwa seems to be more egalitarian in citing. They have a lower self-citation probability (0.5251) and a higher probability of citing UCSC - University of Colombo (0.3513). This results in a probability of citing others at 0.1236. Both the Johns Hopkins University and University of Edinburgh prefer to cite Google instead of work from Sri Lanka. Faculty of IT - University of Moratuwa prefers to cite UCSC - University of Colombo (0.2339) rather than Department of CSE - University of Moratuwa (0.0968) which is from the same parent institution.
Figure 8: Authors with at least 10 papers in the Sinhala NLP domain, mapped to their research interests denoted by the subsection titles of the Section 3 of this paper.

### v22

@@
 However, due to various reasons, these attempts seem to lack coordination and awareness of each other.
-The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize contributions of their peers.
-As such, we shall be uploading this paper to arXiv and perpetually update it periodically to reflect the advances made in the field.
+The objective of this paper is to fill that gap of a comprehensive literature survey of the publicly available Sinhala natural language tools and research so that the researchers working in this field can better utilize the contributions of their peers.
+As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field.
+This manuscript is at version $5.4.0$.
 Index Terms:
@@
 When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7].
-To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that Sinhala language belongs to the same the Indo-European language tree [8, 9].
+A strong correlation between the island of Sri Lanka and the usage of Sinhala has been observed [8] [^2] Dunn and Edwards-Brown [8] in their work to augment Language Identification (LID) task with geography observe that for Sinhala LID does not benefit from the geography model given that the to models already has an alignment of 99%..
+This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala.
+To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10].
 However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch.
@@
 By extension, this makes Sinhala Script a member of the Aramaic family of scripts [18, 19].
-Thus by inheritance, Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [19].
+Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [20].
 We show the evolutionary eras of the Sinhala script in Appendix A.
-It should be noted that the modern Sinhala language have loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [20].
-Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [21, 22]), modern natural language processing tools for the Sinhala language are scarce [23].
+It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [21].
+Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [22, 23]), modern natural language processing tools for the Sinhala language are scarce [24, 25].
+Further, Sinhala is a diglossic language [26] which has two variations: the literary form and the spoken form.
 Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages.
@@
 Some forms of anaphora (coreference) resolution [31, 32, 33, 34, 35] fall into this application.
-Figure 1: NLP layers and tasks [24]
+Figure 1: Traditional NLP layers and tasks (Derived from the work of Wijeratne et al. [27] and extended to encapsulate a wider range by this work.)
 On the other hand, Wimalasuriya and Dou [29] categorize NLP tools and research by utility.
@@
 However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [37, 38] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available.
-We also discuss how the non-compliance of policies to put data and code online [36], after the research is concluded and the paper is published, have negatively impacted the growth and sustainability of Sinhala NLP.
+We also discuss how the non-compliance of policies to put data and code online [39], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP.
 Hence, it is our hope that this work will help future researchers who are engaged in Sinhala NLP research to conduct their literature surveys efficiently and comprehensively.
 For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io.
-This manuscript is at version $5.3.0$.
+This manuscript is at version $5.4.0$.
 The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
 Figure 6 in Appendix B shows the most prolific researchers in the domain of Sinhala NLP.
-The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that researcher has authored.
-The edges between two researchers are labeled with the number of Sinhala NLP papers the relevant pair of researchers have co-authored.
+The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that the researcher has authored.
+The edges between the two researchers are labelled with the number of Sinhala NLP papers the relevant pair of researchers have co-authored.
 When selecting authors, we have applied a threshold of 3 Sinhala NLP publications.
-Given that the objective of the visualization is to portray the co-operation between researchers, we have also added the strongest edge that connects each researcher to the rest of the researchers in the graph.
-The few isolated nodes are researchers that have authored at least 3 Sinhala NLP publications but do not have any coauthored papers with anyone else on the graph.
+Given that the objective of the visualization is to portray the cooperation between researchers, we have also added the strongest edge that connects each researcher to the rest of the researchers in the graph.
+The few isolated nodes are researchers who have authored at least 3 Sinhala NLP publications but do not have any coauthored papers with anyone else on the graph.
 We have also added labels to clusters in cases where all or the majority of researchers in those clusters have the same affiliation.
@@
 Figure 7 in Appendix B shows the probability of studies from the institutions to which the most prolific authors from Figure 6 are affiliated citing institutions of the same set.
-We observe both interesting and disturbing trends which we discuss in the figure caption
+We observe both interesting and disturbing trends which we discuss in the figure caption.
 Figure 8 in Appendix B shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper.
@@
 [99] [^22] A later further overview on the corpus is also available [100].Dhananjaya et al.
-[98] created sin-cc-15M corpus2222223https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus.
+[101] created sin-cc-15M corpus2323233https://tinyurl.com/42un7a9y, which they claim, at the time of their publication to be the largest monolingual Sinhala corpus.
 A text corpus collected from the Sihala blog Kalaya [^24] http://www1.kalaya.org/ along with the relevant code is available on Github [^25] https://bit.ly/KalayaCorpus.
+A Sinhala NEWS corpus of more than 500,000 articles collected from various online free news websites was released [^26] https://github.com/Sinhala-NLP/NSINA by Hettiarachchi et al.
+[102].
 As for Sinhala-Tamil corpora, Hameed et al.
@@
 Zhang et al.
-[120] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100404040https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT).
+[124] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100424242https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT).
 Thier code is also available on GitHub [^43] https://github.com/bzhangGo/zero.
@@
 Kudugunta et al.
-[136] released MADLAD-400727272https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages.
+[140] released MADLAD-400747474https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages.
 For Sinhala, they report that they have 788K noisy docs, 349.2K clean docs, 22.1M noisy sentences, 16M clean sentences, 3.4B noisy tokens, and 1.9B clean tokens.
@@
 The OSCAR [^78] https://oscar-project.org/ dataset [144, 145] contains web mined Sinhala data consisting of 172,755,385 words in 301,066 documents.
+The Dakshina dataset [^79] https://github.com/google-research-datasets/dakshina by [146] contains Native Sinhala script data (200k train and 28k validation sentences) along with Romanized (Singlish) data (10k sentences) from Wikipedia.
+It also contains a Romanization Lexicon (25k train and 5k test).
+Ranathunga et al.
+[147] have publicly released [^80] https://github.com/nlpcuom/quality-matters their data, code, and models for Sinhala-English translation.
+Very importantly, this includes a human-cleaned portion of the NLLB dataset [127] in En-Si [^81] https://huggingface.co/datasets/NLPC-UOM/nllb-top25k-ensi-cleaned.
+Singh et al.
+[148] introduced the Aya Data set [^82] https://huggingface.co/datasets/CohereForAI/aya_dataset, a human-curated multilingual instruction fine-tuning (IFT) data set for large language models (LLM).
+They also introduced Aya Collection [^83] https://huggingface.co/datasets/CohereForAI/aya_collection, 44 instruction-style datasets that were created by transforming existing NLP datasets into pairs of prompt and completion and then using NMT on them to obtain data in multiple languages.
+Both of these data sets include Sinhala.
+Further, they also released Aya Evaluation suite [^84] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality.
 3.3 Dictionaries
@@
 [101].
-Both their code [^84] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus8585853https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^86] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^87] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^88] https://bit.ly/3PARuUS) are publicly available.
+Both their code [^92] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9393933https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^94] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^95] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^96] https://bit.ly/3PARuUS) are publicly available.
 The study by Wijayarathna and Jayalal [234] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter.
@@
 The dataset contains 318,308 Sinhala tweets annotated with emotions.
+The study by Dhananjaya et al.
+[254] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala.
 3.12 Hate Speech Detection
@@
 [122] analyses Sinhala hate speech propagation on Twitter.
-They have also published a dataset [^92] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
+They have also published a dataset100100100https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
 Rajapaksha et al.
@@
 [271] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala.
+The study by Ehelepola [272] analyzed Sinhala text from social media and e-commerce sites for hate speech.
+The work by Dikwatta et al.
+[273] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes).
+Wickramaarachchi et al.
+[274] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech.
 After the release of the SOLD dataset by Ranasinghe et al.
@@
 Fernando and Deng [275] used traditional machine learning algorithms to predict Sinhala hate speech.
-Ojo et al.
-[265] used mBERT to identify hate and offensive content in Sinhala under HASOC 2023, Task 1A [^93] https://hasocfire.github.io/hasoc/2023/task1.html which included the SOLD dataset.
 Ranasinghe and Zampieri [276] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi.
 The work by Bestgen [277] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali.
-The study by Narayan et al.
+Ranasinghe et al.
+[278] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task101101101https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
+In completing the task, the study by Narayan et al.
 [279] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection.
 They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%.
+For the same task, Ojo et al.
+[280] used mBERT to identify hate and offensive content in Sinhala.
 3.13 Fake News Detection
@@
 The study by Adihetti and Jayalal [284] used autoencoders to detect Sinhala fake news from social media posts.
+The study by Wickramaarachchi et al.
+[274] detected fake content in Sinhala youtube videos by comparing their title and description to the audio using LSTM on BART.
 3.14 Word Sense Disambiguation
@@
 [166] have proposed a system named Aruth based on the Lesk Algorithm[290].
-An online tool [^94] http://aruth.herokuapp.com/, an API [^95] https://bit.ly/3sJEYbS of the algorithm, and code along with data on github [^96] https://github.com/jseanm1/aruthSWSD are available.
+An online tool102102102http://aruth.herokuapp.com/, an API103103103https://bit.ly/3sJEYbS of the algorithm, and code along with data on github104104104https://github.com/jseanm1/aruthSWSD are available.
 For the same task, Marasinghe et al.
@@
 Hasan et al.
-[137] introduced the XL-sum [^97] https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website [^98] https://www.bbc.com/sinhala.
-The M3LS [^99] https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
-[138] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website100100100https://www.bbc.com/sinhala.
+[141] introduced the XL-sum105105105https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website106106106https://www.bbc.com/sinhala.
+The M3LS107107107https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
+[142] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website108108108https://www.bbc.com/sinhala.
 Further, they claim that MT5 [299] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
@@
 [302] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems.
-The Massively Multilingual Speech (MMS)101101101https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
+The Massively Multilingual Speech (MMS)109109109https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
 [135] has Sinhala data for the Spoken Language Identification (LID) task.
+The meta-study by Al-Fraihat et al.
+[303] compared the status of Sinhala speech recognition research against 17 other languages.
 3.17.1 Text-to-Speech
@@
 [323] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system.
+The study by Madhusha et al.
+[324] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students.
 Anuradha and Thelijjagoda [325] proposed a machine translation system to convert Sinhala and English Braille documents into voice.
@@
 TacoSi introduced by Arachchige and Weerasinghe [353] is based on Tacotron [354] and has been evaluated with 10 human evaluators to determine its text-to-speech quality.
-Nanayakkara [341] used DeepSpeech102102102https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
+Nanayakkara [355], Nanayakkara and Weerasinghe [356] used DeepSpeech110110110https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
 Gunarathne et al.
-[342] used an earlier version of the CMUSphinx toolkit103103103https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
+[357] used an earlier version of the CMUSphinx toolkit111111111https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
 The later work by Akesh and Meegama [358] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [359, 360] to automatically generate Sinhala subtitles from Speech.
@@
 [367] presented a Sinhala speech command classification system which can be used for downstream applications.
+The voice assistant system created by Senarathne et al.
+[368] is capable of handling Sinhala voice commands.
 The work by Welarathna et al.
 [369] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children.
-The voice assistant system created by Senarathne et al.
-[354] is capable of handling Sinhala voice commands.
+Sundarapperuma [370] created a speech emotion detection system for Sinhala.
 3.17.6 Lip Synchronization
@@
 3.18.1 Printed Text
-The XTREME-UP104104104https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP112112112https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [134] contains a Sinhala data set for the OCR task.
@@
 Anuradha et al.
-[382] used Tesseract 3105105105https://tesseract-ocr.github.io/ [383] for Sinhala OCR.
+[398] used Tesseract 3113113113https://tesseract-ocr.github.io/ [399] for Sinhala OCR.
 The later study by [400] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.
@@
 The study by Withana and Rupasinghe [424] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia.
+The study by eka [425] has used CNN to recognize Sinhala handwritten text.
 3.18.3 Ancient Text
@@
 They used datasets from OPUS [123] for model building and utilized Flores-200 for fine-tuning.
-The NMT for Indic languages study conduced by Sheshadri et al.
+The NMT for Indic languages study conducted by Sheshadri et al.
 [439] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India.
@@
 As mentioned in Section 3.1, the study by [96] raised questions on the quality of the existing Sinhala-English corpora.
+In a follow-up study by Ranathunga et al.
+[147], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil.
+These models sometimes were better performing than a model trained on the full dataset.
+They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering.
+Thus they concluded that using expensive human labour for this task might not be efficient.
 Sen et al.
@@
 Attigala and Weerasinghe [468] conducted an analysis on the effectiveness of ChatGPT [469] in translating Sinhala songs to English.
+Utsa et al.
+[470] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [92] and reported better results than mBART [471].
 3.19.3 Singlish to Sinhala (Transliteration)
-The XTREME-UP106106106https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP114114114https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [134] contains a Sinhala data set for the transliteration task.
@@
 [485] claim to have built Sinhala-English translators but in actuality are just invoking Google API for their translation tasks.
+The work by Kirov et al.
+[486] used the Dakshina [146] dataset and compared the accuracy of Singlish to Sinhala transliteration with vanilla LSTM, vanilla Transformers, mT5 [299], ByT5 [487], and a non-neural finite-state transducer (FST) based on work by Bisani and Ney [488].
+They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole.
+Their code is available on GitHub115115115https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
+Khiu et al.
+[489] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
 3.19.4 Singlish and English (Translation via Transliteration)
@@
 For NMT they have used base transformer, turned transformer, and mT5 [299].
+The work by Su et al.
+[514] which compared 8 Parameter-efficient fine-tuning (PEFT) methods on the Si-Ta language pair using the Government corpus [454] and NLLB [127] reports that Houlsby adapter [224], with a 33.34 SacreBLEU [515] score, to be the best PEFT method.
+However, they also note that the Pfeiffer adapter [226] runs the fastest at 52.59 hours with a reasonable score of 31.24.
+(Comparatively, the Houlsby adapter runs for 78.65 hours).
 There have been attempts to link Sinhala NLP with Japanese by Herath et al.
@@
 [525].
-A rule-based Sinhala spell checker named SinSpell based on Hunspell107107107http://hunspell.github.io/ was introduced by Liyanapathirana et al.
+A rule-based Sinhala spell checker named SinSpell based on Hunspell116116116http://hunspell.github.io/ was introduced by Liyanapathirana et al.
 [526].
-They have also made the tool available108108108http://nlp-tools.uom.lk/sinspell/ online for use.
+They have also made the tool available117117117http://nlp-tools.uom.lk/sinspell/ online for use.
 The study by Sithamparanathan and Uthayasanker [527] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil.
@@
 A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al.
-[516] where they used, the fasttext model trained by Facebook [109, 110, 111], on a RASA109109109https://rasa.com/ [517] chat bot.
+[540] where they used, the fasttext model trained by Facebook [113, 114, 115], on a RASA118118118https://rasa.com/ [541] chat bot.
 A Sinhala chat bot for train information was proposed by Harshani [542].
@@
 [568] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language.
+The study by Fernando et al.
+[569] introduced a mobile app capable of translating Sinhala Sign language to text.
+They further enhance the results of their system with a facial emotion detection system created with CNNs.
+The study by Gamage et al.
+[570] also combines Sinhala Sign language recognition and emotion recognition.
+Krishnananthan et al.
+[571] introduced HANDTALK web-based application designed to teach sign-language.
+As Such it has a component that translates the sign language gestures of the learner into text.
+In their study Gedaragoda et al.
+[572] used traditional machine learning approaches to recognize Sinhala Sign Language.
+A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [573].
+The work by Perera [574] uses CNN for Sri Lanka Sign Language recognition.
 Strides have been made in the domains of computer interpreting for written Sinhala [575] and animation of finger-spelled words and number signs [576].
+The work by Idushan et al.
+[577] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
 Kumar et al.
@@
 Rishan et al.
-[532] uses Leap Motion Controller110110110https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
+[556] uses Leap Motion Controller119119119https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
 Dissanayake et al.
@@
 The study by Fazal and Farook [597] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook.
+The study by uda [598] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection.
+3.27 Language Acquisition and Evaluation
+Sandathara et al.
+[599] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala.
+The learning bot MiMi proposed by Vithana et al.
+[600] assists children to learn to speak without stuttering.
+Nethmi et al.
+[601] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala.
+The mobile application developed by Saranga et al.
+[602] helps individuals learn Sinhala letters with the help of images, colours, and shapes.
+The application also attempts to generate stories in Sinhala.
+The Katha App developed by Wijesooriya et al.
+[603] uses SVM and LSTm to identify phonological disorders from Sinhala speech.
+3.28 Embeddings and Other Language Models
+A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al.
+[604].
+According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding.
+Dhananjaya et al.
+[101] has published two Sinhala BERT models: SinBERT-large120120120https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small121121121https://huggingface.co/NLPC-UOM/SinBERT-small.
+A smaller BERT model named SinhalaBERTo122122122https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
+This BERT model has been created by training vanilla RoBERTa [605] on the OSCAR [144, 145] dataset.
+Üstün et al. [606] introduced Aya, a massively multilingual generative language model123123123https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [148].
+Wickramasinghe and de Silva [138] used Sinhala-English dictionary data [137] for obtaining Sinhala-English embedding alignment.
+Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
+The relevant code and data can be accessed on GitHub124124124https://bit.ly/3t3SKu7.
 3.29 Miscellaneous Applications
@@
 [621].
-Sandathara et al.
-[582] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala.
-The learning bot MiMi proposed by Vithana et al.
-[583] assists children to learn to speak without stuttering.
-Nethmi et al.
-[584] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala.
-Ranasinghe et al.
-[585] trained a Sinhala image caption generator model on Google translated MS COCO [586] dataset and tested it on Google translated Flickr8K [587] and MS COCO datasets.
+Ranasinghe et al. [622] trained a Sinhala image caption generator model on Google translated MS COCO [623] dataset and tested it on Google translated Flickr8K [624] and MS COCO datasets.
 Rajitha et al.
@@
 Rajitha et al.
-[588] used the data set that they introduced in their previous work [94] to prove that task specific supervised distance learning metrics outperform their unsupervised counterparts, for document alignment.
+[625] used the data set that they introduced in their previous work [97] to prove that task-specific supervised distance learning metrics outperform their unsupervised counterparts, for document alignment.
 Fernando et al.
@@
 Kumari and Hettiarachchi [626] proposed an algorithm for Sinhala topic modelling based on LDA [232] and RAKE [627]Batawalaarachchi [628] proposed two methods using statistical features to select words to be included in Sinhala document titles.
+The study by pal [629] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results.
 Arambewela et al.
 [630] proposed a Sinhala writing assistant tool utilising CNNS.
+The study by Bandara et al.
+[631] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.
 Jayaweera et al.
@@
 But they provide no information on where to obtain this corpus.
-A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al.
-[594].
-According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding.
-Dhananjaya et al.
-[98] has published two Sinhala BERT models: SinBERT-large111111111https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small112112112https://huggingface.co/NLPC-UOM/SinBERT-small.
-A smaller BERT model named SinhalaBERTo113113113https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
-This BERT model has been created by training vanilla RoBERTa [595] on the OSCAR [140, 141] dataset.
-Wickramasinghe and de Silva [134] used Sinhala-English dictionary data [133] for obtaining Sinhala-English embedding alignment.
-Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
-The relevant code and data can be accessed on GitHub114114114https://bit.ly/3t3SKu7.
 Liyanage et al.
@@
 Burchell et al.
-[132].show that they obtain better results for the LID task for most of the languages compared to NLLB [123].
+[136] show that they obtain better results for the LID task for most of the languages compared to NLLB [127].
 However, for Sinhala the results by Burchell et al.
 [136] are the same as that of NLLB.
+Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model.
+However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%.
 Minixhofer et al.
@@
 A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [638].
+Senarathne [639] introduced a model to translate Sinhala pseudocode to computer code written in C#.
+Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [597] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook.
+Hewapathirana and Sumanathilaka [640] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
 4 Primary Sources
@@
 In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [155, 151, 154], Sinhala-Sinhala [150, 153], and English-Sinhala-Tamil [152].
-A number of NLP applications have utilized first sources intended to teach children [616, 617, 618, 619, 620] or foreigners [621, 186, 622]115115115Note that [622] is an extension of [186]..
+A number of NLP applications have utilized first sources intended to teach children [655, 656, 657, 658, 659] or foreigners [660, 193, 661]125125125Note that [661] is an extension of [193]..
 This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@
 The corpora, tools, algorithm, and anything else that were discovered through these research are either locked away as properties of individual research groups or worse lost to the time with crashed ancient servers, lost hard drives, and expired web hosts.
-This reason and probably academic/research rivalry have caused these separate research groups not to cite or build upon the works of each-other.
-In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig.
+This reason and probably academic/research rivalry has caused these separate research groups not to cite or build upon the works of each other.
+In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource-rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig.
 7 in Appendix B).
-This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP work-flow.
-In their analysis of low resourced languages, Ranathunga and de Silva [642] observed that only 11.43% of Sinhala NLP papers have released the relevant data sets.
+This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP workflow.
+In their analysis of low-resourced languages, Ranathunga and de Silva [25] observed that only 11.43% of Sinhala NLP papers have released the relevant data sets.
 Further, according to them, code being released sits at 9.71% while tools being released sits at 5.71%.
-While the 11.43% figure of data release may induce a feeling of availability, Ranathunga and de Silva [642] further observes that it is only 1.14% that has been released in public repositories.
+While the 11.43% figure of data release may induce a feeling of availability, Ranathunga and de Silva [25] further observes that it is only 1.14% has been released in public repositories.
 Research publications promising access to data and code only to be found lacking later is a common academic shortfall according to Gabelica et al.
@@
 In conclusion, what can be said is that, even though there are islands of implementations done for Sinhala NLP, they are of very small scale and/or are usually not readily accessible for further use and research by other researchers.
-Thus, so far, sadly, Sinhala stays a resource poor language.
+Thus, so far, sadly, Sinhala stays a resource-poor language.
 Acknowledgments
@@
 300 BCE - 1 CE
-  1. ${}^{{1}}$
+  1. 1
     Periyankulama (207-197 BCE)
-  2. ${}^{{2}}$
+  2. 2
     Mihintale (207-197 BCE)
-  3. ${}^{{3}}$
+  3. 3
     Situlpawwa (161-137 BCE)
-  4. ${}^{{4}}$
+  4. 4
     Korawakgala (77-63 BCE)
-  5. ${}^{{5}}$
+  5. 5
     Ritigala Weweltanne (22-7 BCE)
-  6. ${}^{{6}}$
+  6. 6
     Yatahalena Vihara (22-7 BCE)
-  7. ${}^{{7}}$
+  7. 7
     Gallena Vihara (22-7 BCE)
-  8. ${}^{{8}}$
+  8. 8
     Nuwaragala (22-7 BCE)
-  9. ${}^{{9}}$
+  9. 9
     Ritigala Andiayakanna (22-7 BCE)
-  10. ${}^{{10}}$
+  10. 10
     Boowattegala (22-7 BCE)
-  11. ${}^{{11}}$
+  11. 11
     Rajagala (44-22 BCE)
 1 CE - 300 CE
-  1. ${}^{{12}}$
+  1. 12
     Anuradhapura (1-7 CE)
-  2. ${}^{{13}}$
+  2. 13
     Situlpawwa (1-7 CE)
-  3. ${}^{{14}}$
+  3. 14
     Maharatmale (7-18 CE)
-  4. ${}^{{15}}$
+  4. 15
     Wallipuram (67-111 CE)
-  5. ${}^{{16}}$
+  5. 16
     Viharagala (60-67 CE)
-  6. ${}^{{17}}$
+  6. 17
     Pahala Kainattama (60-67 CE)
 300 CE - 500 CE
-  1. ${}^{{18}}$
+  1. 18
     Tonigala (301-328 CE)
-  2. ${}^{{19}}$
+  2. 19
     Ruwanweliseya (337-365 CE)
-  3. ${}^{{20}}$
+  3. 20
     Thissamaharama (406-428 CE)
-  4. ${}^{{21}}$
+  4. 21
     Anuradhapura (437-452 CE)
 500 CE - 700 CE
-  1. ${}^{{22}}$
+  1. 22
     Kandanadu (517-518 CE)
-  2. ${}^{{23}}$
+  2. 23
     Dhakshinathupa (639-650 CE)
-  3. ${}^{{24}}$
+  3. 24
     Baron Paviliyan (639-650 CE)
-  4. ${}^{{25}}$
+  4. 25
     Kuchchaweli (639-650 CE)
-  5. ${}^{{26}}$
+  5. 26
     Murutawa (639-650 CE)
 700 CE - 1000 CE
-  1. ${}^{{27}}$
+  1. 27
     Thiriyaya (733-771 CE)
-  2. ${}^{{28}}$
+  2. 28
     Viyaulpotha (853-887 CE)
-  3. ${}^{{29}}$
+  3. 29
     Dorabewila (915-923 CE)
-  4. ${}^{{30}}$
+  4. 30
     Baddulla (946-954 CE)
-  5. ${}^{{31}}$
+  5. 31
     Polonnaruwa (982-1029 CE)
-  6. ${}^{{32}}$
+  6. 32
     Indikatuseya (982-1029 CE)
@@
 Figure 7: The Probability of research from an institution citing that of other institutions.
-Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext117117117https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
+Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext127127127https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
 With all those limitations in mind, we still can make a few Interesting observations.

### v23

@@
 As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field.
-This manuscript is at version $5.4.0$.
+This manuscript is at version $5.5.0$.
 Index Terms:
@@
 Further, Sinhala is a diglossic language [26] which has two variations: the literary form and the spoken form.
+The Vedda (Veddah) language is considered a Sinhala creole [27, 28] of a language isolate [29].
 Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages.
@@
 For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io.
-This manuscript is at version $5.4.0$.
+This manuscript is at version $5.5.0$.
 The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@
 Being a multi-modal dataset, M3LS additionally contains the relevant images collected from the same source.
+Hewapathirana et al.
+[146] created the first Sinhala multi-document summarizing dataset, M2DS [^78] https://huggingface.co/datasets/KushanH/m2ds which is consistent of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary.
+They have also released the code and pre-trained models on github [^79] https://github.com/KushanMH/m2ds.
 A dynamic word-level Sinhala Sign Language video dataset of 50 classes has been published by Charuka et al.
@@
 Further, they also released Aya Evaluation suite [^86] https://huggingface.co/datasets/CohereForAI/aya_evaluation_suite to measure multilingual open-ended generation quality.
+Kudugunta et al.
+[153] introduced MADLAD-400878787https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents.
+They have also released the trained checkpoints on GitHub [^88] https://github.com/google-research/google-research/tree/master/madlad_400.
+A multilingual visual Question Answering benchmark which includes Sinhala was released by Romero et al.
+[154] along with data [^89] https://huggingface.co/datasets/afaji/cvqa and a leader board [^90] https://eval.ai/web/challenges/challenge-page/2305/overview.
 3.3 Dictionaries
@@
 Kanduboda [48] have worked on Sinhala differential object markers relevant for parsing.
+A meta-study was conducted by Stephen and Zeman [199] on the Sinhala data from Universal Dependencies (UD) for identifying light verb constructions in Sinhala.
 The first attempt at a Sinhala parser, as mentioned above, was by Hettige and Karunananda [72] where they created prototype Sinhala morphological analyzer and a parser as part of their larger project to build an end-to-end translator system.
@@
 [207] used support vector machines to detect Sinhala named entities of types person, location and, organization.
+Even though the main task reported in the study by [208] is clustering Sinhala news articles, they used NER tools to do so.
+The work by Gurgurov et al.
+[209] compared the performance of LLMs for the Sinhala NER task.
+The highest results are obtained for this task in the Masked Language Modeling objective for Sinhala.
+They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
 3.9 Semantic Similarity
@@
 Another study by Kumari and Haddela [216] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [217] for human interpretability of Sinhala document classification.
-However, they too do not provide access corpus.
+However, they too do not provide access to the corpus.
 Nanayakkara and Ranathunga [218] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.
@@
 [104].
-Both their code [^92] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9393933https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^94] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^95] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^96] https://bit.ly/3PARuUS) are publicly available.
+Both their code [^98] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9999993https://tinyurl.com/42un7a9y, Sinhala News source classification data set100100100https://bit.ly/3PBTyfo, Sinhala News category classification data set101101101https://bit.ly/3AsfK7u, and Sinhala writing style classification data set102102102https://bit.ly/3PARuUS) are publicly available.
 The study by Wijayarathna and Jayalal [243] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter.
@@
 A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [248] based on their prior work [249].
-A word2vec based tool [^97] http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
+A word2vec based tool103103103http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
 A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al.
@@
 [223]Aththanayaka and Naleer [262] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media.
-The ACTSEA dataset [^98] https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
+The ACTSEA dataset104104104https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
 [131].
@@
 [263] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala.
+A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [264].
+However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
+The work by Gurgurov et al.
+[209] compared the performance of LLMs for the Sinhala sentiment analysis task.
+They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
 3.12 Hate Speech Detection
@@
 A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [266].
-A feature model and a data set [^99] https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [121]Sandaruwan et al.
+A feature model and a data set105105105https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [124]Sandaruwan et al.
 [267] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques.
@@
 [125] analyses Sinhala hate speech propagation on Twitter.
-They have also published a dataset100100100https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
+They have also published a dataset106106106https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
 Rajapaksha et al.
@@
 [284] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech.
+The study by Muthuthanthri and Smith [285] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2.
+They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
 After the release of the SOLD dataset by Ranasinghe et al.
@@
 Ranasinghe et al.
-[278] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task101101101https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
+[289] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task107107107https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
 In completing the task, the study by Narayan et al.
@@
 [172] have proposed a system named Aruth based on the Lesk Algorithm[301].
-An online tool102102102http://aruth.herokuapp.com/, an API103103103https://bit.ly/3sJEYbS of the algorithm, and code along with data on github104104104https://github.com/jseanm1/aruthSWSD are available.
+An online tool108108108http://aruth.herokuapp.com/, an API109109109https://bit.ly/3sJEYbS of the algorithm, and code along with data on github110110110https://github.com/jseanm1/aruthSWSD are available.
 For the same task, Marasinghe et al.
@@
 Hasan et al.
-[141] introduced the XL-sum105105105https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website106106106https://www.bbc.com/sinhala.
-The M3LS107107107https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
-[142] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website108108108https://www.bbc.com/sinhala.
+[144] introduced the XL-sum111111111https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website112112112https://www.bbc.com/sinhala.
+The M3LS113113113https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
+[145] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website114114114https://www.bbc.com/sinhala.
 Further, they claim that MT5 [310] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
+The first Sinhala multi-document summarizing dataset, M2DS115115115https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
+[146].
+The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary.
+Thier code and pre-trained models are also available on github116116116https://github.com/KushanMH/m2ds.
 3.16 Other Semantic Tools
@@
 [313] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems.
-The Massively Multilingual Speech (MMS)109109109https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
+The Massively Multilingual Speech (MMS)117117117https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
 [138] has Sinhala data for the Spoken Language Identification (LID) task.
@@
 However, it was shown that the new model slightly underperforms compared to the state-of-the-art result.
+Later, they conducted further development on their model in a follow-up work [358].
 Karunathilaka et al.
@@
 TacoSi introduced by Arachchige and Weerasinghe [365] is based on Tacotron [366] and has been evaluated with 10 human evaluators to determine its text-to-speech quality.
-Nanayakkara [355], Nanayakkara and Weerasinghe [356] used DeepSpeech110110110https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
+Nanayakkara [367], Nanayakkara and Weerasinghe [368] used DeepSpeech118118118https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
 Gunarathne et al.
-[357] used an earlier version of the CMUSphinx toolkit111111111https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
+[369] used an earlier version of the CMUSphinx toolkit119119119https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
 The later work by Akesh and Meegama [370] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [371, 372] to automatically generate Sinhala subtitles from Speech.
+The work by [373] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children.
 3.17.3 Speech-to-Speech
@@
 [375] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end.
+The work by Athas and Pirapuraj [376] uses Google APIs (speech-to-text API, text-to-text-translator API, and text-to-speech synthesizer API) to achieve voice translation for Sinhala and Tamil.
 3.17.4 Speech-to-Intent
@@
 3.18.1 Printed Text
-The XTREME-UP112112112https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP120120120https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [137] contains a Sinhala data set for the OCR task.
@@
 Anuradha et al.
-[398] used Tesseract 3113113113https://tesseract-ocr.github.io/ [399] for Sinhala OCR.
+[412] used Tesseract 3121121121https://tesseract-ocr.github.io/ [413] for Sinhala OCR.
 The later study by [414] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.
@@
 Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents.
+The work by Thamarasee and Surendra [417] used histogram-oriented gradient descriptor (HOG descriptor) [418] and SVMs to recognize 400+ variations of Sinhala characters.
 An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al.
@@
 The study by eka [441] has used CNN to recognize Sinhala handwritten text.
+Karunarathne et al.
+[442] used Gabor-initialized CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
 3.18.3 Ancient Text
@@
 Heenkenda and Fernando [448] used Inception-v3 [449], VGG-19 [450], and ResNet-50 [451] to classify Sinhala inscriptions to historical time periods.
+The work by Gunasekara et al.
+[452] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters.
 3.19 Translators
@@
 [459] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala.
+Yet another meta-analysis by Halpege [460] compared the commercially available English-Sinhala translation systems provided by Google and Bing through a comprehensive error analysis.
+This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage.
 As mentioned in Section 3.1, the study by [99] raised questions on the quality of the existing Sinhala-English corpora.
@@
 Thus they concluded that using expensive human labour for this task might not be efficient.
+The study by Team et al.
+[461] conducted a study on the quality of the translator models trained on their NLLB data set [130].
+Naturally, this includes Sinhala.
 Sen et al.
@@
 3.19.3 Singlish to Sinhala (Transliteration)
-The XTREME-UP114114114https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP122122122https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [137] contains a Sinhala data set for the transliteration task.
@@
 Sumanathilaka et al.
-[478, 479] proposed a Trie [480] data structure-based algorithm for word suggestion in Sinhala transliterations.
+[498, 499] proposed a Trie [500] data structure-based algorithm for word suggestion in Sinhala transliterations named Swa-Bhasha.
 An extended analysis of the same work was presented in a later work [501].
+A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [502] used NMT to eliminate word selection ambiguity in the transliteration process.
 Amarasekara et al.
@@
 They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole.
-Their code is available on GitHub115115115https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
+Their code is available on GitHub123123123https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
 Khiu et al.
@@
 (Comparatively, the Houlsby adapter runs for 78.65 hours).
+The work by Pramodya et al.
+[537] conducted a comparative analysis of the available models for Sinhala-Tamil translation and reaffirmed the observations of Duh et al.
+[538] that, in low-resource scenarios, SMT and NMT both may work similarly, however, NMT needs more careful tuning to fit performance.
 There have been attempts to link Sinhala NLP with Japanese by Herath et al.
@@
 3.20 Spelling and Grammar
-The open-source data driven approach proposed by Wasala et al.
+The open-source data-driven approach proposed by Wasala et al.
 [545, 546] claims to be able to check and correct spelling errors in Sinhala.
@@
 [548].
-A rule-based Sinhala spell checker named SinSpell based on Hunspell116116116http://hunspell.github.io/ was introduced by Liyanapathirana et al.
+A rule-based Sinhala spell checker named SinSpell based on Hunspell124124124http://hunspell.github.io/ was introduced by Liyanapathirana et al.
 [549].
-They have also made the tool available117117117http://nlp-tools.uom.lk/sinspell/ online for use.
+They have also made the tool available125125125http://nlp-tools.uom.lk/sinspell/ online for use.
 The study by Sithamparanathan and Uthayasanker [550] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil.
@@
 The work by Jayasuriya et al.
-[536] used a rule-based approach along with Google Translation to correct grammar is Sinhala text.
+[559] used a rule-based approach along with Google Translation to correct grammar in Sinhala text.
 A rule-based system to convert Sinhala sentences from active voice to passive voice while correcting grammatical errors was proposed by Ilukkumbura and Rupasinghe [560].
@@
 Navoda et al.
-[539] also claim to have created an automated tool to check Sinhala spelling and grammar.
+[562] also claims to have created an automated tool to check Sinhala spelling and grammar.
 3.21 Chat Bots
@@
 A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al.
-[540] where they used, the fasttext model trained by Facebook [113, 114, 115], on a RASA118118118https://rasa.com/ [541] chat bot.
+[563] where they used, the fasttext model trained by Facebook [116, 117, 118], on a RASA126126126https://rasa.com/ [564] chat bot.
 A Sinhala chat bot for train information was proposed by Harshani [565].
@@
 As far as we can observe, there does not seem to be a technical or functional difference between the two.
+The work by Chithrani et al.
+[580] compares a few Sinhala sign language recognition works [581, 582, 583, 584, 585, 586, 579, 587] against work conducted for other languages.
 In the domain of Sinhala sign language, Liyanaarachchi et al.
@@
 Charuka et al.
-[143] have published a dynamic word-level Sinhala Sign Language video dataset of 50 classes.
+[147] has published a dynamic word-level Sinhala Sign Language video dataset of 50 classes.
 The dataset contains 1110 videos.
@@
 The work by Perera [602] uses CNN for Sri Lanka Sign Language recognition.
+The work by Nadeesha and Wasalthilaka [603] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos.
+They reported a success rate of 86.7%.
 Strides have been made in the domains of computer interpreting for written Sinhala [604] and animation of finger-spelled words and number signs [605].
@@
 [606] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
+The work by Samarasinghe et al.
+[582] used a CNN model to convert Sinhala sign language to Sinhala text.
 Kumar et al.
 [607] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language.
+The work by dha [586] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs.
+Dahanayaka et al.
+[584] also used CNN in their multi-model approach to recognize Sinhala sign language.
 Both the approaches by Perera and Kumarika [608] and Herath and Ishanka [609] integrated MediaPipe [610] with Deep Neural Networks to interpret Sri Lankan sign language.
@@
 Rishan et al.
-[556] uses Leap Motion Controller119119119https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
+[579] uses Leap Motion Controller127127127https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
 Dissanayake et al.
@@
 The study by Fazal and Farook [625] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook.
-The study by uda [598] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection.
+The study by Udawatta et al.
+[626] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection.
+Uthpala and Thirukumaran [264] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos.
+However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
 3.27 Language Acquisition and Evaluation
@@
 [631] uses SVM and LSTm to identify phonological disorders from Sinhala speech.
-3.28 Embeddings and Other Language Models
+3.28 Embeddings, LLMs, and Other Language Models
 A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al.
@@
 Dhananjaya et al.
-[101] has published two Sinhala BERT models: SinBERT-large120120120https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small121121121https://huggingface.co/NLPC-UOM/SinBERT-small.
-A smaller BERT model named SinhalaBERTo122122122https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
+[104] has published two Sinhala BERT models: SinBERT-large128128128https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small129129129https://huggingface.co/NLPC-UOM/SinBERT-small.
+A smaller BERT model named SinhalaBERTo130130130https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
 This BERT model has been created by training vanilla RoBERTa [633] on the OSCAR [148, 149] dataset.
-Üstün et al. [606] introduced Aya, a massively multilingual generative language model123123123https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [148].
+Üstün et al. [634] introduced Aya, a massively multilingual generative language model131131131https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [152].
 Wickramasinghe and de Silva [141] used Sinhala-English dictionary data [140] for obtaining Sinhala-English embedding alignment.
 Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
-The relevant code and data can be accessed on GitHub124124124https://bit.ly/3t3SKu7.
+The relevant code and data can be accessed on GitHub132132132https://bit.ly/3t3SKu7.
+Jayakody and Dias [635] conducted a comparative analysis on LLMs for Sinhala.
+In it, they observed that the small (7̃B) models of Claude133133133https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o134134134https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)135135135https://llama.meta.com/docs/get-started/ and Mistral 7B136136136https://docs.mistral.ai/.
+Further, they reported that they had to exclude Falcon 2137137137https://falconllm.tii.ae/ and Gemini138138138https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+3.29 Text Simplification
+The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [636] included the NSINA [105] Sinhala corpus.
+The aggregated data from the shared task is available online as Train-Test data139139139https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data140140140https://huggingface.co/datasets/MLSP2024/MLSP2024.
+Enomoto et al.
+[637] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages.
+Goswami et al.
+[638] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model.
+Seneviratne and Suominen [639] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task.
+In the submission by Cristea and Nisioi [640] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task.
+Their Sihala models are trained only using Sinhala text Google translated to English.
+As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
 3.30 Miscellaneous Applications
-In this section, we discuss NLP tools and research which are either hard to categorize under above sections or reasonably involve multiples of them.
+In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them.
 Dissanayake and Hettige [641] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs.
@@
 [642] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha.
-On the other hand, Amarasinghe and Ranathunga [609] used an ontology-based approch to generate Sinhala essay questions.
-The work by Liyanage and Ranathunga [610, 611] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages.
-The follow-up work by Niyarepola et al.
-[612] used pre-trained mBART [471] and mT5 [299, 613] models for the same task and showed better results.
+Romero et al.
+[154] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala.
+They have publicly released both their data141141141https://huggingface.co/datasets/afaji/cvqa and leader board142142142https://eval.ai/web/challenges/challenge-page/2305/overview.
+On the other hand, Amarasinghe and Ranathunga [643] used an ontology-based approach to generate Sinhala essay questions.
 The study by Kao and Ilmini [644] used LSTM to generate Sinhala lyrics.
@@
 Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API.
+The work by Liyanage and Ranathunga [646, 647] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages.
+The follow-up work by Niyarepola et al.
+[648] used pre-trained mBART [491] and mT5 [310, 649] models for the same task and showed better results.
 Fernando [650] proposed a method for inexact matching of Sinhala proper names.
@@
 Hewapathirana and Sumanathilaka [674] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
+The study by Herath and Wijayasiriwardhane [675] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts.
+The study by [208] used NER to cluster Sinhala news articles.
 4 Primary Sources
@@
 In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [161, 157, 160], Sinhala-Sinhala [156, 159], and English-Sinhala-Tamil [158].
-A number of NLP applications have utilized first sources intended to teach children [655, 656, 657, 658, 659] or foreigners [660, 193, 661]125125125Note that [661] is an extension of [193]..
+A number of NLP applications have utilized first sources intended to teach children [690, 691, 692, 693, 694] or foreigners [695, 200, 696]143143143Note that [696] is an extension of [200]..
 This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@
 Figure 7: The Probability of research from an institution citing that of other institutions.
-Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext127127127https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
+Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext145145145https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
 With all those limitations in mind, we still can make a few Interesting observations.

### v24

@@
 As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field.
-This manuscript is at version $5.5.0$.
+This manuscript is at version $5.6.0$.
 Index Terms:
@@
 This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala.
-To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10].
+To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [9, 10, 11].
 However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch.
@@
 It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22].
+An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [23] using the SinMin Sinhala corpus created by Upeksha et al.
+[24, 25] is given in Appendix B.
 Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [26, 27]), modern natural language processing tools for the Sinhala language are scarce [28, 29].
 Further, Sinhala is a diglossic language [30] which has two variations: the literary form and the spoken form.
-The Vedda (Veddah) language is considered a Sinhala creole [27, 28] of a language isolate [29].
+The Vedda (Veddah) language is considered a Sinhala creole [31, 32, 33] or a language isolate [34].
 Natural Language Processing (NLP) is a broad area covering all computational processing and analysis of human languages.
@@
 For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io.
-This manuscript is at version $5.5.0$.
+This manuscript is at version $5.6.0$.
 The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
-Figure 6 in Appendix B shows the most prolific researchers in the domain of Sinhala NLP.
+Given that this work was initially published in 2019, when citing this work, it is recommended to cite the most recent version to avoid the reviewers questioning its continued contemporary relevance.
+An example of an updated reference is given in Code Block LABEL:lst:ref.
+@article{de2025survey,
+title={{Survey on Publicly Available Sinhala Natural Language Processing Tools and Research}},
+author={de Silva, Nisansa},
+journal={arXiv preprint arXiv:1906.02358v24},
+year={2025}
+Code Block 1: Example of an updated reference
+Figure 7 in Appendix C shows the most prolific researchers in the domain of Sinhala NLP.
 The nodes contain the name of the researcher along with the total number of Sinhala NLP papers that the researcher has authored.
@@
 It is observable that the cluster from the Department of Computer Science & Engineering, University of Moratuwa is the most prolific in Sinhala NLP research.
-Figure 7 in Appendix B shows the probability of studies from the institutions to which the most prolific authors from Figure 6 are affiliated citing institutions of the same set.
+Figure 8 in Appendix C shows the probability of studies from the institutions to which the most prolific authors from Figure 7 are affiliated citing institutions of the same set.
 We observe both interesting and disturbing trends which we discuss in the figure caption.
-Figure 8 in Appendix B shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper.
+Figure 9 in Appendix C shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper.
 If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication.
@@
 For any language, the key for NLP applications and implementations is the existence of adequate corpora.
-On this matter, a relatively substantial Sinhala text corpus [^10] https://osf.io/a5quv/ was created by Upeksha et al.
+On this matter, a relatively substantial Sinhala text corpus named SinMin [^10] https://osf.io/a5quv/ was created by Upeksha et al.
 [24, 25] by web crawling.
@@
 [108].
+ImaniGooghari et al.
+[109] published Glot500272727https://github.com/cisnlp/Glot500, a corpus of 500 languages that includes Sinhala.
+The have 7,293,178 sentences in their Sinhala corpus.
+Further, they train a Sinhala language model and report the results of various tasks such as: sentence retrieval, NER, PoS, Text Classification, and round-trip alignment.
+They also report their result on perplexity analysis.
 As for Sinhala-Tamil corpora, Hameed et al.
@@
 Zhang et al.
-[127] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100424242https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT).
+[131] sampled data from OPUS for 100 languages (including Sinhala) and created the OPUS100434343https://github.com/EdinburghNLP/opus-100-corpus data set as a benchmark for neural machine translation (NMT).
 Thier code is also available on GitHub [^44] https://github.com/bzhangGo/zero.
@@
 [24] and derived a Sinhala writing style classification data set [^61] https://bit.ly/3PARuUS.
+Another Sinhala writing style identification data set [^62] https://github.com/cipherdragon/SimpleAA for author identification has been created by Faumi et al.
+[136].
 A data set of Sinhala swear and/or obscene words is publicly available in both Unicode [^63] https://bit.ly/3R1bdOc and Singlish [^64] https://bit.ly/3S3Z1O2 formats.
@@
 In the same paper, they also introduce the dataset SemiSOLD [^68] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach.
-The relevant code is also available on Github  [^67] https://github.com/Sinhala-NLP/SOLD.
+The relevant code is also available on Github 696969%****␣03_Sinhala.tex␣Line␣75␣****https://github.com/Sinhala-NLP/SOLD.
+The subsequent work by Dmonte et al.
+[141] used the SOLD data set in the evaluation process of their new MT-Offense [^70] https://github.com/LanguageTechnologyLab/MT-Offense data set which they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
 Sinhala is included in the 2800+ language metadata set composed by van Esch et al.
@@
 Kudugunta et al.
-[143] released MADLAD-400747474https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages.
+[150] released MADLAD-400777777https://github.com/google-research/google-research/tree/master/madlad_400 a manually audited, general domain, large document level dataset spanning 419 languages.
 For Sinhala, they report that they have 788K noisy docs, 349.2K clean docs, 22.1M noisy sentences, 16M clean sentences, 3.4B noisy tokens, and 1.9B clean tokens.
@@
 Kudugunta et al.
-[153] introduced MADLAD-400878787https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents.
+[160] introduced MADLAD-400909090https://huggingface.co/datasets/allenai/MADLAD-400/tree/main/data/si a large audited dataset of multilingual documents.
 They have also released the trained checkpoints on GitHub [^91] https://github.com/google-research/google-research/tree/master/madlad_400.
@@
 [161] along with data [^92] https://huggingface.co/datasets/afaji/cvqa and a leader board [^93] https://eval.ai/web/challenges/challenge-page/2305/overview.
+Pushpananda et al.
+[162] claims to have created a manually annotated parallel corpus named TamSiPara with 25k parallel sentences in the Si-Ta direction and 22k parallel sentences in the Ta-Si direction.
+But it is not freely and openly accessible.
+While the stated purpose of Ranasinghe [163] is creating a tokenizer for Sinhala and not creating a text classification data set, the repository does contain [^94] https://github.com/Buddhilive/sltk/blob/main/data/sin.csv a data set of 26k Sinhala news titles with classification labels.
+A visual Q&A benchmark data set named ALM-Bench [^95] https://huggingface.co/datasets/MBZUAI/ALM-Bench was created by Vayani et al.
+[164] to evaluate Q&A performance of LLMs on 100 languages including Sinhala.
+They initially machine translated LLaVA-Bench (In-the-Wild) dataset [165] data set and then extended it by searching for images under country name, language name, and cultural category on the web so that they can include culturally relevant questions and answer for each language.
+In their own experiment results, they report that GPT-4o [166] performance on Sinhala is superior to other tested LLMs on this data set.
+Ranathunga et al.
+[167] has created a text simplification dataset named SiTSE [^96] https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation.
+It has 1000 complex Sinhala texts along with two separate simplifications for each of them.
+Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data.
+Ranathunga et al.
+[168] created multiNER [^97] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil.
+It contains 99k tagged Sinhala words.
+They have used PER, ORG, LOC, and MISC tags as well as O.
+Singh et al.
+[169] introduced Global-MMLU [^98] https://huggingface.co/datasets/CohereForAI/Global-MMLU, a multilingual extension of the original English Massive Multitask Language Understanding (MMLU) dataset [170].
+They claim to have improved the data set beyond what Üstün et al.
+[171] did before.
+Chavinda and Thayasivam [172] introduces two publicly available Sinhala hate speech data sets.
+One is from Facebook [^99] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter100100100https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech.
+A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released101101101https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [173].
+One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish.
 3.3 Dictionaries
@@
 The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs.
-Both their code and the two versions of the dictionary are publicly available on github [^91] https://github.com/kasunw22/sinhala-para-dict/tree/main.
-A Sinhala-Sinhala dictionary named Wahara [^92] http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
+Both their code and the two versions of the dictionary are publicly available on github102102102https://github.com/kasunw22/sinhala-para-dict/tree/main.
+A Sinhala-Sinhala dictionary named Wahara103103103http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
 There exists the government-sponsored trilingual dictionary [184], which matches Sinhala, English, and Tamil.
@@
 For a time it was hosted on [190] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al.
-[172] have cloned to use in their application uploaded on github [^93] https://github.com/jseanm1/aruthSWSD.
+[191] have cloned to use in their application uploaded on github104104104https://github.com/jseanm1/aruthSWSD.
 However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product.
@@
 [202] compared the effectiveness of using RNN, LSTM, and GRU for Sinhala morphological analysis of Sinhala and Sinhlish deatasets and found the BiGRU gives the highest accuracy of 87.96%.
+The comparative study by [203] compared Sinhala morphology to that of a number of other languages in hopes of making subsequent machine translation work easier.
 TABLE V: Morphological Analyzers comparison
@@
 Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [211].
-A RESTFul PoS tagging web service created by Jayaweera and Dias [192] using the above research can still be accessed [^94] http://bit.ly/2F0jKid via POST and GET.
+A RESTFul PoS tagging web service created by Jayaweera and Dias [212] using the above research can still be accessed105105105http://bit.ly/2F0jKid via POST and GET.
 A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al.
@@
 The Ananya system by Manamini et al.
-[114] is available to download at GitHub  [^95] http://bit.ly/2XrwCoK.
+[118] is available to download at GitHub 106106106http://bit.ly/2XrwCoK.
 The data and code for the approaches by Dahanayaka and Weerasinghe [222] and by Senevirathne et al.
@@
 They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
+A data set for Sinhala NER named multiNER107107107https://github.com/suralk/multiNER was introduced by Ranathunga et al.
+[168].
+It includes parallel NER data for Sinhala, Tamil, and English.
 3.9 Semantic Similarity
@@
 [233].
+Deepal et al.
+[234] proposed to use Siamese network which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity.
+They report an F1 of 0.9041 for their model calculated by cosine distance.
+They discuss how they used a number of publicly available data sets and refined them to be used in their task, nevertheless, they do not publicly release the resulting data set.
 3.10 Text Classification
@@
 A smaller implementation of Sinhala news classification has been attempted by de Silva [28].
-As mentioned in Section 3.2, their news corpus is publicly available [^96] https://osf.io/tdb84/.
+As mentioned in Section 3.2, their news corpus is publicly available108108108https://osf.io/tdb84/.
 Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [236].
@@
 A simple TF-IDF based Sinhala text classification system was proposed by Koralage [251].
-The study by [231] used Genetic Algorithm on Lucene [^97] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
+The study by [252] used Genetic Algorithm on Lucene109109109https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
 The approach proposed by Rathnayake et al.
@@
 [107].
-Both their code [^98] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9999993https://tinyurl.com/42un7a9y, Sinhala News source classification data set100100100https://bit.ly/3PBTyfo, Sinhala News category classification data set101101101https://bit.ly/3AsfK7u, and Sinhala writing style classification data set102102102https://bit.ly/3PARuUS) are publicly available.
+Both their code110110110https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1111111113https://tinyurl.com/42un7a9y, Sinhala News source classification data set112112112https://bit.ly/3PBTyfo, Sinhala News category classification data set113113113https://bit.ly/3AsfK7u, and Sinhala writing style classification data set114114114https://bit.ly/3PARuUS) are publicly available.
+Faumi et al.
+[136] also worked on Sinhala writing style identification (author identification) using their own data set115115115https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
 The study by Wijayarathna and Jayalal [264] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter.
@@
 A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [269] based on their prior work [270].
-A word2vec based tool103103103http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
+A word2vec based tool116116116http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
 A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al.
 [271]Demotte et al.
-[251] proposed a sentiment analysis system based on sentence-state LSTM Networks for Sinhala news comments.
+[272] proposed a sentiment analysis system117117117https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments.
 In the subsequent work [273], they used word similarity to generate a Sinhala semantic lexicon.
@@
 [244]Aththanayaka and Naleer [283] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media.
-The ACTSEA dataset104104104https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
+The ACTSEA dataset118118118https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
 [135].
@@
 They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
+Liyanaarachchi et al.
+[286] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative.
+They claim that the 1-of-m embedding outperforms Log-m embedding for this task.
+Bandaranayake and Usoof [287] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [244].
+They report that XLMR-large model achieves the best result.
+Fernando and Ranathunga [288] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations.
 3.12 Hate Speech Detection
@@
 A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [290].
-A feature model and a data set105105105https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [124]Sandaruwan et al.
+A feature model and a data set119119119https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al.
 [291] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques.
@@
 [129] analyses Sinhala hate speech propagation on Twitter.
-They have also published a dataset106106106https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
+They have also published a dataset120120120https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
 Rajapaksha et al.
@@
 They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
+Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook121121121https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter122122122https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets.
+They report TwHIN-BERT-base [310] model to have obtained the best results.
 After the release of the SOLD dataset by Ranasinghe et al.
@@
 Ranasinghe et al.
-[289] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task107107107https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
+[314] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task123123123https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
 In completing the task, the study by Narayan et al.
@@
 [316] used mBERT to identify hate and offensive content in Sinhala.
+The work by Rostamkhani and Eetemadi [317] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models.
+They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD.
+Dmonte et al.
+[141] used the SOLD data set to verify the new MT-Offense124124124https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
+They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].
 3.13 Fake News Detection
@@
 [191] have proposed a system named Aruth based on the Lesk Algorithm[327].
-An online tool108108108http://aruth.herokuapp.com/, an API109109109https://bit.ly/3sJEYbS of the algorithm, and code along with data on github110110110https://github.com/jseanm1/aruthSWSD are available.
+An online tool125125125http://aruth.herokuapp.com/, an API126126126https://bit.ly/3sJEYbS of the algorithm, and code along with data on github127127127https://github.com/jseanm1/aruthSWSD are available.
 For the same task, Marasinghe et al.
@@
 Hasan et al.
-[144] introduced the XL-sum111111111https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website112112112https://www.bbc.com/sinhala.
-The M3LS113113113https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
-[145] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website114114114https://www.bbc.com/sinhala.
+[151] introduced the XL-sum128128128https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website129129129https://www.bbc.com/sinhala.
+The M3LS130130130https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
+[152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website131131131https://www.bbc.com/sinhala.
 Further, they claim that MT5 [336] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
-The first Sinhala multi-document summarizing dataset, M2DS115115115https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
+The first Sinhala multi-document summarizing dataset, M2DS132132132https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
 [153].
 The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary.
-Thier code and pre-trained models are also available on github116116116https://github.com/KushanMH/m2ds.
+Thier code and pre-trained models are also available on github133133133https://github.com/KushanMH/m2ds.
 3.16 Other Semantic Tools
@@
 [339] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems.
-The Massively Multilingual Speech (MMS)117117117https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
+The Massively Multilingual Speech (MMS)134134134https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
 [145] has Sinhala data for the Spoken Language Identification (LID) task.
@@
 [361] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students.
+Praveen et al.
+[362] claim to have created a Sinhala TTS system with a custom neural network that archives 98% accuracy.
 Anuradha and Thelijjagoda [363] proposed a machine translation system to convert Sinhala and English Braille documents into voice.
@@
 TacoSi introduced by Arachchige and Weerasinghe [392] is based on Tacotron [393] and has been evaluated with 10 human evaluators to determine its text-to-speech quality.
-Nanayakkara [367], Nanayakkara and Weerasinghe [368] used DeepSpeech118118118https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
+Nanayakkara [394], Nanayakkara and Weerasinghe [395] used DeepSpeech135135135https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
 Gunarathne et al.
-[369] used an earlier version of the CMUSphinx toolkit119119119https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
+[396] used an earlier version of the CMUSphinx toolkit136136136https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
 The later work by Akesh and Meegama [397] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [398, 399] to automatically generate Sinhala subtitles from Speech.
-The work by [373] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children.
+The work by Wickramaarachchi et al.
+[400] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children.
+Dissanayaka et al.
+[401] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking.
+tha [402] have introduced a Sinhala text to Speech dataset137137137https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes.
+As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [403].
+While the data set link points to a github repository, it does not contain the data set; it points to a project website138138138https://sita-speakerdiarization.github.io/ which then points to a download page139139139https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
 3.17.3 Speech-to-Speech
@@
 [416] came up with Phoneme-Viseme mapping for Sinhala speech that they intended to be used for future applications of robotics.
+3.17.7 Music to Notation
+The work by Dulmi and Hettige [417] has created a pipeline which converts audio recordings to western music notation which is then converted to Sinhala musical notations using an API call to a Large Language Model with a custom-tailored prompt.
 3.18 Optical Character Recognition Applications
@@
 3.18.1 Printed Text
-The XTREME-UP120120120https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP140140140https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [144] contains a Sinhala data set for the OCR task.
@@
 Anuradha et al.
-[412] used Tesseract 3121121121https://tesseract-ocr.github.io/ [413] for Sinhala OCR.
+[443] used Tesseract 3141141141https://tesseract-ocr.github.io/ [444] for Sinhala OCR.
 The later study by [445] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.
@@
 [483] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters.
+The work by Surasinghe and Thanikasalam [484] used CNN models to predict the historical period of ancient Sinhala text.
+In their subsequent work [485], they extended this work by adding attention to the CNN models.
 3.19 Translators
@@
 [523] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [98] and reported better results than mBART [524].
+The study by Ranathungaa et al.
+[525] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models.
+In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [288] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations.
+Sinhala is one of the languages considered in the short survey by Perera and Sumanathilaka [526] on machine translation and transliteration for Indo-Aryan Languages.
+They mention a sub-set of works [505, 509, 527, 508, 528, 529] discussed in this survey for EnSi translation.
 3.19.3 Singlish to Sinhala (Transliteration)
-The XTREME-UP122122122https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP142142142https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [144] contains a Sinhala data set for the transliteration task.
@@
 [535] introduced an English-to-Sinhala transliteration system.
+Athukorala and Sumanathilaka [536] proposed a fuzzy logic-based Sinhala transliteration system named Swa-Bhasha.
 Sumanathilaka et al.
-[498, 499] proposed a Trie [500] data structure-based algorithm for word suggestion in Sinhala transliterations named Swa-Bhasha.
+[537, 538] then proposed a Trie [539] data structure-based algorithm for word suggestion for the Swa-Bhasha system.
 An extended analysis of the same work was presented in a later work [540].
 A further extension, Swa-Bhasha 2.0, by Dharmasiri and Sumanathilaka [541] used NMT to eliminate word selection ambiguity in the transliteration process.
+Subsequently, Sumanathilaka et al.
+[542] claim to have released the Swa-Bhasha Dataset.
 Amarasekara et al.
@@
 They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole.
-Their code is available on GitHub123123123https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
+Their code is available on GitHub143143143https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
 Khiu et al.
 [550] fine-tuned mBART on Sinhala Government Corpus and Bible corpus to predict the NMT performance dependence on domain similarity.
+De Mel et al.
+[551] conducted a comparative analysis for Sinhala transliteration between an extension of the rule-based approach proposed by Tennage et al.
+[552] and the transliteration as a translation model proposed by Deselaers et al.
+[553].
+They reported across-the-board better performance with the latter.
+All thier code as well as the test-sets are publicly avaialble144144144https://github.com/kasunw22/Sinhala-Transliterator/.
+Perera et al.
+[554] attempted the same task with BERT.
+Their code is also publicly available145145145https://github.com/Sameera2001Perera/Singlish-Transliterator.
+In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [526] covers a number of works [504, 552, 531, 532, 534, 536, 535, 538, 541, 542] on Sinhala transliteration which we have already discussed above.
 3.19.4 Singlish and English (Translation via Transliteration)
@@
 [579] that, in low-resource scenarios, SMT and NMT both may work similarly, however, NMT needs more careful tuning to fit performance.
+The short survey [526] on Indo-Aryan Languages machine translation and transliteration, discusses a few works [527, 572, 509, 571, 569, 573] under SiTa translation which have already been discussed in this survey.
 There have been attempts to link Sinhala NLP with Japanese by Herath et al.
@@
 Jayasinghe [581] discusses the importance of translation between Sinhala and the liturgical language of Buddhism, Pali [582, 583, 584].
-[544] attempted to use a dictionary-based machine-translation method for this task.
+Shalini and Hettige [585] attempted to use a dictionary-based machine-translation method for this task.
 The study by Anuradha and Thelijjagoda [363] uses machine translation on the unique application of converting Sinhala and English Braille documents, which they have run OCR on, into voice.
@@
 [589].
-A rule-based Sinhala spell checker named SinSpell based on Hunspell124124124http://hunspell.github.io/ was introduced by Liyanapathirana et al.
+A rule-based Sinhala spell checker named SinSpell based on Hunspell146146146http://hunspell.github.io/ was introduced by Liyanapathirana et al.
 [590].
-They have also made the tool available125125125http://nlp-tools.uom.lk/sinspell/ online for use.
+They have also made the tool available147147147http://nlp-tools.uom.lk/sinspell/ online for use.
 The study by Sithamparanathan and Uthayasanker [591] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil.
@@
 [594] to correct real-word errors in Sinhala text.
+Praveen et al.
+[362] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.
 A model for detecting grammatical mistakes in Sinhala was developed by Pabasara and Jayalal [595].
@@
 A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al.
-[563] where they used, the fasttext model trained by Facebook [116, 117, 118], on a RASA126126126https://rasa.com/ [564] chat bot.
+[604] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA148148148https://rasa.com/ [605] chat bot.
 A Sinhala chat bot for train information was proposed by Harshani [606].
@@
 Rishan et al.
-[579] uses Leap Motion Controller127127127https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
+[620] uses Leap Motion Controller149149149https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
 Dissanayake et al.
@@
 They later used it for NMT in Sinhala-English Code-Mixed text [517].
+Shakir and Deuber [666] claims to have developed a corpus of South Asian languages in the context of code-mixing which includes Sinhala.
+However, this is not publicly available.
+It should further be noted that they report Sinhala data using the heading Sri Lanka rather than Sinhala.
+In their follow-up work Shakir and Deuber [667] claim to extend this code-mixed data set to include meme data.
 As discussed in Section 3.10Rathnayake et al.
@@
 [674] uses SVM and LSTm to identify phonological disorders from Sinhala speech.
+Francis et al.
+[675] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students.
+Praveen et al.
+[362] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language.
+They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system.
 3.28 Embeddings, LLMs, and Other Language Models
@@
 Dhananjaya et al.
-[104] has published two Sinhala BERT models: SinBERT-large128128128https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small129129129https://huggingface.co/NLPC-UOM/SinBERT-small.
-A smaller BERT model named SinhalaBERTo130130130https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
+[107] has published two Sinhala BERT models: SinBERT-large150150150https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small151151151https://huggingface.co/NLPC-UOM/SinBERT-small.
+A smaller BERT model named SinhalaBERTo152152152https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
 This BERT model has been created by training vanilla RoBERTa [677] on the OSCAR [155, 156] dataset.
-Üstün et al. [634] introduced Aya, a massively multilingual generative language model131131131https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [152].
+Demotte et al.
+[272] have released the word embeddings153153153https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code154154154https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer.
+Gurgurov et al.
+[678] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository.
+They credit Demotte et al.
+[272] as their source for Sinhala embeddings.
 Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment.
 Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
-The relevant code and data can be accessed on GitHub132132132https://bit.ly/3t3SKu7.
+The relevant code and data can be accessed on GitHub155155155https://bit.ly/3t3SKu7.
+Subsequent work by Rathnayake et al.
+[679] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [680]) as the baseline.
+They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [681] archives the best result of 29.49 for En-Si FastText embedding.
+Üstün et al.
+[171] introduced Aya, a massively multilingual generative language model156156156https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159].
 Jayakody and Dias [682] conducted a comparative analysis on LLMs for Sinhala.
-In it, they observed that the small (7̃B) models of Claude133133133https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o134134134https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)135135135https://llama.meta.com/docs/get-started/ and Mistral 7B136136136https://docs.mistral.ai/.
-Further, they reported that they had to exclude Falcon 2137137137https://falconllm.tii.ae/ and Gemini138138138https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+In it, they observed that the small (7̃B) models of Claude157157157https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o158158158https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)159159159https://llama.meta.com/docs/get-started/ and Mistral 7B160160160https://docs.mistral.ai/.
+Further, they reported that they had to exclude Falcon 2161161161https://falconllm.tii.ae/ and Gemini162162162https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+Chang et al.
+[683] introduced Goldfish, a collection of monolingual language models163163163https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109].
+The work by Ayesha et al.
+[684] used a fined-tuned Llama 3.1 (8B) model [685] to predict student performance by analyzing textual entries created by students in Sinhala and English.
+Vayani et al.
+[164] analyzed the performance of a number of LLMs over 100 languages that included Sinhala.
+The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165].
+This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category.
+This dataset, which they named ALM-Bench164164164https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available.
+They made a number of observations in this aspect.
+First, they observed that for the Q&A task, the performance of GPT-4o [166] for Sinhala is better than Qwen2-VL [686].
+Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%.
+In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
 3.29 Text Simplification
 The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [687] included the NSINA [108] Sinhala corpus.
-The aggregated data from the shared task is available online as Train-Test data139139139https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data140140140https://huggingface.co/datasets/MLSP2024/MLSP2024.
+The aggregated data from the shared task is available online as Train-Test data165165165https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data166166166https://huggingface.co/datasets/MLSP2024/MLSP2024.
 Enomoto et al.
@@
 As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
+3.30 Tokenizing
+The study by Petrov et al.
+[692] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala.
+They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest).
+They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range.
+They report reasonable results on MBart50 [693] and mT5 [336, 693] with $1.35$ and $1.66$ respectively.
+The best result ($1.00$), is given by CANINE [694].
+The work by Velayuthan and Sarveswaran [695] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala.
+Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available167167167https://github.com/Buddhilive/sltk.
+However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words.
+While the authors have made the dictionary publicly available168168168https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries.
+So while it may be adequate to handle Sri Lanka-centric news title data set169169169https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
+3.31 Detection or Management of Mental Illnesses
+Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [668] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook.
+Hewapathirana and Sumanathilaka [696] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
+The study by Herath and Wijayasiriwardhane [697] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts.
+Fazal and Farook [698] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [699] obtains the best results.
+Welarathna et al.
+[413] used CNNs to classify emotions of Sinhala speech by Autistic children.
+They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry.
+Jayawardena et al.
+[700] used a NAO Robot [701] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children.
+3.32 Sinhala Text Generation
+The study by Kao and Ilmini [702] used LSTM to generate Sinhala lyrics.
+Bandara et al.
+[703] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala.
+However, they did not train a GTP-2 model in Sinhala.
+Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API.
+Amarasinghe and Ranathunga [704] used an ontology-based approach to generate Sinhala essay questions.
+Ranasinghe et al.
+[705] trained a Sinhala image caption generator model on Google translated MS COCO [706] dataset and tested it on Google translated Flickr8K [707] and MS COCO datasets.
+The work by Liyanage and Ranathunga [708, 709] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages.
+The follow-up work by Niyarepola et al.
+[710] used pre-trained mBART [524] and mT5 [336, 693] models for the same task and showed better results.
 3.33 Miscellaneous Applications
@@
 [161] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala.
-They have publicly released both their data141141141https://huggingface.co/datasets/afaji/cvqa and leader board142142142https://eval.ai/web/challenges/challenge-page/2305/overview.
-On the other hand, Amarasinghe and Ranathunga [643] used an ontology-based approach to generate Sinhala essay questions.
-The study by Kao and Ilmini [644] used LSTM to generate Sinhala lyrics.
-Bandara et al.
-[645] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala.
-However, they did not train a GTP-2 model in Sinhala.
-Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API.
-The work by Liyanage and Ranathunga [646, 647] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages.
-The follow-up work by Niyarepola et al.
-[648] used pre-trained mBART [491] and mT5 [310, 649] models for the same task and showed better results.
+They have publicly released both their data170170170https://huggingface.co/datasets/afaji/cvqa and leader board171171171https://eval.ai/web/challenges/challenge-page/2305/overview.
+Vayani et al.
+[164] created a dataset ALM-Bench172172172https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web.
+They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [166] is superior to other tested LLMs.
+Wansekara and Jayasekara [713] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text.
 Fernando [714] proposed a method for inexact matching of Sinhala proper names.
@@
 [528].
-Ranasinghe et al. [656] trained a Sinhala image caption generator model on Google translated MS COCO [657] dataset and tested it on Google translated Flickr8K [658] and MS COCO datasets.
 Rajitha et al.
@@
 However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8.
-The study by Petrov et al.
-[669] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala.
-They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest).
-They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range.
-They report reasonable results on MBart50 [649] and mT5 [310, 649] with $1.35$ and $1.66$ respectively.
-The best result ($1.00$), is given by CANINE [670].
 A meta-study on how language data from social media are used in research was conducted by Hewapathirana [729].
@@
 Senarathne [731] introduced a model to translate Sinhala pseudocode to computer code written in C#.
-Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [625] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook.
-Hewapathirana and Sumanathilaka [674] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
-The study by Herath and Wijayasiriwardhane [675] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts.
-The study by [208] used NER to cluster Sinhala news articles.
+The work by De Silva and Athukorala [732] claims to be able to provide Sinhala language explanations to JAVA code as well as generate JAVA code snippets from Sinhala descriptions.
+The study by [228] performs best in this regard.
+used NER to cluster Sinhala news articles.
+Subasinghe and Samarawickrama [733] analysed the Noto Sans Sinhala173173173https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens.
+Zhao et al.
+[734, 735] discussed a Sinhala LLM-based system they utilized to conduct participatory research in Sri Lanka.
+While they discussed how the LLM system helped their goals and briefly touched on how they utilised a hybrid translation approach of NMT and rule-based methods to circumvent the lack of training data, the specifics of their LLM system are not given.
 4 Primary Sources
@@
 In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [180, 176, 179], Sinhala-Sinhala [175, 178], and English-Sinhala-Tamil [177].
-A number of NLP applications have utilized first sources intended to teach children [690, 691, 692, 693, 694] or foreigners [695, 200, 696]143143143Note that [696] is an extension of [200]..
+A number of NLP applications have utilized first sources intended to teach children [750, 751, 752, 753, 754] or foreigners [755, 220, 756]174174174Note that [756] is an extension of [220]..
 This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@
 In many cases where similar work is done, it is a re-hashing on the same ideas adopted from resource-rich languages because of, the unavailability of (or the reluctance to), referring and building on work done by another group (Refer Fig.
-7 in Appendix B).
+8 in Appendix C).
 This has resulted in multiple groups building multiple foundations behind closed doors but no one ending up with a completed end-to-end NLP workflow.
@@
 1998 from the Iskoola Pota UNICODE font by Microsoft [786].
-Appendix B Author Meta Analysis
+Appendix B Distribution of Sinhala Letters Towards the Beginning, Middle, or End of Words
+Figure 6: An adaptation to Sinhala of the character position visualization proposed by ProoffReader [23].
+For this, we have used a portion of the SinMin corpus created by Upeksha et al.
+[24, 25].
+Our sample included 3,990,838 unique words which appear more than once in the corpus.
+While the Spearman correlation used in this method dilutes the notability aspect such as pure Sinhala vowels only appearing as the first character, it brings to the front how diacritics, on the other hand, are generally biased towards the end.
+Appendix C Author Meta Analysis
 Figure 7: Co-author graph of the most prolific researchers in the Sinhala NLP domain (Selected at the threshold of at least 3 publications)
 Figure 8: The Probability of research from an institution citing that of other institutions.
-Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext145145145https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
+Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext176176176https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
 With all those limitations in mind, we still can make a few Interesting observations.

### v25

@@
 As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field.
-This manuscript is at version $5.6.0$.
+This manuscript is at version $5.7.0$.
 Index Terms:
@@
 However, unlike English, which is part of the Germanic branch, Sinhala belongs to the Indo-Aryan branch.
+An extremely simplified version of the Indo-European language tree focusing mainly on Sinhala and its close relatives is shown in Fig 1.
 Further, Sinhala, unlike English, which borrowed the Latin alphabet, has its own writing system, which is a descendant of the Indian Brahmi script [12, 13, 14, 15, 16, 17, 18].
@@
 Thus by inheritance, the Sinhala writing system is abugida (alphasyllabary), which to say, that consonant-vowel sequences are written as single units [21].
-We show the evolutionary eras of the Sinhala script in Appendix A.
+An extremely simplified evolution tree of the Sinhala script is shown in Fig 2.
+In it, the Sinhala script is shown as a singular derivation from Brahmi directly appearing with the contemporary script.
+But this is a simplification.
+We show the actual evolutionary eras of the Sinhala script in Appendix A where the gradual change from Brahmi to the contemporary script is clearer.
 It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22].
@@
 [24, 25] is given in Appendix B.
+Figure 1: A portion of the Indo-European language tree with a focus on Sinhala and its close relatives.
+Only the Elu sub-tree is shown exhaustively.
+All other sub-trees only list some languages for representation purposes chosen randomly or otherwise.
+Some of the languages shown as Extinct have archaeological proof (eg: Prakrit) while others were retroactively reconstructed by linguists by observing the commonalities of languages that they begot (eg: Proto-Indo-Aryan).
+The status of some languages has spread beyond linguistics to be a social and political question.
+In this depiction, we have attempted to side with academic consensus when available and the majority when not.
+As an example of one contentions point on this diagram note the positioning of Pali.
+While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
+Figure 2: A portion of the writing systems (scripts) tree with a focus on Sinhala and its close relatives.
+It is important to note that this is distinct from the language tree shown in Fig 1.
+The writing systems used by a language being related, not being related, or the degree to which they are related is not directly correlated to the relevant languages being related, not being related, or the degree to which they are related.
+As an example, note that the writing systems of Sinhala and Tamil are reasonably closely related.
+But Tamil is not shown on Fig 1 given that it does not belong to the Indo-European language family.
+It belongs to the Dravidian language family.
+A language that is shown as a single entity in Fig 1 may have multiple entries in this diagram.
+An example is the language shown as Hindustani splits into Hindi and Urdu where the former uses the Devanagari script and the latter a derivation of Arabic script.
+Conversely, languages that may be shown as distinct entities in the full language tree, may map to a singular entity in the script tree.
+An example is the fact that English which would go under Proto-Germanic and Italian which would go under Proto-Italic in and extended version of Fig 1, would both map to Latin in this diagram given that both of them use the Latin script.
 Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [26, 27]), modern natural language processing tools for the Sinhala language are scarce [28, 29].
@@
 For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io.
-This manuscript is at version $5.6.0$.
+This manuscript is at version $5.7.0$.
 The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@
 author={de Silva, Nisansa},
-journal={arXiv preprint arXiv:1906.02358v24},
+journal={arXiv preprint arXiv:1906.02358v25},
 year={2025}
@@
 We observe both interesting and disturbing trends which we discuss in the figure caption.
-Figure 9 in Appendix C shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests denoted by the subsection titles of the Section 3 of this paper.
+Figure 11 in Appendix C shows the mapping between authors with at least 10 papers in the Sinhala NLP domain and their research interests, denoted by the subsection titles of the Section 3 of this paper.
 If two or more authors on the diagram have co-authored a paper, it gets counted for each of the authors separately without any bias on the author order listed on the publication.
@@
 For example, the paper Sinhala Text Classification: Observations from the Perspective of a Resource Poor Language [28] which introduces a new Sinhala text classification data set is counted for both Data Sets (Section 3.2) and Text Classification (Section 3.10).
-The remainder of this survey is organized as follows; Section 2 introduces some important properties and conventions of the Sinhala language which are important for the development and understanding of Sinhala NLP.
+The remainder of this survey is organized as follows: Section 2 introduces some important properties and conventions of the Sinhala language, which are important for the development and understanding of Sinhala NLP.
 Section 3 discusses the various tools and research available for Sinhala NLP.
-In this section we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work.
+In this section, we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work.
 We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka.
@@
 One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish.
+Burchell et al.
+[174] introduced HPLT v2102102102https://huggingface.co/datasets/HPLT/HPLT2.0_cleaned a multilingual monolingual and parallel corpora that includes 273k Sinhala Sentences.
+These sentences were extracted from the Internet Archive(IA) and the Common Crawl (CC).
+They have also publicly released their translation models103103103https://huggingface.co/collections/HPLT/hplt-20-uni-direction-translation-models-67f2fc7ae54845f9b182957a.
 3.3 Dictionaries
@@
 The study by Wickramasinghe and De Silva [147] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Sipairs.
-Both their code and the two versions of the dictionary are publicly available on github102102102https://github.com/kasunw22/sinhala-para-dict/tree/main.
-A Sinhala-Sinhala dictionary named Wahara103103103http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
+Both their code and the two versions of the dictionary are publicly available on github104104104https://github.com/kasunw22/sinhala-para-dict/tree/main.
+A Sinhala-Sinhala dictionary named Wahara105105105http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
 There exists the government-sponsored trilingual dictionary [185], which matches Sinhala, English, and Tamil.
@@
 For a time it was hosted on [191] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al.
-[191] have cloned to use in their application uploaded on github104104104https://github.com/jseanm1/aruthSWSD.
+[192] have cloned to use in their application uploaded on github106106106https://github.com/jseanm1/aruthSWSD.
 However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product.
@@
 Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [212].
-A RESTFul PoS tagging web service created by Jayaweera and Dias [212] using the above research can still be accessed105105105http://bit.ly/2F0jKid via POST and GET.
+A RESTFul PoS tagging web service created by Jayaweera and Dias [213] using the above research can still be accessed107107107http://bit.ly/2F0jKid via POST and GET.
 A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al.
@@
 The Ananya system by Manamini et al.
-[118] is available to download at GitHub 106106106http://bit.ly/2XrwCoK.
+[118] is available to download at GitHub 108108108http://bit.ly/2XrwCoK.
 The data and code for the approaches by Dahanayaka and Weerasinghe [223] and by Senevirathne et al.
@@
 They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
-A data set for Sinhala NER named multiNER107107107https://github.com/suralk/multiNER was introduced by Ranathunga et al.
+A data set for Sinhala NER named multiNER109109109https://github.com/suralk/multiNER was introduced by Ranathunga et al.
 [168].
 It includes parallel NER data for Sinhala, Tamil, and English.
+Rathnayake and Asanka [231] simply uses LLaMA for Sinhala NER.
 3.9 Semantic Similarity
@@
 A smaller implementation of Sinhala news classification has been attempted by de Silva [28].
-As mentioned in Section 3.2, their news corpus is publicly available108108108https://osf.io/tdb84/.
+As mentioned in Section 3.2, their news corpus is publicly available110110110https://osf.io/tdb84/.
 Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [238].
@@
 A simple TF-IDF based Sinhala text classification system was proposed by Koralage [253].
-The study by [252] used Genetic Algorithm on Lucene109109109https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
+The study by [254] used Genetic Algorithm on Lucene111111111https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
 The approach proposed by Rathnayake et al.
@@
 [107].
-Both their code110110110https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1111111113https://tinyurl.com/42un7a9y, Sinhala News source classification data set112112112https://bit.ly/3PBTyfo, Sinhala News category classification data set113113113https://bit.ly/3AsfK7u, and Sinhala writing style classification data set114114114https://bit.ly/3PARuUS) are publicly available.
+Both their code112112112https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1131131133https://tinyurl.com/42un7a9y, Sinhala News source classification data set114114114https://bit.ly/3PBTyfo, Sinhala News category classification data set115115115https://bit.ly/3AsfK7u, and Sinhala writing style classification data set116116116https://bit.ly/3PARuUS) are publicly available.
 Faumi et al.
-[136] also worked on Sinhala writing style identification (author identification) using their own data set115115115https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
+[136] also worked on Sinhala writing style identification (author identification) using their own data set117117117https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
 The study by Wijayarathna and Jayalal [266] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter.
@@
 [270] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
+The work by Gunawardana and Thamarasee [271] attempted to predict the subsequent popularity of Sinhala language YouTube videos by analysing a number of associated factors.
+Ilangeshwaran et al.
+[272] introduced Saliency-based token Swap (SSwap), a data augmentation technique to improve Sinhala text classification.
+The work by Rathnayake and Asanka [231] uses LLaMA for Named entity classification in Sinhala documents and then runs Fuzzy C-Means clustering on it.
+Next, they utilise LLaMA again to propose names to the created clusters.
 3.11 Sentiment Analysis
 A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [273] based on their prior work [274].
-A word2vec based tool116116116http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
+A word2vec based tool118118118http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
 A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al.
 [275]Demotte et al.
-[272] proposed a sentiment analysis system117117117https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments.
+[276] proposed a sentiment analysis system119119119https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments.
 In the subsequent work [277], they used word similarity to generate a Sinhala semantic lexicon.
@@
 [246]Aththanayaka and Naleer [287] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media.
-The ACTSEA dataset118118118https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
+The ACTSEA dataset120120120https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
 [135].
@@
 Fernando and Ranathunga [292] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations.
+The work by Peiris and Sumanathilaka [293] analyses emotions in Sinhala songs both in the directions of lyrics and audio.
+In their analysis of lyrics, they use RoBERTa to represent the Sinhala words.
+Rizvi et al.
+[294] used XLM-RoBERTa to extract keywords and classify aspects in Sinhala Code-Mixed data.
+Then in their follow up work [295] they again used XLM-RoBERTa, this time for sentiment analysis and, on top of that used SHAP [296] and LIME [240] to make the model explainable.
 3.12 Hate Speech Detection
@@
 A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [298].
-A feature model and a data set119119119https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al.
+A feature model and a data set121121121https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al.
 [299] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques.
@@
 [129] analyses Sinhala hate speech propagation on Twitter.
-They have also published a dataset120120120https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
+They have also published a dataset122122122https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
 Rajapaksha et al.
@@
 They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
-Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook121121121https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter122122122https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets.
+Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook123123123https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter124124124https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets.
 They report TwHIN-BERT-base [318] model to have obtained the best results.
+The work by Jahnavi and Chaturvedi [319] used BERT with CNN and LSTM to detect hate speech in Sinhala.
+Perera et al.
+[320] uses XLM-RoBERTa to classify Sinhala youtube video titles and comments to severity levels of hate.
+This, by definition is a multi-class classification.
+The work by Tanjim [321] analyses multiple low-resourced languages in the context of Cyberbullying detection.
+Ultimately, for the case of Sinhala as the target language, they recommend using Assamese, Bengali, Nepali, Urdu, Hindi, or Bodo as the source language for transfer learning this task.
+, title=Multilingual Cyberbullying Detection in Low-Resource Languages and Dialects
 After the release of the SOLD dataset by Ranasinghe et al.
@@
 The work by Bestgen [324] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali.
+Dmonte et al.
+[325] used the SOLD data set to determine whether machine translation impacts the performance of hate speech detection.
+Haturusinghe et al.
+[326] introduces a series of models named Subasa that are tested on the SOLD dataset.
+Their trained PMLs125125125https://github.com/haturusinghe/subasa-plm and LLMs126126126https://github.com/haturusinghe/subasa-llm are separately publicly available.
+Ou6t of these models, they report that their Subasa-XLM-R model outperforms even the state-of-the-art commercial LLMs when evaluated on the SOLD data set.
 Ranasinghe et al.
-[314] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task123123123https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
+[327] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task127127127https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
 In completing the task, the study by Narayan et al.
@@
 Dmonte et al.
-[141] used the SOLD data set to verify the new MT-Offense124124124https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
+[141] used the SOLD data set to verify the new MT-Offense128128128https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
 They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].
@@
 [192] have proposed a system named Aruth based on the Lesk Algorithm[340].
-An online tool125125125http://aruth.herokuapp.com/, an API126126126https://bit.ly/3sJEYbS of the algorithm, and code along with data on github127127127https://github.com/jseanm1/aruthSWSD are available.
+An online tool129129129http://aruth.herokuapp.com/, an API130130130https://bit.ly/3sJEYbS of the algorithm, and code along with data on github131131131https://github.com/jseanm1/aruthSWSD are available.
 For the same task, Marasinghe et al.
@@
 Hasan et al.
-[151] introduced the XL-sum128128128https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website129129129https://www.bbc.com/sinhala.
-The M3LS130130130https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
-[152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website131131131https://www.bbc.com/sinhala.
+[151] introduced the XL-sum132132132https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website133133133https://www.bbc.com/sinhala.
+The M3LS134134134https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
+[152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website135135135https://www.bbc.com/sinhala.
 Further, they claim that MT5 [349] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
-The first Sinhala multi-document summarizing dataset, M2DS132132132https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
+The first Sinhala multi-document summarizing dataset, M2DS136136136https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
 [153].
 The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary.
-Thier code and pre-trained models are also available on github133133133https://github.com/KushanMH/m2ds.
+Thier code and pre-trained models are also available on github137137137https://github.com/KushanMH/m2ds.
 3.16 Other Semantic Tools
@@
 [352] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems.
-The Massively Multilingual Speech (MMS)134134134https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
+The Massively Multilingual Speech (MMS)138138138https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
 [145] has Sinhala data for the Spoken Language Identification (LID) task.
@@
 TacoSi introduced by Arachchige and Weerasinghe [405] is based on Tacotron [406] and has been evaluated with 10 human evaluators to determine its text-to-speech quality.
-Nanayakkara [394], Nanayakkara and Weerasinghe [395] used DeepSpeech135135135https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
+Nanayakkara [407], Nanayakkara and Weerasinghe [408] used DeepSpeech139139139https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
 Gunarathne et al.
-[396] used an earlier version of the CMUSphinx toolkit136136136https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
+[409] used an earlier version of the CMUSphinx toolkit140140140https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
 The later work by Akesh and Meegama [410] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [411, 412] to automatically generate Sinhala subtitles from Speech.
@@
 [414] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking.
-tha [402] have introduced a Sinhala text to Speech dataset137137137https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes.
+tha [415] have introduced a Sinhala text to Speech dataset141141141https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes.
 As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [416].
-While the data set link points to a github repository, it does not contain the data set; it points to a project website138138138https://sita-speakerdiarization.github.io/ which then points to a download page139139139https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
+While the data set link points to a github repository, it does not contain the data set; it points to a project website142142142https://sita-speakerdiarization.github.io/ which then points to a download page143143143https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
 3.17.3 Speech-to-Speech
@@
 3.17.5 Speech classification
+Luger et al.
+[423] introduced a data set named unsupervised_peoples_speech144144144https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data.
 The Sinhala speech classification system proposed by Buddhika et al.
@@
 [426] is capable of handling Sinhala voice commands.
+Kathriarachchi and Pulasinghe [427] developed a system to classify and predict Sinhala speech clips of children into age groups.
 The work by Welarathna et al.
@@
 3.18.1 Printed Text
-The XTREME-UP140140140https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP145145145https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [144] contains a Sinhala data set for the OCR task.
@@
 Anuradha et al.
-[443] used Tesseract 3141141141https://tesseract-ocr.github.io/ [444] for Sinhala OCR.
+[458] used Tesseract 3146146146https://tesseract-ocr.github.io/ [459] for Sinhala OCR.
 The later study by [460] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.
@@
 [469] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka.
-However, the paper provides no URLs and we were not able to find the data set on the NSF website either.
+However, the paper provides no URLs, and we were not able to find the dataset on the NSF website either.
+Weraduwa et al.
+[470] claims to have created a dataset of 373 Sinhala handwriting images from 84 participants where 73 are dysgraphic and 300 are non-dysgraphic.
+Their objective isthat this dataset may be used to detect dysgraphia in children.
 The work by Karunanayaka et al.
@@
 The study by Rowel et al.
-[470] frames their work as an E-Learning platform for hearing impaired children.
+[486] frames their work as an E-Learning platform for hearing-impaired children.
 However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23.
-What they do have is an OCR system that they claim to recognize letters and digits.
+What they do have is an OCR system that they claim to recognise letters and digits.
 Even there, we are only given an example of a recognised digit.
-Whether or not their system can recognize Sinhala letters is not explicitly discussed.
+Whether or not their system can recognise Sinhala letters is not explicitly discussed.
 As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al.
@@
 Karunarathne et al.
-[473] used Gabor-initialized CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
+[489] used Gabor-initialised CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
 3.18.3 Ancient Text
@@
 This study used academic articles as its domain and ultimately concluded that both translation systems need significant improvements before they can be recommended for widespread usage.
+The survey by Tafa et al.
+[510] compared the performance of a number of MT models on Low-Resource languages, including Sinhala.
 As mentioned in Section 3.1, the study by [102] raised questions on the quality of the existing Sinhala-English corpora.
@@
 In their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking, Fernando and Ranathunga [292] used SiTa and EnSi NMT on NLLB and CCAligned as one of the evaluations.
+Thillainathan et al.
+[543] proposed to use continual pre-training and intermediate task transfer learning for Fine-tuning what they call multilingual sequence-to-sequence large language models (msLLMs).
+However, in actuality, what their experiments are on is mBART.
+Fernando et al.
+[544] in their meta-study on low-resource language NMT, introduced debiasing heuristics for web mined parallel corpora that improve EnSi translation quality.
+They showed the improvements on CCMatrix and CCAligned datasets.
+The meta study by Velayuthan et al.
+[545] proposes an improved sequence level Knowledge Distillation method that improves encoder knowledge transfer.
+They used Sinhala-English NMT to discuss the $\alpha$ value ablation by using CCAligned [546], OpenSubtitles [133], and SITA (Gov) [524] corpora.
 Sinhala is one of the languages considered in the short survey by Perera and Sumanathilaka [547] on machine translation and transliteration for Indo-Aryan Languages.
@@
 3.19.3 Singlish to Sinhala (Transliteration)
-The XTREME-UP142142142https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP147147147https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [144] contains a Sinhala data set for the transliteration task.
@@
 They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole.
-Their code is available on GitHub143143143https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
+Their code is available on GitHub148148148https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
 Khiu et al.
@@
 They reported across-the-board better performance with the latter.
-All thier code as well as the test-sets are publicly avaialble144144144https://github.com/kasunw22/Sinhala-Transliterator/.
+All thier code as well as the test-sets are publicly avaialble149149149https://github.com/kasunw22/Sinhala-Transliterator/.
 Perera et al.
 [575] attempted the same task with BERT.
-Their code is also publicly available145145145https://github.com/Sameera2001Perera/Singlish-Transliterator.
+Their code is also publicly available150150150https://github.com/Sameera2001Perera/Singlish-Transliterator.
 In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [547] covers a number of works [521, 573, 552, 553, 555, 557, 556, 559, 562, 563] on Sinhala transliteration which we have already discussed above.
@@
 [610].
-A rule-based Sinhala spell checker named SinSpell based on Hunspell146146146http://hunspell.github.io/ was introduced by Liyanapathirana et al.
+A rule-based Sinhala spell checker named SinSpell based on Hunspell151151151http://hunspell.github.io/ was introduced by Liyanapathirana et al.
 [611].
-They have also made the tool available147147147http://nlp-tools.uom.lk/sinspell/ online for use.
+They have also made the tool available152152152http://nlp-tools.uom.lk/sinspell/ online for use.
 The study by Sithamparanathan and Uthayasanker [612] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil.
@@
 A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al.
-[604] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA148148148https://rasa.com/ [605] chat bot.
+[625] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA153153153https://rasa.com/ [626] chat bot.
 A Sinhala chat bot for train information was proposed by Harshani [627].
@@
 The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
+The end product of work by Dasanayaka et al.
+[634] is a Sinhala chatbot that can answer patient questions on orthopantomography radiology reports which were previously generated by the system itself by analysing dental panoramic tomography (DPT) images.
 3.22 News and/or Social Media Recommendation
@@
 Meyler [641] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language.
-Further, [620] claim that SSL was derived from British Sign Language (BSL).
+Further, Rishan et al.
+[642] claim that SSL was derived from British Sign Language (BSL).
 However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL.
@@
 The dataset contains 1110 videos.
+Abhishek and Sumanathilaka [653] claims to introduce SSL400, a word-level Sinhala sign language dataset consisting of 384 commonly used words.
+However, given that their paper is behind a paywall, it is not possible to verify whether the dataset is publicly available or not.
 Priyankara et al.
@@
 The study by Haputhanthri et al.
-[637] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language.
+[660] used ResNet followed by a transformer encoder with multi-head attention to recognise sentence-level Sinhala sign language.
 The study by Fernando et al.
@@
 They reported a success rate of 86.7%.
-Strides have been made in the domains of computer interpreting for written Sinhala [645] and animation of finger-spelled words and number signs [646].
+Rajapakshe et al.
+[668] used a hybrid CNN-LSTM model and reports 94% accuracy on the 20 Sinhala signs they tested on.
+The work by Navodya et al.
+[669] reports accuracies between 87% and 92% for recognising Sinhala sign language.
+Thennakon et al.
+[670] on the other hand reports 96% accuracy in recognising real-time Sinhala Sign Language signs.
+The work by Kavinda et al.
+[671] proposed a multi-modal system that converts both sign language and miming language to Sinhala text.
+Further, they have a component to convert Sinhala audio to miming language and Sinhala text as well.
+Strides have been made in the domains of computer interpreting for written Sinhala [672] and animation of finger-spelt words and number signs [673].
+The survey by Ahinsa et al.
+[674] discusses text-to-Sinhala sign language in the contexts of: finger-spelling and translation.
+It should be noted here that the translation they describe here is text-to-sign language translation and not translation between any two or more natural languages.
 The work by Idushan et al.
@@
 Kumar et al.
-[648] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language.
-The work by dha [627] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs.
+[676] introduced EasyTalk, a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language.
+The work by dha [649] used CNN to classify features of static Sinhala signs, while dynamic signs were modelled as a tree structure of static signs.
 Dahanayaka et al.
@@
 Rishan et al.
-[620] uses Leap Motion Controller149149149https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
+[642] uses Leap Motion Controller154154154https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
 Dissanayake et al.
@@
 [682] as well as the study by Weerasinghe [683] also focused on creating an automated Braille to Sinhala recognition system.
+Dampalessa and Ragel [684] used a data set of 230 characters to train a model to transliterate Braille letters into Sinhala text.
 Anuradha and Thelijjagoda [376] uses OCR on Sinhala and English Braille documents on which they then run a machine translation system in order to convert them into voice.
+Udana et al.
+[685] introduces a prototype tactile Braille display for Sinhala.
+They claim to be able to convert digital Sinhala text into Braille using an electromechanical process that also utilises a CNN.
 3.25 Plagiarism Detection
@@
 However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
+Senanayaka et al.
+[700] proposes SingRAG, a RAG and translation framework built on LLaMA that converts code-mixed Sinhala to English for processing and then converts whatever the English results is of the attached backend to code-mixed Sinhala before sending it back to the user.
 3.27 Language Acquisition and Evaluation
@@
 They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system.
+Weerasinghe et al.
+[707] introduced LexiLearn, a gamified system intended to help Sinhala learners with dyslexia.
+The survey by Perera and Sumanathilaka [708] discusses the lack of annotated datasets and NLP resources to cater for Sinhala-speaking adults with dyslexia.
 3.28 Embeddings, LLMs, and Other Language Models
@@
 Dhananjaya et al.
-[107] has published two Sinhala BERT models: SinBERT-large150150150https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small151151151https://huggingface.co/NLPC-UOM/SinBERT-small.
-A smaller BERT model named SinhalaBERTo152152152https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
+[107] has published two Sinhala BERT models: SinBERT-large155155155https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small156156156https://huggingface.co/NLPC-UOM/SinBERT-small.
+A smaller BERT model named SinhalaBERTo157157157https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
 This BERT model has been created by training vanilla RoBERTa [710] on the OSCAR [155, 156] dataset.
 Demotte et al.
-[272] have released the word embeddings153153153https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code154154154https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer.
+[276] have released the word embeddings158158158https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code159159159https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer.
 Gurgurov et al.
@@
 [276] as their source for Sinhala embeddings.
+Gurgurov et al.
+[712]160160160https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs.
 Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment.
 Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
-The relevant code and data can be accessed on GitHub155155155https://bit.ly/3t3SKu7.
+The relevant code and data can be accessed on GitHub161161161https://bit.ly/3t3SKu7.
 Subsequent work by Rathnayake et al.
@@
 Üstün et al.
-[171] introduced Aya, a massively multilingual generative language model156156156https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159].
+[171] introduced Aya, a massively multilingual generative language model162162162https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159].
 Jayakody and Dias [716] conducted a comparative analysis on LLMs for Sinhala.
-In it, they observed that the small (7̃B) models of Claude157157157https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o158158158https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)159159159https://llama.meta.com/docs/get-started/ and Mistral 7B160160160https://docs.mistral.ai/.
-Further, they reported that they had to exclude Falcon 2161161161https://falconllm.tii.ae/ and Gemini162162162https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+In it, they observed that the small (7̃B) models of Claude163163163https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o164164164https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)165165165https://llama.meta.com/docs/get-started/ and Mistral 7B166166166https://docs.mistral.ai/.
+Further, they reported that they had to exclude Falcon 2167167167https://falconllm.tii.ae/ and Gemini168168168https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+Hansaka and Sumanathilaka [717] also claims to have conducted a review on fine-tuning approaches for LLMs such as GPT-3, GPT-4, and LlaMA in the context of Sinhala.
 Chang et al.
-[683] introduced Goldfish, a collection of monolingual language models163163163https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109].
+[718] introduced Goldfish, a collection of monolingual language models169169169https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109].
 The work by Ayesha et al.
 [719] used a fined-tuned Llama 3.1 (8B) model [720] to predict student performance by analyzing textual entries created by students in Sinhala and English.
+Tang et al.
+[721] evaluated the susceptibility of LLMs being tricked into answering illegal questions in low-resource languages when the same questions would have been successfully rejected had they been asked in a high-resourced language.
+They provide the framework for testing with the results on eight low-resourced languages, which include Sinhala.
 Vayani et al.
@@
 This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category.
-This dataset, which they named ALM-Bench164164164https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available.
+This dataset, which they named ALM-Bench170170170https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available.
 They made a number of observations in this aspect.
@@
 3.29 Text Simplification
+Ranathunga et al.
+[723] introduced SiTSE, a Sinhala text simplification dataset and the associated benchmark.
+They claim that the data set is publicly available on github171171171https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation; however, currently it resolves to a 404.
 The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [724] included the NSINA [108] Sinhala corpus.
-The aggregated data from the shared task is available online as Train-Test data165165165https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data166166166https://huggingface.co/datasets/MLSP2024/MLSP2024.
+The aggregated data from the shared task is available online as Train-Test data172172172https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data173173173https://huggingface.co/datasets/MLSP2024/MLSP2024.
 Enomoto et al.
@@
 The work by Velayuthan and Sarveswaran [732] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala.
-Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available167167167https://github.com/Buddhilive/sltk.
+Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available174174174https://github.com/Buddhilive/sltk.
 However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words.
-While the authors have made the dictionary publicly available168168168https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries.
-So while it may be adequate to handle Sri Lanka-centric news title data set169169169https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
+While the authors have made the dictionary publicly available175175175https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries.
+So while it may be adequate to handle Sri Lanka-centric news title data set176176176https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
 3.31 Detection or Management of Mental Illnesses
@@
 [737] used a NAO Robot [738] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children.
+The system by Kulasooriya et al.
+[739] used SVM and simple ANN to detect psychological stress in Sinhala speech.
 3.32 Sinhala Text Generation
@@
 [748] used pre-trained mBART [541] and mT5 [349, 730] models for the same task and showed better results.
-3.33 Miscellaneous Applications
-In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them.
-Dissanayake and Hettige [711] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs.
+3.33 Question Answering Systems (Q&A)
+Dissanayake and Hettige [749] implemented a question-and-answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs.
 Jayakody et al.
@@
 [161] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala.
-They have publicly released both their data170170170https://huggingface.co/datasets/afaji/cvqa and leader board171171171https://eval.ai/web/challenges/challenge-page/2305/overview.
+They have publicly released both their data177177177https://huggingface.co/datasets/afaji/cvqa and leader board178178178https://eval.ai/web/challenges/challenge-page/2305/overview.
 Vayani et al.
-[164] created a dataset ALM-Bench172172172https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web.
+[164] created a dataset ALM-Bench179179179https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web.
 They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [166] is superior to other tested LLMs.
 Wansekara and Jayasekara [751] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text.
+Dhanawardhana et al.
+[752] frame their work just as a multi-model deep learning approach.
+But in actuality, their lost-and-found application is utilising techniques that are relevant to Sinhala VQA.
+The system created by Dasanayaka et al.
+[634] is claimed to be capable of first, generating radiology reports for orthopantomography (related to dentistry) using Llama on captions generated by a Blip-2-based caption generator on dental panoramic tomography (DPT) images and then serve a Sinhala chatbot where the patients may ask questions on the said report and get answers.
+3.34 Language Identification (LID) Task
+Burchell et al.
+[146] show that they obtain better results for the LID task for most of the languages compared to NLLB [134].
+However, for Sinhala the results by Burchell et al.
+[146] are the same as that of NLLB.
+Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model.
+However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%.
+Kargaran et al.
+[753] introduces a tool named GlotScript which is used for written LangID it achieves above 80% accuracy for Sinhala.
+3.35 Miscellaneous Applications
+In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them.
 Fernando [754] proposed a method for inexact matching of Sinhala proper names.
@@
 [767] have created a small Sinhala treebank of 100 sentences.
-Burchell et al.
-[146] show that they obtain better results for the LID task for most of the languages compared to NLLB [134].
-However, for Sinhala the results by Burchell et al.
-[146] are the same as that of NLLB.
-Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model.
-However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%.
 Minixhofer et al.
@@
 used NER to cluster Sinhala news articles.
-Subasinghe and Samarawickrama [733] analysed the Noto Sans Sinhala173173173https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens.
+Subasinghe and Samarawickrama [773] analysed the Noto Sans Sinhala180180180https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens.
 Zhao et al.
@@
 In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [181, 177, 180], Sinhala-Sinhala [176, 179], and English-Sinhala-Tamil [178].
-A number of NLP applications have utilized first sources intended to teach children [750, 751, 752, 753, 754] or foreigners [755, 220, 756]174174174Note that [756] is an extension of [220]..
+A number of NLP applications have utilized first sources intended to teach children [790, 791, 792, 793, 794] or foreigners [795, 221, 796]181181181Note that [796] is an extension of [221]..
 This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@
 Figure 10: The Probability of research from an institution citing that of other institutions.
-Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext176176176https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
+Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext183183183https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
 With all those limitations in mind, we still can make a few Interesting observations.

### v26

@@
 As such, we shall be uploading this paper to arXiv and perpetually update it periodically as a living research article to reflect the advances made in the field.
-This manuscript is at version $5.7.0$.
+This manuscript is at version $5.7.1$.
 Index Terms:
-Sinhala, Natural Language Processing, Resource Poor Language
+Sinhala, Natural Language Processing, Low-Resource Languages
 1 Introduction
-Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 16 million people [5, 6].
+Sinhala [^1] Englebretson and Genetti [1] observe that in some contexts the Sinhala language is also referred as Sinhalese, Singhala, and Singhalese language, being the native language of the Sinhalese people [2, 3, 4], who make up the largest ethnic group of the island country of Sri Lanka, enjoys being reported as the mother tongue (L1) of approximately 17 million people [5, 6].
 When both L1 and L2 speakers are counted, $79.7\%$ of the total Sri Lankan population are literate in Sinhala [7].
@@
 This implies two things: 1) the majority of the Sinhala linguistic sources are located in the geographical area of Sri Lanka, and 2) the majority of the available linguistic sources in Sri Lanka are in Sinhala.
+Sri Lanka is the only country where Sinhala is recognised as an official language [9].
 To give a brief linguistic background for the purpose of aligning the Sinhala language with the baseline of English, primarily it should be noted that the Sinhala language belongs to the same Indo-European language tree [10, 11, 12].
@@
 We show the actual evolutionary eras of the Sinhala script in Appendix A where the gradual change from Brahmi to the contemporary script is clearer.
-It should be noted that the modern Sinhala language has loanwords from languages such as Tamil, English, Portuguese, and Dutch due to various historical reasons [22].
+It should be noted that, due to various historical reasons [23], the modern Sinhala language has loanwords from languages such as Tamil [24], English, Portuguese, and Dutch.
 An analysis of the distribution of Sinhala letters towards the beginning, middle, or end of words conducted in the style of proposed by ProoffReader [25] using the SinMin Sinhala corpus created by Upeksha et al.
@@
 Only the Elu sub-tree is shown exhaustively.
-All other sub-trees only list some languages for representation purposes chosen randomly or otherwise.
+All other sub-trees only list some languages for representation purposes, chosen randomly or otherwise.
 Some of the languages shown as Extinct have archaeological proof (eg: Prakrit) while others were retroactively reconstructed by linguists by observing the commonalities of languages that they begot (eg: Proto-Indo-Aryan).
-The status of some languages has spread beyond linguistics to be a social and political question.
+The status of some languages have spread beyond linguistics to be a social and political question.
 In this depiction, we have attempted to side with academic consensus when available and the majority when not.
-As an example of one contentions point on this diagram note the positioning of Pali.
-While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
+As an example of one contentions point on this diagram, note the positioning of Pali.
+While what we show here is the academically accepted relationship, there are other schools of thought: 1) Some argue that Pali is a constructed language (ConLang) of Sinhala and thus should be depicted as a child of Sinhala., 2) Others claim it to be a sibling of Sinhala, 3) Further, some claim it to be either be the link between Elu and Sinhala or the parent of Elu, 4) Some claim it needs to be merged with Gandhari, and 5) Some claim it to be the same as Magadhi Prakrit given that in some Buddhist traditions, it is noted Buddha to have spoken Magadhi.
 Figure 2: A portion of the writing systems (scripts) tree with a focus on Sinhala and its close relatives.
 It is important to note that this is distinct from the language tree shown in Fig 1.
-The writing systems used by a language being related, not being related, or the degree to which they are related is not directly correlated to the relevant languages being related, not being related, or the degree to which they are related.
+The writing systems used by a language being related, not being related, or the degree to which they are related, is not directly correlated to the said languages being related, not being related, or the degree to which they are related.
 As an example, note that the writing systems of Sinhala and Tamil are reasonably closely related.
@@
 A language that is shown as a single entity in Fig 1 may have multiple entries in this diagram.
-An example is the language shown as Hindustani splits into Hindi and Urdu where the former uses the Devanagari script and the latter a derivation of Arabic script.
-Conversely, languages that may be shown as distinct entities in the full language tree, may map to a singular entity in the script tree.
-An example is the fact that English which would go under Proto-Germanic and Italian which would go under Proto-Italic in and extended version of Fig 1, would both map to Latin in this diagram given that both of them use the Latin script.
+An example is the language shown as Hindustani in Fig 1 splits into Hindi and Urdu in this diagram, where the former uses the Devanagari script and the latter a derivation of Arabic script.
+Conversely, languages that may be shown as distinct entities in Fig 1, may map to a singular entity in the script tree.
+An example is English which would go under Proto-Germanic and Italian which would go under Proto-Italic in an extended version of Fig 1, would both map to Latin in this diagram, given that both of them use the Latin script.
 Regardless of the rich historical array of literature spanning several millennia (starting between $3^{rd}$ to $2^{nd}$ century BCE [28, 29]), modern natural language processing tools for the Sinhala language are scarce [30, 31].
@@
 The morphological layer analyses words at their smallest units of meaning.
-As such, analysis on word lemmas and prefix-suffix-based inflection are handled in this layer.
+As such, analysis of word lemmas and prefix-suffix-based inflection is handled in this layer.
 Lexical layer handles individual words.
-Therefore tasks such as Part of Speech (PoS) tagging happens here.
-The next layer, syntactic, takes place at the phrase and sentence level where grammatical structures are utilized to obtain meaning.
+Therefore, tasks such as Part of Speech (PoS) tagging happens here.
+The next layer, syntactic, takes place at the phrase and sentence level, where grammatical structures are utilised to obtain meaning.
 Semantic layer attempts to derive the meanings from the word level to the sentence level.
-Starting with Named Entity Recognition (NER) at the word level and working its way up by identifying the contexts they are set in until arriving at overall meaning.
+Starting with Named Entity Recognition (NER) at the word level and working its way up by identifying the contexts they are set in until arriving at the overall meaning.
 The discourse layer handles meaning in textual units larger than a sentence.
-In this, the function of a particular sentence maybe contextualized within the document it is set in.
+In this, the function of a particular sentence maybe contextualised within the document it is set in.
 Finally, the pragmatic layer handles contexts read into contents without having to be explicitly mentioned [37, 38].
@@
 [37].
-However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [45, 46] which continuously gets updated as new research and tools for Sinhala language are created and made publicly available.
+However, our hope is to host this survey at arXiv as a perpetually evolving, living research article [47, 48] which continuously gets updated as new research and tools for the Sinhala language are created and made publicly available.
 We also discuss how the non-compliance of policies to put data and code online [49], after the research is concluded and the paper is published, has negatively impacted the growth and sustainability of Sinhala NLP.
@@
 For the success of this survey, we shall also consider the Sri Lankan NLP tools repository, lknlp [^4] https://github.com/lknlp/lknlp.github.io.
-This manuscript is at version $5.7.0$.
+This manuscript is at version $5.7.1$.
 The latest version of the manuscript can be obtained from arXiv [^5] https://arxiv.org/abs/1906.02358 or ResearchGate [^6] http://bit.ly/31AhvvR.
@@
 An example of an updated reference is given in Code Block LABEL:lst:ref.
-@article{de2025survey,
+@article{de2026survey,
 title={{Survey on Publicly Available Sinhala Natural Language Processing Tools and Research}},
 author={de Silva, Nisansa},
-journal={arXiv preprint arXiv:1906.02358v25},
-year={2025}
+journal={arXiv preprint arXiv:1906.02358v26},
+year={2026}
 Code Block 1: Example of an updated reference
@@
 In this section, we would discuss both pure Sinhala NLP tools and research as well as hybrid Sinhala-English work.
-We will also discuss research and tools which contributes to Sinhala NLP either along with or by the help of Tamil, the other official language of Sri Lanka.
+We will also discuss research and tools that contribute to Sinhala NLP, either along with or with the help of Tamil, the other official language of Sri Lanka.
 Section 4 gives a brief introduction to the primary language sources used by the studies discussed in this work.
@@
 Sinhala grammar has two forms: written (literary) and spoken.
-These forms differ from each-other in their core grammatical structures [49, 1, 50].
+These forms differ from each-other in their core grammatical structures [32, 51, 1, 52].
 The written form strictly adheres to the SOV (Subject, Object, and Verb) configuration [53, 54].
@@
 Sinhala is also a head-final language, where the complements and modifiers would appear before their heads [57] this is similar to that of English and dissimilar to that of French.
-In total, according to Abhayasinghe [56], there are 25 types of simple sentence structures in Sinhala.
+In total, according to Abhayasinghe [58], there are 25 types of simple sentence structures in Sinhala and according to Madushani [59] there are 8 types of Verb Phrases.
 Similar to many Indo Aryan languages, animacy plays a major role in Sinhala grammar in syntactic and semantic roles [60, 61, 62].
@@
 [66].
-(a) English
-(b) Sinhala
-(c) Hindi
-(d) French
+((a)) English
+((b)) Sinhala
+((c)) Hindi
+((d)) French
 Figure 4: Parse trees for the sentence “I eat a red apple” in four languages.
@@
 [111].
-ImaniGooghari et al.
+Imani et al.
 [112] published Glot500272727https://github.com/cisnlp/Glot500, a corpus of 500 languages that includes Sinhala.
@@
 In the same paper, they also introduce the dataset SemiSOLD [^68] https://huggingface.co/datasets/sinhala-nlp/SemiSOLD which contains 145,000 Sinhala tweets annotated with the same classes but with a semi-supervised approach.
-The relevant code is also available on Github 696969%****␣03_Sinhala.tex␣Line␣75␣****https://github.com/Sinhala-NLP/SOLD.
+The relevant code is also available on Github  [^69] https://github.com/Sinhala-NLP/SOLD.
 The subsequent work by Dmonte et al.
@@
 Ranathunga et al.
-[167] has created a text simplification dataset named SiTSE [^96] https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation.
-It has 1000 complex Sinhala texts along with two separate simplifications for each of them.
-Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data.
-Ranathunga et al.
 [170] created multiNER [^96] https://github.com/suralk/multiNER, a parallel NER data set for Sinhala, English, and Tamil.
@@
 Chavinda and Thayasivam [174] introduces two publicly available Sinhala hate speech data sets.
-One is from Facebook [^99] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter100100100https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech.
-A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released101101101https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [173].
+One is from Facebook [^98] https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and the other is from Twitter [^99] https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech.
+A large data set of 63,471 Sinhala youtube comments collected from Sinhala videos was released100100100https://github.com/Yomald93/Linguistic-Analysis-of-Sinhala-YouTube-Comments-on-Sinhala-Music-Videos by De Mel and de Silva [175].
 One important point is that while 35,428 of them are original Sinhala comments collected as was, 28,043 are either transliterated from Singlish.
+By further manually tagging this dataset, De Mel and de Silva [176] introduced GeeSanBhava101101101https://bit.ly/SinhalaYoutubeComments, the first Sinhala sentiment analysis dataset tagged using Russell’s Valence-Arousal model [177] as opposed to arbitrary sentiment classes.
 Burchell et al.
@@
 They have also publicly released their translation models103103103https://huggingface.co/collections/HPLT/hplt-20-uni-direction-translation-models-67f2fc7ae54845f9b182957a.
+Chang et al.
+[179] created Global PIQA104104104https://mrlbenchmarks.github.io/, a multilingual extension to PIQA [180] in which they include physical commonsense reasoning triples covering 100+ languages such that the examples are culturally aligned to the relevant languages.
+The objective of this dataset is to function as a benchmark to measure the cultural alignment of LLMs.
+Senaratna [181] has created and perpetually maintains Sri Lanka Document Datasets in which he curates data sets which include Sinhala text such as: Hansards105105105https://github.com/nuuuwan/lk_hansard/, Police Press Releases106106106https://github.com/nuuuwan/lk_police_press_releases/, Legal Documents107107107https://github.com/nuuuwan/lk_legal_docs/, Cabinet Decisions108108108https://github.com/nuuuwan/lk_cabinet_decisions/, Treasury Press Releases109109109https://github.com/nuuuwan/lk_treasury/, Sri Lanka Presidential Media Division press releases110110110https://github.com/nuuuwan/lk_pmd/, News111111111https://github.com/nuuuwan/lk_news/, and Educational Publications from the Educational Publications Department112112112https://github.com/nuuuwan/lk_edupub/.
+A collection of Sinhala-Sinhala, Sinhala-English, English-Sinhala, and Pali-Sinhala dictionaries is available on GitHub113113113https://github.com/pnfo/arutha.lk/tree/main/dev/dict-input courtesy of aruth.lk.
+Another Pali-Sinhala dictionary is available on Github114114114https://github.com/pnfo/pali-sinhala-dictionary by the Path Nirvana project.
+The same project provides a large repository of Buddhist books with Sinhala and Pali text 115115115https://github.com/pathnirvana/pitaka.lk/tree/master/books116116116https://github.com/pathnirvana/sinhala-tools/tree/master/dev/buddhist-books/text.
+Further, they have some News data collected from Silumina117117117https://github.com/pathnirvana/sinhala-tools/tree/master/dev/silumina, Mawbima118118118https://github.com/pathnirvana/sinhala-tools/tree/master/dev/mawbima, and AdaDerana119119119https://github.com/pathnirvana/sinhala-tools/tree/master/dev/adaderana.
+An associated profile contains a large Sinhala TTS data set120120120https://github.com/pnfo/sinhala-tts-dataset consisting of 6248 sentences with 13.8 hours of recordings over two speakers: one male and one female.
+A wiki version of the Sinhala Encyclopedia maintained by the Sri Lankan government is available online121121121https://t.co/k8gFT9puM0.
+It seems it has not been updated recently.
+However, some of its articles are of higher quality than the same of Sinhala Wikipedia122122122https://si.wikipedia.org/.
+Google has made available a comprehensive IPA conversion (pronunciation) dictionary of a large Sinhala vocabulary123123123https://raw.githubusercontent.com/google/language-resources/master/si/data/lexicon.tsv.
+A PoS tagged data set of 250,000+ Sinhala words is available through Google Groups124124124https://t.co/jpQz74YEdQ.
+A rudimentary Sinhala Tokenizer, Stemmer, and a PoSTagger have been published on GitHub125125125https://github.com/ysenarath/sinling by the user ysenarath.
+While the research paper by Ekanayaka et al.
+[182] seems to be lost, the news corpus created for that study, which contains 1000 labelled and 50,000 unlabelled Sinhala news reports collected from various sources, has been published to github126126126https://github.com/rksk/sinhala-news-analysis/tree/master/sinhala-news-corpus.
+The repository also contains some code developed for the relevant work 127127127https://github.com/rksk/sinhala-news-analysis/.
 3.3 Dictionaries
@@
 The study by Wickramasinghe and De Silva [150] introduced a large Sinhala-English dictionary which has $1,368,416$ unfiltered and $195,255$ filtered En-Si pairs.
-Both their code and the two versions of the dictionary are publicly available on github104104104https://github.com/kasunw22/sinhala-para-dict/tree/main.
-A Sinhala-Sinhala dictionary named Wahara105105105http://crawler.nlpc.uom.lk/ was built and is publicly maintained by the University of Moratuwa.
+Both their code and the two versions of the dictionary are publicly available on github128128128https://github.com/kasunw22/sinhala-para-dict/tree/main.
+A Sinhala-Sinhala dictionary named Wahara129129129http://crawler.nlpc.uom.lk/ was built by Iniyage et al.
+[193].
+It is publicly maintained by the University of Moratuwa.
 There exists the government-sponsored trilingual dictionary [194], which matches Sinhala, English, and Tamil.
@@
 For a time it was hosted on [200] but it too is now defunct and all the data and applications are lost other than what Arukgoda et al.
-[192] have cloned to use in their application uploaded on github106106106https://github.com/jseanm1/aruthSWSD.
+[201] have cloned to use in their application uploaded on github130130130https://github.com/jseanm1/aruthSWSD.
 However, even at its peak, due to the lack of volunteers for the crowd-sourced methodology of populating the WordNet, it was at best an incomplete product.
@@
 Further, this group presented a comparison of few Sinhala PoS taggers that are available to them [221].
-A RESTFul PoS tagging web service created by Jayaweera and Dias [213] using the above research can still be accessed107107107http://bit.ly/2F0jKid via POST and GET.
+A RESTFul PoS tagging web service created by Jayaweera and Dias [222] using the above research can still be accessed131131131http://bit.ly/2F0jKid via POST and GET.
 A hybrid PoS tagger for Sinhala language was proposed by Gunasekara et al.
@@
 The Ananya system by Manamini et al.
-[118] is available to download at GitHub 108108108http://bit.ly/2XrwCoK.
+[121] is available to download at GitHub 132132132http://bit.ly/2XrwCoK.
 The data and code for the approaches by Dahanayaka and Weerasinghe [232] and by Senevirathne et al.
@@
 They have used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA).
-A data set for Sinhala NER named multiNER109109109https://github.com/suralk/multiNER was introduced by Ranathunga et al.
+A data set for Sinhala NER named multiNER133133133https://github.com/suralk/multiNER was introduced by Ranathunga et al.
 [170].
@@
 Rathnayake and Asanka [240] simply uses LLaMA for Sinhala NER.
+Sinhala NER is discussed along with other Indic languages in the survey by Deshmukh et al.
+[241].
 3.9 Semantic Similarity
@@
 Data and tools for these projects are not publicly available.
-A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [234] where they demonstrate their results for Sinhala and Tamil.
+A sentence similarity measurement which uses Siamese neural networks was developed by Nilaxan and Ranathunga [244], where they demonstrate their results for Sinhala and Tamil.
 A cross-lingual document similarity measurement using the use-case of Sinhala and English was developed by Isuranga et al.
@@
 Deepal et al.
-[236] proposed to use Siamese network which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity.
-They report an F1 of 0.9041 for their model calculated by cosine distance.
+[246] proposed to use a Siamese network, which is a hybrid of LSTM and CNN to calculate Sinhala sentence similarity.
+They report an F1 of 0.9041 for their model, calculated by cosine distance.
 They discuss how they used a number of publicly available data sets and refined them to be used in their task, nevertheless, they do not publicly release the resulting data set.
+Ranasinghe et al.
+[247] introduced MUSTS134134134https://github.com/TharinduDR/MUSTS, a multilingual semantic similarity benchmark that includes Sinhala.
+They use the Sinhala short sentence similarity data set created by Kadupitiya et al.
+[243] as the Sinhala source for this.
 3.10 Text Classification
@@
 A smaller implementation of Sinhala news classification has been attempted by de Silva [30].
-As mentioned in Section 3.2, their news corpus is publicly available110110110https://osf.io/tdb84/.
+As mentioned in Section 3.2, their news corpus is publicly available135135135https://osf.io/tdb84/.
 Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [249].
@@
 A simple TF-IDF based Sinhala text classification system was proposed by Koralage [264].
-The study by [254] used Genetic Algorithm on Lucene111111111https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
+The study by [265] used Genetic Algorithm on Lucene136136136https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents.
 The approach proposed by Rathnayake et al.
@@
 [110].
-Both their code112112112https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1131131133https://tinyurl.com/42un7a9y, Sinhala News source classification data set114114114https://bit.ly/3PBTyfo, Sinhala News category classification data set115115115https://bit.ly/3AsfK7u, and Sinhala writing style classification data set116116116https://bit.ly/3PARuUS) are publicly available.
+Both their code137137137https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus1381381383https://tinyurl.com/42un7a9y, Sinhala News source classification data set139139139https://bit.ly/3PBTyfo, Sinhala News category classification data set140140140https://bit.ly/3AsfK7u, and Sinhala writing style classification data set141141141https://bit.ly/3PARuUS) are publicly available.
 Faumi et al.
-[136] also worked on Sinhala writing style identification (author identification) using their own data set117117117https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
+[139] also worked on Sinhala writing style identification (author identification) using their own data set142142142https://github.com/cipherdragon/SimpleAA which includes Romanized (transliterated) Sinhala text.
+Ranathunga et al.
+[277] also worked on writing style identification; however their work is specifically on classifying Sinhala answers as AI-generated or human-written.
+They reported 86% accuracy for LSTM running on TF-IDF features.
 The study by Wijayarathna and Jayalal [278] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter.
@@
 [284] introduced Saliency-based token Swap (SSwap), a data augmentation technique to improve Sinhala text classification.
+Wickramasinghe et al.
+[285] proposed a Sinhala news personalisation model built on SinBERT for classification.
 The work by Rathnayake and Asanka [240] uses LLaMA for Named entity classification in Sinhala documents and then runs Fuzzy C-Means clustering on it.
@@
 A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [286] based on their prior work [287].
-A word2vec based tool118118118http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
+A word2vec based tool143143143http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available.
 A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al.
 [288]Demotte et al.
-[276] proposed a sentiment analysis system119119119https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments.
+[289] proposed a sentiment analysis system144144144https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments.
 In the subsequent work [290], they used word similarity to generate a Sinhala semantic lexicon.
@@
 [257]Aththanayaka and Naleer [300] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media.
-The ACTSEA dataset120120120https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
+The ACTSEA dataset145145145https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al.
 [138].
@@
 Then in their follow up work [308] they again used XLM-RoBERTa, this time for sentiment analysis and, on top of that used SHAP [309] and LIME [251] to make the model explainable.
+The short study on the sentiments of Sinhala comments by Mohamed et al.
+[310] reported best results overall to be from XLM-R and SinBERT, while BiLSTM and Logistic Regression showing reasonable results on certain tasks.
 3.12 Hate Speech Detection
@@
 A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [312].
-A feature model and a data set121121121https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al.
+A feature model and a data set146146146https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [131]Sandaruwan et al.
 [313] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques.
@@
 [132] analyses Sinhala hate speech propagation on Twitter.
-They have also published a dataset122122122https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
+They have also published a dataset147147147https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets.
 Rajapaksha et al.
@@
 [329] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes).
+Later, they extended their work [330] by creating a Sinhala meme data set annotated with both the text and visual components.
+They then showed that their SVM-based method obtains 95% accuracy on this data set.
 Wickramaarachchi et al.
@@
 They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data.
-Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook123123123https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter124124124https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets.
+Chavinda and Thayasivam [174] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook148148148https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter149149149https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets.
 They report TwHIN-BERT-base [333] model to have obtained the best results.
@@
 Ultimately, for the case of Sinhala as the target language, they recommend using Assamese, Bengali, Nepali, Urdu, Hindi, or Bodo as the source language for transfer learning this task.
-, title=Multilingual Cyberbullying Detection in Low-Resource Languages and Dialects
 After the release of the SOLD dataset by Ranasinghe et al.
@@
 [341] introduces a series of models named Subasa that are tested on the SOLD dataset.
-Their trained PMLs125125125https://github.com/haturusinghe/subasa-plm and LLMs126126126https://github.com/haturusinghe/subasa-llm are separately publicly available.
+Their trained PMLs150150150https://github.com/haturusinghe/subasa-plm and LLMs151151151https://github.com/haturusinghe/subasa-llm are separately publicly available.
 Ou6t of these models, they report that their Subasa-XLM-R model outperforms even the state-of-the-art commercial LLMs when evaluated on the SOLD data set.
 Ranasinghe et al.
-[327] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task127127127https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
+[342] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task152152152https://hasocfire.github.io/hasoc/2023/task1.html they introduced.
 In completing the task, the study by Narayan et al.
@@
 Dmonte et al.
-[141] used the SOLD data set to verify the new MT-Offense128128128https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142].
+[144] used the SOLD data set to verify the new MT-Offense153153153https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [145].
 They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [137].
@@
 [331] detected fake content in Sinhala youtube videos by comparing their title and description to the audio using LSTM on BART.
+Atthanayake and Ruzaik [350] conducted a shallow comparison between the English Fakedit fake news data set [351] and the LIRNEasia Sinhala fake news data set [129].
 3.14 Word Sense Disambiguation
@@
 [201] have proposed a system named Aruth based on the Lesk Algorithm[357].
-An online tool129129129http://aruth.herokuapp.com/, an API130130130https://bit.ly/3sJEYbS of the algorithm, and code along with data on github131131131https://github.com/jseanm1/aruthSWSD are available.
+An online tool154154154http://aruth.herokuapp.com/, an API155155155https://bit.ly/3sJEYbS of the algorithm, and code along with data on github156156156https://github.com/jseanm1/aruthSWSD are available.
 For the same task, Marasinghe et al.
@@
 Hasan et al.
-[151] introduced the XL-sum132132132https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website133133133https://www.bbc.com/sinhala.
-The M3LS134134134https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
-[152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website135135135https://www.bbc.com/sinhala.
+[154] introduced the XL-sum157157157https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website158158158https://www.bbc.com/sinhala.
+The M3LS159159159https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al.
+[155] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website160160160https://www.bbc.com/sinhala.
 Further, they claim that MT5 [366] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
-The first Sinhala multi-document summarizing dataset, M2DS136136136https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
+Patabandi et al.
+[367] used BART and DistilBART to summarise Romanised Sinhala (Singlish) messages collected from WhatsApp group chats.
+The first Sinhala multi-document summarizing dataset, M2DS161161161https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al.
 [156].
 The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary.
-Thier code and pre-trained models are also available on github137137137https://github.com/KushanMH/m2ds.
+Thier code and pre-trained models are also available on github162162162https://github.com/KushanMH/m2ds.
 3.16 Other Semantic Tools
@@
 [370] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems.
-The Massively Multilingual Speech (MMS)138138138https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
+The Massively Multilingual Speech (MMS)163163163https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al.
 [148] has Sinhala data for the Spoken Language Identification (LID) task.
@@
 [393] claim to have created a Sinhala TTS system with a custom neural network that archives 98% accuracy.
+The work by Senarath [394] discusses using existing TTS systems for Sinhala.
+They especially highlight VAENAR-TTS [395] as the best candidate for this task.
+Nayanathara et al.
+[396], on the other hand, propose to use the VITS architecture [397]164164164https://github.com/jaywalnut310/vits for the same task.
 Anuradha and Thelijjagoda [398] proposed a machine translation system to convert Sinhala and English Braille documents into voice.
@@
 TacoSi introduced by Arachchige and Weerasinghe [427] is based on Tacotron [428] and has been evaluated with 10 human evaluators to determine its text-to-speech quality.
-Nanayakkara [407], Nanayakkara and Weerasinghe [408] used DeepSpeech139139139https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
+Nanayakkara [429], Nanayakkara and Weerasinghe [430] used DeepSpeech165165165https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition.
 Gunarathne et al.
-[409] used an earlier version of the CMUSphinx toolkit140140140https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
+[431] used an earlier version of the CMUSphinx toolkit166166166https://cmusphinx.github.io/ to transcribe Sinhala speech to text.
 The later work by Akesh and Meegama [432] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [433, 434] to automatically generate Sinhala subtitles from Speech.
@@
 [436] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking.
-tha [415] have introduced a Sinhala text to Speech dataset141141141https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes.
+tha [437] have introduced a Sinhala text to Speech dataset167167167https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes.
 As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [438].
-While the data set link points to a github repository, it does not contain the data set; it points to a project website142142142https://sita-speakerdiarization.github.io/ which then points to a download page143143143https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
+While the data set link points to a github repository, it does not contain the data set; it points to a project website168168168https://sita-speakerdiarization.github.io/ which then points to a download page169169169https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
+Omnilingual ASR Team et al.
+[439] includes Sinhala170170170https://huggingface.co/datasets/facebook/omnilingual-asr-corpus/tree/main/data/sin_Sinh among the 1600+ languages for which they have created a multilingual ASR system171171171https://github.com/facebookresearch/omnilingual-asr.
 3.17.3 Speech-to-Speech
@@
 The work by Athas and Pirapuraj [442] uses Google APIs (speech-to-text API, text-to-text-translator API, and text-to-speech synthesizer API) to achieve voice translation for Sinhala and Tamil.
+Dilshani and Chandrasena [443] have conducted a survey on Speech-to-Speech systems available for Sinhala and Tamil translation.
+In their speech-to-speech chatbot, Weerakoon et al.
+[444] uses Wav2Vec2-BERT [445] for ASR and VITS [397]172172172https://github.com/jaywalnut310/vits model for TTS.
+The Text-to-Text chatbot component is the middle is handled by an LLM with RAG.
 3.17.4 Speech-to-Intent
@@
 Luger et al.
-[423] introduced a data set named unsupervised_peoples_speech144144144https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data.
+[449] introduced a data set named unsupervised_peoples_speech173173173https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data.
 The Sinhala speech classification system proposed by Buddhika et al.
@@
 3.18 Optical Character Recognition Applications
-While it is not necessarily a component of the NLP stack shown in Fig 3, which follows the definition by Liddy [36], it is possible to swap out the bottom most phonological layer of the stack in favour of an Optical Character Recognition (OCR) and text rendering layer.
+While it is not necessarily a component of the NLP stack shown in Fig 3, which follows the definition by Liddy [38], it is possible to swap out the bottom-most phonological layer of the stack in favour of an Optical Character Recognition (OCR) and text rendering layer.
+Gunathilaka et al.
+[459] introduced SinOCR174174174https://github.com/SriDoc/datasets, consisting of 100,000 images which include 1,135 handwritten Sinhala texts175175175https://drive.google.com/file/d/16X6lAzinZTIgiy6iw8Z0mLEyPyEwZJkL/view?usp=drive_link and texts printed in 200 different Sinhala fonts176176176https://drive.google.com/file/d/1uuPNN8M4jHLuZLF94CApkiG3kysDALia/view?usp=drive_link.
+They also released SinFUND177177177https://drive.google.com/file/d/1hfAfCJcZciEqMgO_wCuvTQ4pWI8t0L5D/view?usp=drive_link, a fully annotated dataset of 100 manually filled Sinhala forms.
 3.18.1 Printed Text
-The XTREME-UP145145145https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP178178178https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [147] contains a Sinhala data set for the OCR task.
@@
 Anuradha et al.
-[458] used Tesseract 3146146146https://tesseract-ocr.github.io/ [459] for Sinhala OCR.
+[485] used Tesseract 3179179179https://tesseract-ocr.github.io/ [486] for Sinhala OCR.
 The later study by [487] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.
-Maduranga and Jayalal [461] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters.
+Maduranga and Jayalal [488] used an Artificial Neural Network (ANN) based on Universe of Discourse and Self-Organisation Map methods to recognise multi-style printed Sinhala characters.
 The study by De Silva et al.
@@
 Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents.
-The work by Thamarasee and Surendra [463] used histogram-oriented gradient descriptor (HOG descriptor) [464] and SVMs to recognize 400+ variations of Sinhala characters.
+The work by Thamarasee and Surendra [490] used a histogram-oriented gradient descriptor (HOG descriptor) [491] and SVMs to recognise 400+ variations of Sinhala characters.
+After conducting a survey of existing methods [492], Hulathdoowage and Kumara [493] proposed a U-Net [494] based system for Sinhala printed document layout analysis.
+For this task they have created a data set using Sinhala textbooks.
+However, they do not provide a public link to this data set.
+Jayatilleke and de Silva [495] conducted a comparative analysis on zero-shot Sinhala OCR using a synthetic Sinhala OCR data set180180180https://huggingface.co/datasets/Ransaka/sinhala_synthetic_ocr-large.
+They concluded that Surya181181181https://github.com/VikParuchuri/surya to be the best Sinhala OCR system in this configuration.
+This aligns with the previous short study conducted by Velayuthan and Ambegoda [496].
+However, later in their subsequent work, Jayatilleke and de Silva [497], they showed that for real-world applications Google Document AI182182182https://cloud.google.com/document-ai/ obtains better results for Sinhala.
 An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al.
@@
 [522] used Gabor-initialised CNN (GCNN) on a dataset of 6000 handwritten Sinhala character images.
+Even though the work by Ramanayake and Weerakoon [523] applies OCR on Sinhala handwritten text, its ultimate objective is not recognising the text itself, but the gender of the author.
+For this, they use several classical ML models as well as GoogleNet and ResNet.
 3.18.3 Ancient Text
-Summarizing optically recognized old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al.
+Summarising optically recognised old Sinhala text for the purpose of archival search and preservation was explored by Rathnasena et al.
 [524].
@@
 Building upon the architecture proposed by Ruwanmini et al.
-[492], a neural network-based method for recognizing ancient Sinhala inscriptions was proposed by Karunarathne et al.
+[526], a neural network-based method for recognising ancient Sinhala inscriptions was proposed by Karunarathne et al.
 [527].
-The study by Wickramarathna and Ranathunga [494] created a system to recognize Brahmi characters, correct errors, and generate Sinhala meanings.
+The study by Wickramarathna and Ranathunga [528] created a system to recognise Brahmi characters, correct errors, and generate Sinhala meanings.
 Heenkenda and Fernando [529] used Inception-v3 [530], VGG-19 [531], and ResNet-50 [532] to classify Sinhala inscriptions to historical time periods.
+They later extended their work [533] by proposing to use YOLOv5183183183https://github.com/ultralytics/yolov5 to detect ancient Sinhala inscriptions.
 The work by Gunasekara et al.
-[499] claims to have created a mobile app that is capable of recognizing and translating early Brahmi characters.
+[534] claims to have created a mobile app that is capable of recognising and translating early Brahmi characters.
 The work by Surasinghe and Thanikasalam [535] used CNN models to predict the historical period of ancient Sinhala text.
@@
 [545] compared the performance of a number of MT models on Low-Resource languages, including Sinhala.
+De Silva and Hansadi [546] have also conducted another shallow study on rule-based, statistical, neural, and hybrid MT models for Low-Resource languages.
 As mentioned in Section 3.1, the study by [105] raised questions on the quality of the existing Sinhala-English corpora.
@@
 Kugathasan and Sumathipala [570] proposed an NMT system for Sinhala-English Code-Mixed text using the standardized Sinhala Code-Mixed text they proposed earlier [571].
+Later, they published a slightly different version of the work at a different venue but under the same title [572].
 Nguyen et al.
@@
 [577] used transfer learning and back translation as well as focal loss of the Sinhala-English dataset from FLoRes [101] and reported better results than mBART [578].
-The study by Ranathungaa et al.
+The study by Ranathunga et al.
 [579] presents an evaluation of using parallel data from auxiliary domains to enhance EnSI NMT by fine-tuning or further pre-training the models.
@@
 However, in actuality, what their experiments are on is mBART.
+The work by Zebaze et al.
+[581] introduces a new NMT pardigm that they call CompTra (compositional translation) which is based on LLMs.
+They report the results for Sinhala using the FLORES-200 [137] data set, where CompTra applied on NLLB [137] gives the best result.
 Fernando et al.
@@
 3.19.3 Singlish to Sinhala (Transliteration)
-The XTREME-UP147147147https://github.com/google-research/xtreme-up data set created by Ruder et al.
+The XTREME-UP184184184https://github.com/google-research/xtreme-up data set created by Ruder et al.
 [147] contains a Sinhala data set for the transliteration task.
+Perera and Sumanathilaka [589] introduced a data set185185185https://github.com/Sameera2001Perera/Romanized-Sinhala-Transliteration-Disambiguation-Dataset to disambiguate Sinhala transliteration.
+The data set contains 600 sentences across 22 ambiguous Romanised Sinhala words.
 The early work by Goonetilleke et al.
@@
 [602] claim to have released the Swa-Bhasha Dataset.
+Swa-bhasha Resource Hub [603] was then created by putting the previous work together for easy access.
 Amarasekara et al.
@@
 They discuss in detail how the zero-width joiner (ZWJ, U+200D) character cause issues in this task and Sinhala rendering as a whole.
-Their code is available on GitHub148148148https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
+Their code is available on GitHub186186186https://github.com/google-research/google-research/tree/master/context_aware_transliteration.
 Khiu et al.
@@
 They reported across-the-board better performance with the latter.
-All thier code as well as the test-sets are publicly avaialble149149149https://github.com/kasunw22/Sinhala-Transliterator/.
+All thier code as well as the test-sets are publicly avaialble187187187https://github.com/kasunw22/Sinhala-Transliterator/.
 Perera et al.
 [615] attempted the same task with BERT.
-Their code is also publicly available150150150https://github.com/Sameera2001Perera/Singlish-Transliterator.
+Their code is also publicly available188188188https://github.com/Sameera2001Perera/Singlish-Transliterator.
 In their short survey discussing Indo-Aryan Languages machine translation and transliteration, Perera and Sumanathilaka [585] covers a number of works [557, 613, 591, 592, 594, 596, 595, 598, 601, 602] on Sinhala transliteration which we have already discussed above.
+The survey by Deshpande and Deshpande [616] covers the works by Perera et al.
+[615] and De Mel et al.
+[612] for Sinhala transliteration, which we have discussed above.
 3.19.4 Singlish and English (Translation via Transliteration)
@@
 [651].
-A rule-based Sinhala spell checker named SinSpell based on Hunspell151151151http://hunspell.github.io/ was introduced by Liyanapathirana et al.
+A rule-based Sinhala spell checker named SinSpell based on Hunspell189189189http://hunspell.github.io/ was introduced by Liyanapathirana et al.
 [652].
-They have also made the tool available152152152http://nlp-tools.uom.lk/sinspell/ online for use.
+They have also made the tool available190190190http://nlp-tools.uom.lk/sinspell/ online for use.
 The study by Sithamparanathan and Uthayasanker [653] extended the Generic Environment for context-aware spell correction to handle Sinhala and Tamil.
@@
 [393] claim to have developed a Sinhala Spell checker with the Random Forest algorithm that archives 82% accuracy.
+[657] introduced LMSpell191191191https://anonymous.4open.science/r/lm-spell-A2D1, an LLM-based toolkit for Sinhala spelling correction.
 A model for detecting grammatical mistakes in Sinhala was developed by Pabasara and Jayalal [658].
@@
 A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al.
-[625] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA153153153https://rasa.com/ [626] chat bot.
+[667] where they used, the fasttext model trained by Facebook [123, 124, 125], on a RASA192192192https://rasa.com/ [668] chat bot.
 A Sinhala chat bot for train information was proposed by Harshani [669].
@@
 [676] is a Sinhala chatbot that can answer patient questions on orthopantomography radiology reports which were previously generated by the system itself by analysing dental panoramic tomography (DPT) images.
+Kandamby et al.
+[677] introduced Cloud Era, a multilingual chatbot utilising LightRAG [678] for agentic behaviour in multi-cloud setups.
+They claim to have achieved 81% recall for Sinhala.
+The work by Weerakoon et al.
+[444] set out to build a Sinhala speech-to-speech chatbot.
+They select Wav2Vec2-BERT [445] for ASR and utilise a RAG approach for the chatbot.
+The TTS component is handled with the VITS [397]193193193https://github.com/jaywalnut310/vits model.
 3.22 News and/or Social Media Recommendation
@@
 It should be noted here that the translation they describe here is text-to-sign language translation and not translation between any two or more natural languages.
+Thrimahavithana et al.
+[719] used 3D avatars to show Sinhala sign language on screen.
+Their system was evaluated with a usability score, error rate in text-to-sign language conversion, and overall user satisfaction.
 The work by Idushan et al.
@@
 Rishan et al.
-[642] uses Leap Motion Controller154154154https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
+[686] uses Leap Motion Controller194194194https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text.
 Dissanayake et al.
 [690] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
+Jayasekara et al.
+[725] uses EfficientNet-B0 model [726] to detect static Sri Lankan Sign Language.
 3.24 Sinhala Braille
@@
 The survey by Perera and Sumanathilaka [755] discusses the lack of annotated datasets and NLP resources to cater for Sinhala-speaking adults with dyslexia.
+The work by Perera and Sumanathilaka [756] created a pipeline with: Whisper for speech-to-text, SinBERT for dyslexic error detection, mT5 and Mistral to generate corrected text, and finally GoogleTTS (gTTS) for text-to-speech.
 3.28 Embeddings, LLMs, and Other Language Models
@@
 Dhananjaya et al.
-[107] has published two Sinhala BERT models: SinBERT-large155155155https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small156156156https://huggingface.co/NLPC-UOM/SinBERT-small.
-A smaller BERT model named SinhalaBERTo157157157https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
+[110] has published two Sinhala BERT models: SinBERT-large195195195https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small196196196https://huggingface.co/NLPC-UOM/SinBERT-small.
+A smaller BERT model named SinhalaBERTo197197197https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper.
 This BERT model has been created by training vanilla RoBERTa [758] on the OSCAR [158, 159] dataset.
 Demotte et al.
-[276] have released the word embeddings158158158https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code159159159https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer.
+[289] have released the word embeddings198198198https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code199199199https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer.
 Gurgurov et al.
@@
 Gurgurov et al.
-[712]160160160https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs.
+[760]200200200https://github.com/d-gurgurov/Knowledge-Driven-Adaptation-LLMs argued that mLMs such as mBERT and XLM-R are better suited for low-resourced languages such as Sinhala as opposed to LLMs and suggested parameter-efficient adapter-based methods for adapting such mLMs.
+Enevoldsen et al.
+[761] introduced MMTEB201201201https://github.com/embeddings-benchmark/mteb, a multilingual embedding benchmark along with a leader board202202202https://huggingface.co/spaces/mteb/leaderboard.
+They used the news classification data set created by de Silva [30] for Sinhala.
+Ranasinghe et al.
+[762] introduced SINHALA-GLUE203203203https://github.com/Sinhala-NLP/Sinhala-GLUE, a language understanding benchmark for Sinhala, along with three new trained Sinhala embedding models: Sinhala-BERT (Raja), Sinhala-RoBERTa (Koliya), and Sinhala-Electra (Mahasen).
+These models are not publicly available.
 Wickramasinghe and de Silva [151] used Sinhala-English dictionary data [150] for obtaining Sinhala-English embedding alignment.
 Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese.
-The relevant code and data can be accessed on GitHub161161161https://bit.ly/3t3SKu7.
+The relevant code and data can be accessed on GitHub204204204https://bit.ly/3t3SKu7.
 Subsequent work by Rathnayake et al.
@@
 They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [765] archives the best result of 29.49 for En-Si FastText embedding.
+The following work by Wickramasinghe and de Silva [766] compared multiple BLI techniques across different language pairs and showed that LaBSE [767] with cross-domain similarity local scaling (CSLS) [768] generally obtained the best result.
+Further, they proposed Vocabulary Pruning and stemming when one or both of the languages considered is highly inflected as is the case with Sinhala.
+Yu [769] uses FLORES-200 [137] data set to show that their cross-model semantic alignment method is superior to CLIP [770] and ALIGN [771].
+They specifically point out Sinhala as one of the languages that shows the most improvement for recall@1 and Mean Rank.
+Pramodya et al.
+[772] released, SinhalaMMLU, a data set of more than 7000 Sinhala questions and answers covering secondary to collegiate education levels in Sri Lanaka.
+They used this dataset to evaluate 26 LLMs and reported that Claude 3.5 achieved the highest average accuracy of 67%.
 Üstün et al.
-[171] introduced Aya, a massively multilingual generative language model162162162https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159].
+[173] introduced Aya, a massively multilingual generative language model205205205https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [162].
+Amarakoon and Rajapakse [773] further fine-tuned the Aya model using Sri Lankan business and regulatory datasets to automate multilingual customer interactions and generate context-sensitive business insights in Sinhala and Tamil for Sri Lankan Small and Medium Enterprises (SMEs).
+Aravinda et al.
+[774] introduced SinLlama206206206https://huggingface.co/polyglots/SinLlama_v01, the first pre-trained Sinhala LLM-base model.
+They used continual pre-training with a 10-million-sentence Sinhala corpus on Llama-3-8B to obtain this model.
+They show the utility of this model with three text classification tasks.
+Further, for this task, they created a Sinhala tokeniser, which is also publicly available207207207https://huggingface.co/polyglots/Extended-Sinhala-LLaMA.
 Jayakody and Dias [775] conducted a comparative analysis on LLMs for Sinhala.
-In it, they observed that the small (7̃B) models of Claude163163163https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o164164164https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)165165165https://llama.meta.com/docs/get-started/ and Mistral 7B166166166https://docs.mistral.ai/.
-Further, they reported that they had to exclude Falcon 2167167167https://falconllm.tii.ae/ and Gemini168168168https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
+In it, they observed that the small (7̃B) models of Claude208208208https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o209209209https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)210210210https://llama.meta.com/docs/get-started/ and Mistral 7B211211211https://docs.mistral.ai/.
+Further, they reported that they had to exclude Falcon 2212212212https://falconllm.tii.ae/ and Gemini213213213https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
 Hansaka and Sumanathilaka [776] also claims to have conducted a review on fine-tuning approaches for LLMs such as GPT-3, GPT-4, and LlaMA in the context of Sinhala.
 Chang et al.
-[718] introduced Goldfish, a collection of monolingual language models169169169https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109].
+[777] introduced Goldfish, a collection of monolingual language models214214214https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [163], NLLB [137], and Glot500 [112].
 The work by Ayesha et al.
@@
 They provide the framework for testing with the results on eight low-resourced languages, which include Sinhala.
+In their work to overcome data sparsity for LLMs, Lai et al.
+[781]215215215Later elaborated as Lai [782].
+tested the Sinhala text generation and understanding results of Google Translate against LLaMA, BLOOM, and several other variations.
+Even though they point to a repository of the instruction data set they created216216216https://github.com/boschresearch/ACL24-MLLM, currently, the data set cannot be found at the given URL.
+Puranegedara et al.
+[783] proposed an intermediate layer fusing method to LLMs so that they would perform better for low-resource languages such as Sinhala.
+They show their improvement with Sinhala News Results on the data set created by de Silva [30].
 Vayani et al.
@@
 This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category.
-This dataset, which they named ALM-Bench170170170https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available.
+This dataset, which they named ALM-Bench217217217https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available.
 They made a number of observations in this aspect.
@@
 In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
+Shafique et al.
+[785] when introducing ViMUL-Bench, a multimodal video Q&A benchmark for LLMs that includes Sinhala, show that GPT4o has the best results for Sinhala.
 3.29 Text Simplification
+The MultiLS data set218218218https://huggingface.co/datasets/MLSP2024/MLSP2024 introduced by Shardlow et al.
+[786] for MLSP 2024 Shared Task219219219https://github.com/MLSP2024/MLSP_Data includes a Sinhala text simplification data set with data from News and Religious domains.
 Ranathunga et al.
-[723] introduced SiTSE, a Sinhala text simplification dataset and the associated benchmark.
-They claim that the data set is publicly available on github171171171https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation; however, currently it resolves to a 404.
+[787] introduced SiTSE, a Sinhala text simplification dataset which is publicly available on github220220220https://github.com/brainsharks-fyp17/Sinhala-Text-Simplification-Dataset-and-Evaluation.
+It has 1000 complex Sinhala texts along with two separate simplifications for each of them.
+Further, they provide the evaluation metrics that can be used to evaluate any subsequent research conducted using the same data.
 The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [788] included the NSINA [111] Sinhala corpus.
-The aggregated data from the shared task is available online as Train-Test data172172172https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data173173173https://huggingface.co/datasets/MLSP2024/MLSP2024.
+The aggregated data from the shared task is available online as Train-Test data221221221https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data222222222https://huggingface.co/datasets/MLSP2024/MLSP2024.
 Enomoto et al.
@@
 As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
+Even though the main focus on the work by Nohejl et al.
+[793] is introducing a Japanese text simplification data set and a model, to demonstrate the generalizability of the system, they use the MultiLS data set [786] which includes Sinhala.
 3.30 Tokenizing
@@
 The work by Velayuthan and Sarveswaran [797] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala.
-Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available174174174https://github.com/Buddhilive/sltk.
+Ranasinghe [166] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available223223223https://github.com/Buddhilive/sltk.
 However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words.
-While the authors have made the dictionary publicly available175175175https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries.
-So while it may be adequate to handle Sri Lanka-centric news title data set176176176https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
-3.31 Detection or Management of Mental Illnesses
+While the authors have made the dictionary publicly available224224224https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries.
+So while it may be adequate to handle Sri Lanka-centric news title data set225225225https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
+3.31 Detection or Management of Illnesses
 Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [745] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook.
@@
 [802] used a NAO Robot [803] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children.
+Continuing their work, Jayawardena et al.
+[804] used the RASA framework [668] to build a dialogue system that provides speech therapy to Sinhala-speaking Children with Autism Spectrum Disorder (ASD).
+Further, Manukalpa et al.
+[805] extended this work with cultural alignment.
 The system by Kulasooriya et al.
 [806] used SVM and simple ANN to detect psychological stress in Sinhala speech.
+Rathnayake et al.
+[807] used voice classification by biomarkers to detect Parkinson’s disease with classical machine learning algorithms.
 3.32 Sinhala Text Generation
@@
 [164] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala.
-They have publicly released both their data177177177https://huggingface.co/datasets/afaji/cvqa and leader board178178178https://eval.ai/web/challenges/challenge-page/2305/overview.
+They have publicly released both their data226226226https://huggingface.co/datasets/afaji/cvqa and leader board227227227https://eval.ai/web/challenges/challenge-page/2305/overview.
 Vayani et al.
-[164] created a dataset ALM-Bench179179179https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web.
+[167] created a dataset ALM-Bench228228228https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [168] and extended it by searching for images under country name, language name, and cultural category on the web.
 They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [169] is superior to other tested LLMs.
@@
 [676] is claimed to be capable of first, generating radiology reports for orthopantomography (related to dentistry) using Llama on captions generated by a Blip-2-based caption generator on dental panoramic tomography (DPT) images and then serve a Sinhala chatbot where the patients may ask questions on the said report and get answers.
+Shafique et al.
+[785] introduced ViMUL-Bench a multimodal video benchmark for LLMs that includes Sinhala.
+With 49.6, they show that GPT4o has the best results for Sinhala; however, the Sinhala accuracy is the lowest of all the tested languages across almost all the LLMs.
 3.34 Language Identification (LID) Task
@@
 [821] introduces a tool named GlotScript which is used for written LangID it achieves above 80% accuracy for Sinhala.
+Van Der Goot [822] discussed the current open challenges in LangID task with 2000+ languages.
+However, due to them only having the Sinhala language written with Sinhala characters, all models in their experiment obtained 99%+ F1 vlaues for Sinhala with Textcat [823]229229229https://www.let.rug.nl/vannoord/TextCat/ and GLOT500 [112] obtaining 99.9%.
+It should be noted that in the original GLOT500230230230https://github.com/cisnlp/Glot500 paper, Imani et al.
+[112] only achieved 45.0% accuracy for Sinhala.
 3.35 Miscellaneous Applications
@@
 Kumari and Hettiarachchi [830] proposed an algorithm for Sinhala topic modelling based on LDA [275] and RAKE [831]Batawalaarachchi [832] proposed two methods using statistical features to select words to be included in Sinhala document titles.
-The study by pal [763] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results.
+The study by Pallawala and Haddela [833] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results.
 Arambewela et al.
@@
 used NER to cluster Sinhala news articles.
-Subasinghe and Samarawickrama [773] analysed the Noto Sans Sinhala180180180https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens.
+Subasinghe and Samarawickrama [843] analysed the Noto Sans Sinhala231231231https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens.
 Zhao et al.
@@
 While they discussed how the LLM system helped their goals and briefly touched on how they utilised a hybrid translation approach of NMT and rule-based methods to circumvent the lack of training data, the specifics of their LLM system are not given.
+Wijerathne et al.
+[846] introduced ScheduleMe an LLM multi-agent system to manage a Google calendar.
+They reported a success rate of 70% for Sinhala.
+It should be noted that among the languages tested by them, only Chinese achieved a worse result (65%) than Sinhala.
+Jayatilleke and de Silva [497] introduced SiDiaC232232232https://github.com/NeviduJ/SiDiaC, the first Sinhala diachronic corpus.
+It covers 58k tokens across 46 Sinhala literary works from the 5th to the 20th century CE under Religious, History, Poetry, Language, and Medical genres.
+The work by Jayasiri and Peellawalage [847] attempted to identify Sinhala song genres by classifying their Mel-frequency Cepstral Coefficients (MFCCs) [848] spectrogram images using CNN models such as VGG-19, InceptionV3, EfficientNet, and DenseNet.
 4 Primary Sources
@@
 In addition to that seminal work, previous researchers of Sinhala NLP have utilized a number of other dictionaries of various configurations such as: English-Sinhala [189, 185, 188], Sinhala-Sinhala [184, 187], and English-Sinhala-Tamil [186].
-A number of NLP applications have utilized first sources intended to teach children [790, 791, 792, 793, 794] or foreigners [795, 221, 796]181181181Note that [796] is an extension of [221]..
+A number of NLP applications have utilized first sources intended to teach children [863, 864, 865, 866, 867] or foreigners [868, 230, 869]233233233Note that [869] is an extension of [230]..
 This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.
@@
 Figure 10: The Probability of research from an institution citing that of other institutions.
-Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext183183183https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
+Note that the calculations are limited by 4 factors: (1) The availability of free to download pdf of the paper, (2) The aforementioned pdf containing references list (Some extended abstracts do not come with the references list), (3) The text extraction capabilities of pdftotext235235235https://www.xpdfreader.com/pdftotext-man.html, (4) The accuracy of the research paper title look up.
 With all those limitations in mind, we still can make a few Interesting observations.